A Complete Bibliography of Applied Statistics: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 4.05

Title word cross-reference

1 [777]. 2^{n} [73]. 2 × 2 [621, 89]. 6 [255]. _{N}C_{R} [462]. χ^{2} [465, 39, 854, 871, 872]. K [863, 779, 353]. L_{1} [879, 450, 346, 511, 498]. L_{p} [781, 590]. n [233, 121]. R × C [841]. S [204, 215]. T [874, 883, 875, 2, 747, 299, 427, 31, 782, 878, 382, 783, 124]. T^{2} [153, 421]. τ [339]. U [269, 216]. W [494]. X [65, 115].

*Jr [547].

-Distribution [875, 2, 31]. -line [255]. -Means [863, 779]. -Norm [498, 346]. -Series [65]. -Statistic [216]. -Values [115].

10 [5, 117, 93]. 100 [520, 856]. 101 [554].

102 [555]. 103 [556]. 104 [557]. 105 [585]. 106 [872, 586]. 107 [587]. 108 [746, 879, 588]. 109 [886, 887, 589]. 110 [781, 590]. 111 [591]. 112 [618, 690, 877, 857]. 113 [619]. 114 [620]. 115 [621]. 116 [648]. 117 [649]. 118 [650]. 119 [651]. 120 [686]. 121 [687]. 122 [688, 860]. 123 [689]. 124 [710]. 125 [862, 711]. 126 [712, 865]. 127 [713, 866]. 128 [714, 881]. 129 [715, 858]. 130 [741]. 131 [884, 742]. 132 [743]. 133 [744]. 134 [777]. 135 [778, 870]. 136 [863, 779]. 137 [780]. 138 [806]. 139 [807, 852, 855]. 140 [808]. 141 [843, 864, 809]. 142 [839]. 143 [840]. 144 [841]. 145 [842]. 147 [882, 853, 849, 876]. 153 [851, 871, 872]. 155 [854, 872]. 16 [880, 7, 593]. 169 [873, 868, 859]. 1965 [290]. 1968 [377].

1970 [642]. 1976 [575].

21 [10, 95]. 26 [30, 246]. 27 [31]. 28 [32]. 29 [50, 522]. 2nd [833, 796, 705, 703, 706, 212, 772, 768, 641, 609, 583, 614, 584, 611].

30 [217, 51, 94]. 31 [52]. 32 [68, 882, 876]. 33 [69]. 34 [70]. 35 [88, 138, 218, 622]. 36 [89]. 37 [90, 319]. 38 [91]. 39 [92]. 3rd [770, 185].

4 [1, 9, 53, 299]. 40 [113]. 41 [114]. 42 [115]. 43 [116, 139]. 44 [134]. 45 [135, 270]. 46 [136]. 47 [749, 137, 349]. 48 [162]. 49 [163]. 4th [64, 288, 263].

5 [875, 2, 299]. 50 [164]. 51 [193, 521]. 52 [194]. 53 [214]. 54 [215]. 55 [216]. 56 [242]. 57 [243]. 58 [244, 417]. 59 [245]. 5th [285].

6 [8, 748, 386, 867, 3]. 60 [877, 267, 320]. 61 [268, 652]. 62 [269]. 63 [589, 295, 238]. 64 [589, 296]. 65 [885, 297, 878]. 66 [298]. 67 [314]. 68 [315]. 69 [316].

7 [8, 867, 4, 829]. 70 [317]. 71 [318]. 72 [344]. 73 [345, 861]. 74 [346, 498]. 75 [558, 381]. 76 [844, 883, 874, 747, 782, 878, 382, 783]. 77 [383]. 78 [384, 467]. 79 [385].

80 [413]. 81 [414]. 82 [415]. 83 [416]. 84 [434]. 85 [435]. 86 [436]. 87 [592, 437]. 88 [462]. 89 [463].

90 [464]. 91 [465]. 92 [466]. 93 [494]. 94 [495]. 95 [496]. 96 [497]. 97 [517]. 98 [518, 745]. 99 [519, 856].

A. [830, 705, 412, 491, 212, 670, 770, 289, 828, 646, 265, 543, 264, 740, 284]. A.I.D. [423]. Abridged [85]. Absence [282, 325]. Absolute [743, 746, 879, 588]. Absorption [314]. Abstract [81]. Abstracts [685].

Accelerated [757]. Acceptance [98, 311]. Account [264]. Accounting [130]. Accounts [280]. Acheson [460]. Action [338]. Activity [179]. Actuality [525]. Actuarial [133]. adaptive [695]. Adby [484]. Addendum [154, 825]. Additive [686, 600]. Adult [576]. Advanced [210]. advances [330]. Adverse [142]. Aesthetic [205]. Affairs [642]. Against [375]. Age [47]. Agriculture [288]. Ahead [307]. Aid [722, 723, 504]. Aids [705]. Aitchison [131]. Alan [81, 609, 544]. Albaum [182]. Albert [804]. Alder [285]. Alderson [539]. Algol [144]. Algorithm [29, 845, 754, 873, 869, 863, 868, 779, 79, 497, 201, 628, 686, 859, 172, 268, 843, 844, 743, 746, 778, 777, 618, 555, 619, 621, 712, 748, 839, 592, 591, 840, 496, 556, 886, 887, 318, 463, 465, 68, 464, 688, 841, 883, 689, 466, 648, 435, 715, 858, 348, 874, 875, 114, 136, 1, 2, 5, 9, 53, 115, 117, 589, 854, 714, 884, 269, 494, 50, 385, 386, 557, 558, 851, 857, 30, 245, 246, 842, 345, 90, 91, 381, 462, 741, 518, 384, 415, 870, 193, 243, 779, 344, 4, 3, 713]. Algorithm [52, 88, 92, 138, 218, 298, 519, 520, 622, 747, 749, 880, 745, 93, 809, 319, 32, 711, 853, 849, 620, 651, 742, 69, 587, 652, 295, 296, 413, 414, 434, 436, 437, 70, 495, 314, 416, 517, 649, 217, 554, 588, 710, 497, 137, 349, 216, 317, 808, 215, 885, 242, 162, 163, 316, 780, 113, 297, 686, 346, 687, 881, 586, 214, 51, 134, 135, 94, 244, 267, 270, 320, 417, 194, 498, 590, 859, 7, 593, 31, 10, 95, 89, 164, 782, 878, 383, 522, 585, 650, 806, 807, 852, 855, 315, 116, 139, 382, 744]. Algorithm [865]. Algorithmic [829]. Algorithms [268, 8, 321, 387, 468, 500, 560, 594, 843, 844, 743, 746, 778, 777, 618, 690, 555, 619, 621, 712, 748, 839, 592, 591, 840, 496, 499, 556, 886, 318, 463, 465, 68, 464, 688, 860, 841, 883, 689, 466, 648, 435, 864, 715, 858, 348, 874, 875, 114, 136, 461, 838, 1, 2, 5, 9, 53, 115, 117, 589, 882, 854, 714, 873, 877, 884, 269, 494, 869, 50, 863, 385, 386, 557, 558, 851, 871, 872, 879, 857, 30, 245, 246, 842, 867, 345, 861, 862, 90, 91, 381, 462, 741, 518]. Algorithms [559, 384, 415, 467, 870, 670, 193, 243, 521, 868, 779, 344, 4, 3, 713, 52, 88, 92, 138, 218, 298, 519, 520, 622, 747, 749, 856, 880, 299, 745, 93, 809, 319, 32, 711, 849, 620, 651, 742, 69, 587, 652, 295, 296, 413, 414, 434, 436, 437, 70, 495, 314, 416, 517, 649, 217, 554, 588, 710, 497, 137, 349, 216, 317, 808, 215, 885, 242, 162, 163, 316, 781, 876, 780, 113, 297, 347, 686, 346, 687, 881, 586, 214, 51, 134, 135, 94, 244, 267, 270, 320, 417, 194, 498, 590, 859, 7, 593, 866]. Algorithms [31, 10, 95, 89, 164, 782, 878, 383, 522, 585, 650, 806, 807, 852, 855, 315, 116, 139, 382, 783, 744, 865, 771, 486]. Alice [262]. Allocation [61]. Alternative [603, 564, 427, 273, 814]. Amalgamation [240]. Amble [515]. among [355]. Analyses [390]. Analysing [332, 200]. Analysis [12, 292, 580, 752, 801, 722, 118, 833, 202, 133, 169, 38, 205, 703, 706, 395, 27, 257, 251, 212, 310, 771, 569, 680, 681, 443, 391, 344, 542, 768, 372, 797, 646, 547, 660, 599, 175, 223, 625, 227, 758, 185, 428, 121, 280, 726, 723, 311, 488, 627, 119, 832, 109, 239, 628, 425, 325, 490, 176, 120, 244, 417, 543, 459, 237, 489, 674, 266, 293, 283, 634, 255, 451, 273, 448, 361, 509, 398, 172, 354, 252, 211, 209, 767, 616, 541, 610, 45]. Analytic [567]. Ancient [446]. Anders [266, 293]. Andersen [830]. Anderson [372, 602, 492]. Angular [453]. Animal [174, 515]. Anne [411]. Announcement [433]. Annual [281, 452]. ANOVA [441, 659]. Ansari [494]. Ansari-Bradley [494]. Anthony [671]. any [172]. Aoki [183]. applicable [71]. Application [759, 292, 58, 374, 119, 264, 531, 266, 293, 505, 788]. Applications [183, 87, 310, 159, 544, 187, 157, 615, 584, 533, 289, 612, 265, 709]. Applied [321, 468, 827, 513, 803, 493, 817, 42, 837, 547, 46, 325, 804, 581, 45]. approach [787, 513, 803, 672, 647, 773, 409, 492].

Approaches [97]. Approximate [791, 504, 650]. Approximating [714, 667, 881]. Approximations [734, 574, 795, 725]. Areas [142, 528, 626]. ARIMA [633]. Arising [181]. Armore [27]. Armstrong [831]. Arrangement [146]. Array [30, 246]. Arrays [92]. Arthur [105]. Artificial [175]. Aspects [150, 813, 514, 444]. ASR [877, 878]. Assays [308]. Assessment [576, 421, 203, 394]. Association [443, 391, 334, 620, 598]. Astronomy [119]. Asymmetric [727]. Asymmetry [261]. Asymptotic [307, 648, 60, 308, 507]. Attendances [443, 391]. Attribute [328]. Australia [302]. Authors [811]. Autocorrelation [163, 412]. Autocovariance [431, 608]. Automatic [423]. Automatique [106]. Autoregressive [752, 97, 431, 608, 201, 664, 531, 507]. Autoregressive-Moving [608]. Auxiliary [1, 9, 53, 299]. Average [52, 587, 431, 608]. Avi [234]. Avi-Itzhak [234].

B [708, 545, 81, 704, 614, 677, 285, 158, 675]. B. [234, 213, 184, 133, 49, 207, 189, 210]. Back [165, 195, 219, 247, 271, 300, 322, 351, 388, 418, 438, 469, 501, 523, 561, 595, 623, 653, 691, 716, 750, 784, 810, 846]. Bailey [835, 584]. Balaam [377]. Ballot [367, 440]. Bancroft [191, 262]. Band [70]. Banks [645]. Barnard [765]. Barrie [40, 583, 41]. Barry [516]. Bartlett [14]. Based [55, 450]. Bases [106]. BASIC [241, 458, 129, 558, 381, 641, 609, 546, 740, 342]. Basis [683]. Batch [758]. Bayesian [454, 278, 254, 526]. be [255]. Beard [683]. Beech [737]. Behaviour [57, 665, 657]. Behavioural [67, 643, 646]. Behaviours [476]. Being [607]. Benjamin [133, 540]. Bennett [643]. Best [91, 156, 626, 726, 353, 72]. Beta [777, 886, 689, 57, 589, 659, 295, 296, 125, 887, 336]. Beta-Binomial [57, 659]. Bethea [540]. Between [538, 79, 537, 260, 393, 786, 512]. Beverley [644]. Bi [229, 228, 370]. bi- [370]. Bi-Directionally [228]. Bi-Variate [229]. Bias [100]. Biases [367]. Bibliography [372, 377]. Bidder [607]. Bierman [674]. Bilateral [819]. Bilinear [830]. Binary [839, 169, 390, 162, 163]. Binomial [699, 332, 179, 445, 701, 715, 858, 57, 659, 52, 314, 315, 850, 200]. Binomial-Positive [179]. Biological [43, 87]. Biologists [48]. Biology [208]. Biomathematics [578]. Biomedical [65, 544, 645]. Biometrical [407]. Biometrics [66]. Biostatistics [799, 833, 490, 191, 373]. Bird [601, 825]. Birmingham [449, 81]. Bissell [705]. Bivariate [718, 883, 874, 747, 127, 357, 782, 878, 382, 783, 471, 274, 331]. Blalock [547, 487]. BLUS [557]. Boat [375]. Bofinger [541]. Bonini [489, 674]. Book [47, 830, 292, 551, 761, 762, 132, 458, 234, 580, 827, 487, 111, 63, 211, 183, 64, 801, 40, 513, 640, 803, 799, 579, 548, 800, 129, 209, 43, 83, 213, 235, 371, 493, 835, 833, 612, 87, 373, 184, 796, 374, 679, 705, 130, 774, 776, 412, 133, 25, 703, 639, 28, 491, 576, 672, 673, 767, 188, 24, 682, 706, 644, 48, 708, 484, 545, 67, 27, 49, 106, 212, 407, 408, 647, 676, 670, 671, 770, 771, 802, 773, 772, 836, 26, 81, 375, 680, 681, 456, 542, 42, 65, 286, 207, 289]. Book [454, 582, 737, 828, 768, 191, 643, 372, 455, 837, 797, 704, 84, 646, 641, 547, 62, 82, 208, 159, 241, 342, 514, 410, 409, 616, 609, 707, 549, 829, 769, 544, 577, 185, 66, 86, 189, 678, 85, 187, 157, 379, 378, 376, 20, 23, 546, 190, 583, 578, 488, 642, 131, 614, 832, 109, 110, 239, 238, 290, 411, 677, 615, 46, 457, 541, 490, 285, 645, 108, 44, 584, 287, 738, 41, 380, 831, 516, 486, 539, 158, 291, 210, 160, 182, 265, 340, 543, 552, 610, 611, 739, 341, 550, 804, 540, 709, 262]. Book [264, 288, 492, 515, 740, 675, 459, 485, 834, 107, 237, 377, 186, 240, 581, 45, 105, 263, 683, 553, 489, 617, 613, 685, 674, 684, 775, 21, 22, 284, 460, 798, 266, 293, 236]. Bore [202]. Boullion [540]. Bourke [577]. Bowden [207]. Bowers [643]. Box [604, 404, 25, 473, 231, 354]. Box-Jenkins [604]. Box-Muller [404, 473, 231]. Bradley [494, 226]. Bradley-Terry [226]. Britain [184, 199]. British [277]. Bronchitis [144]. Brown [262]. Buddy [186]. Budgets [149, 239]. Building [291, 21]. Buishand [769]. Burdette [109]. Bus [697]. Business [612, 641, 146, 489, 613, 674, 684, 775, 22]. Buyer [665]. Buyer-Behaviour [665]. Buying [265]. Byrkit [371].

C [640, 799, 235, 48, 106, 772, 82, 549, 614, 210, 265, 543, 684, 775]. C. [830, 209, 796, 491, 375, 680, 681, 542, 290, 677, 615, 457, 674]. Calculating [883, 874, 747, 782, 878, 382, 783, 258]. Calculation [592, 69, 437, 570, 194]. Calculator [768, 574, 794]. calculators [795]. Calibration [661, 457]. Call [528]. Cambridge [575]. Campbell [48]. can [255]. Cancer [598]. Candidates [85]. Carleen [130]. Carlo [637, 177, 279, 765, 275]. Carlson [380]. Carter [547]. Case [811, 126]. Cases [530]. Catalytic [360]. categorical [786, 402]. Categories [620]. Caulcott [582]. caused [430]. Causes [720]. Cautionary [627]. Cell [332, 570]. Censored [862, 880, 625, 720, 711, 357, 230, 6, 7, 333, 593, 806, 807, 852, 855, 358, 472]. Central [2, 871, 747, 299, 782, 878, 382, 783, 617]. Certain [152, 181]. Chain [222]. Champion [129]. change [787]. change-point [787]. Changeable [758]. Changes [429, 11, 531]. Characteristic [52]. Characteristics [587]. Charles [489]. Charts [550, 368]. Chatfield [796, 542]. Checking [146]. Checklists [836]. Chemical [263]. Chi [725, 786]. Chi-Squared [725, 786]. Children [638, 763]. China [446]. Chirp [649]. Chiswick [516]. Choice [309, 571, 550, 329]. Christ [454]. Christian [240]. Christine [552]. Christofides [209]. Circle [626].

Circular [414]. City [81, 411]. Clark [544]. Class [604, 587]. classical [359]. Classification [619, 106]. Cliff [412]. Clinical [109]. Clinton [799]. closer [814]. Cluster [257, 244, 417, 448]. Clustering [863, 779, 601, 825, 808, 316, 486, 449]. Coates [184]. Cochran [738]. Code [884, 742]. Coder [821]. Coefficient [718, 140, 532]. Coefficients [495, 719, 507]. Colin [289, 828]. Collection [455, 685]. Colquhoun [60]. Column [841, 397]. Combination [854, 871, 872]. Combinations [462, 122]. Combining [474]. Comment [882, 792, 206]. Comments [606]. Common [197]. Comparative [639]. compared [200]. Comparing [701, 715, 858, 850, 254]. Comparison [565, 58, 525, 233, 142, 603, 599, 535, 73, 226, 600, 402, 789, 258, 364]. Comparisons [792, 355]. compartment [444]. Compatible [828]. Competing [720]. Competition [563]. Complex [416]. Component [175, 223, 567, 273]. Components [752, 203, 527]. Compound [179]. COMPSTAT [774]. Computation [370, 339, 726, 526]. Computational [387, 72]. Computer [47, 12, 755, 644, 65, 737, 208, 241, 342, 149, 74, 147, 533, 709, 10, 95, 818, 123, 200, 278, 527, 312]. computers [252, 671]. computers-features [252]. Computing [229, 150, 76, 75, 344, 207, 620, 563, 77, 74, 433, 6, 333, 78, 78]. Concentration [734]. Concept [264]. Concurrent [729]. condition [511]. Conditions [142]. Conference [166, 655, 666, 310, 128, 733, 369, 575]. Confidence [621, 197, 506, 508, 554, 89]. Confined [806, 807, 852, 855]. Confounded [636]. Congestion [15]. Congruential [518, 559, 745, 231, 404]. connection [512]. Considerations [211, 76, 428]. Considering [600]. Consistent [281]. Constance [480]. constant [123]. Constrained [335, 879, 350]. Construction [461, 838, 104]. Consulting [528]. Consumer [57]. Contemporary [672, 773]. Contingency [592, 800, 39, 101, 193, 521, 253, 437, 815, 366, 477, 359]. Continuous [55]. Contrasts [636, 482, 668]. Control [188, 802, 645, 291, 460, 330, 789, 40, 640, 188]. controlling [368]. Convenient [478]. Conversational [18]. Convex [61]. Cooper [49]. Correction [690, 887, 850, 339, 467, 521, 443, 608, 393, 783]. Corrections [761, 843, 844, 763]. Correlated [817]. Correlation [592, 718, 648, 474, 140, 437, 605, 719, 100]. Correlations [432]. Correspondence [361]. Corrigenda [8, 321, 9, 138, 853, 94, 95, 139]. Corrigendum [292, 53, 349, 270, 320, 333]. Cost [211]. Cot [658]. Course [545, 284]. Covariance [38, 714, 881, 585]. Covariate [720]. Covariates [862, 711]. coverages [275]. Cox [64, 788, 354]. Craddock [47]. Credibility [148]. Criteria [141]. criterion [71]. Critical [565, 435, 251, 396, 756]. Croft [46]. Crop [17]. Cross [345, 861, 832]. Cross-Spectrum [345, 861]. Cross-Tabulated [832]. Cumulative [667, 574, 55, 104, 403, 795]. Curve [164]. Curves [519, 856, 230]. Customers [475, 313]. cusum [368]. Cyrus [374].

D [551, 762, 513, 209, 87, 412, 491, 737, 739]. D. [63, 801, 579, 800, 371, 212, 643, 704, 185, 23, 677, 46, 288]. Daily [769, 282]. d’Analyse [408]. Daniel [833, 837, 615, 490]. Darling [602]. Darrell [286]. Data [292, 458, 426, 496, 801, 143, 332, 145, 118, 97, 150, 169, 767, 306, 148, 862, 173, 281, 817, 36, 75, 680, 681, 144, 880, 455, 326, 178, 797, 175, 223, 625, 149, 720, 711, 620, 535, 390, 626, 488, 731, 146, 256, 230, 147, 605, 250, 543, 6, 7, 333, 593, 266, 293, 806, 807, 852, 855, 255, 166, 452, 362, 788, 200, 358, 392, 472, 644, 832, 543, 105]. David [458, 835, 188, 647, 454, 376, 578, 110].

Davies [208, 263]. Day [289, 828]. Dealing [105]. Dean [129]. Death [658]. Decision [579, 682, 706, 189, 678]. Decisions [159, 86, 157, 489, 613, 674]. Decomposition [748, 386, 867, 3, 527]. Defining [636, 482, 668]. Definite [867, 4]. deflating [786]. delay [327]. Demand [302, 662, 505]. Demographic [455]. Demography [64]. Dempster [484]. Dendrogram [555]. Dense [753]. Density [17, 780, 658, 834]. Department [642]. Dependent [780, 695]. Derivation [281, 431, 608]. Derived [618, 690, 877, 857, 88, 138, 218, 622]. Derman [374]. Desai [581]. Describe [406]. Description [461, 838, 283]. Descriptions [29]. Descriptive [672]. Design [141, 703, 771, 772, 616, 707, 428, 541, 732, 120, 492, 459, 237, 377, 368, 172]. Designed [660, 305]. Designs [233, 56, 356, 73, 250, 451, 363, 397, 547]. Desmond [188]. detect [396]. Detecting [786, 79]. Detection [722, 423]. Determinant [415]. Determination [534, 633, 440]. Deterministic [695]. Developments [234, 541, 99, 818]. deviates [473]. Deviations [194, 794]. Diagnosis [365, 376]. Diagnostic [530]. Diagram [134]. Diagrams [347]. Dichotomous [306]. Dietrich [457]. Differences [79, 260, 393]. different [313, 509]. Digamma [556]. Digital [671, 207]. Dimensional [744]. Dimensions [632, 92]. Direct [206, 19]. Directed [808]. Directionally [228]. Disaggregation [97, 452]. Discarding [175, 223]. Discrete [759, 416, 517, 649, 564]. Discriminant [722, 723, 627, 365]. Discriminante [408]. Discriminate [176]. Discrimination [565, 814]. Diseases [399, 584]. Dispersion [114, 344, 533]. dispersive [327]. Distances [555]. Distribution [699, 591, 465, 790, 466, 209, 179, 445, 875, 1, 2, 9, 53, 39, 854, 269, 494, 871, 872, 842, 667, 299, 566, 140, 436, 181, 554, 710, 238, 317, 204, 215, 176, 570, 586, 6, 333, 31, 11, 820, 383, 634, 403, 315, 725, 791, 478, 274, 402, 331, 275]. Distribution-Free [466, 554, 710, 176]. Distributions [152, 618, 690, 689, 701, 715, 858, 877, 884, 857, 760, 651, 742, 544, 127, 526, 403, 313, 358, 824, 850, 677]. Districts [240]. Disturbances [664]. Diversity [549]. Divination [446]. Dixon [65]. DO [462]. Donald [614]. Douglas [235, 612]. Down [50, 522]. Draper [25]. Drawing [405, 180]. Drivers [171]. Drugs [538]. Duckworth [28, 770]. Due [367]. Duncan [799, 460]. Duran [540]. Durations [730]. Durbin [851, 871]. Durbin-Watson [851, 871]. Durran [83]. Dutch [78]. Dynamic [63, 548].

E. [184, 24, 82, 707, 549, 285, 683]. Earl [207, 835]. Earnings [292, 148, 266, 293]. Ebdon [647]. Ecological [549, 733]. Ecology [82]. Econometric [776, 110]. Econometrics [516, 553, 581]. Economic [329, 184, 130, 569, 642, 630, 240, 553, 368]. Economics [641, 291, 21, 684, 775]. Economy [107]. ed [185, 285, 212, 288]. Editing [143, 145, 150, 148, 144, 326, 149, 146, 147, 166]. Edition [705, 703, 706, 770, 772, 768, 609, 614, 611, 64, 833, 796, 641, 546, 583, 584]. Editorial [811]. Editors [702, 826, 112, 483, 192, 161, 766, 736, 343, 805, 669, 294]. Edmund [109]. Education [665, 545, 27]. Edwards [264]. Effect [428, 664]. Effects [441, 636, 140, 819, 477, 505]. Efficiency [228, 728, 471]. Efficient [790, 54]. Egg [406]. Ehrenberg [537, 265, 543]. Eilon [209, 26]. elasticity [123]. Elections [367]. electricity [505]. Element [259]. Elementary [803, 548, 41]. Elements [371, 189]. Elliot [676]. Elliott [835]. Ellipse [180]. Ellipses [735]. Ellipsoidal [719]. Empirical [566, 630, 658]. Engineering [235, 772, 62]. Engineers [540, 488]. English-Russian [236]. English/English [236]. Enhanced [741]. Enrick [186]. Enslein [671]. Epidemiological [535]. Epidemiology [316, 539]. Episodes [566]. Equal [115, 629]. Equation [59]. Erdos [84]. Erricker [210]. Error [307, 648, 14, 754, 339, 280, 303, 355, 600, 507]. Error-Rates [754]. Errors [746, 430, 879, 760, 588]. Essays [553, 107]. Essentials [491]. Estimate [592, 437, 472]. Estimated [507]. Estimates [743, 778, 870, 563, 122, 6, 333, 533]. Estimating [697, 362, 424, 206, 601, 825, 19, 315]. Estimation [152, 718, 496, 141, 143, 14, 445, 145, 150, 754, 534, 224, 879, 635, 148, 862, 173, 37, 144, 880, 757, 149, 711, 508, 731, 127, 146, 305, 256, 147, 605, 54, 7, 593, 719, 834, 806, 807, 852, 855, 166, 71, 471, 505, 447, 123, 512, 274, 327, 479, 450, 392, 331, 444, 172, 511]. Estimator [752]. Estimators [261, 177, 279, 60]. Ethical [813, 428]. Euclidean [244, 417]. Europe [642]. Evaluation [518, 559, 745, 314]. Evelyn [582]. Evolutionary [25, 337, 338, 447]. Exact [618, 690, 621, 839, 701, 715, 858, 877, 197, 857, 842, 204, 89, 850]. Examination [85, 657]. Example [118]. Executive [159]. exercise [355]. Exercises [26]. Exeter [411]. expected [363]. Expenditure [149, 290]. Experience [455]. Experiment [332, 599, 121, 221, 541, 377, 200]. Experimental [772, 174, 547, 230]. Experimentalists [49]. Experimentation [615]. Experiments [703, 636, 251, 482, 668, 660, 424, 481, 694, 698, 732, 305, 550, 492, 396, 312, 707, 237]. explicit [512]. Exploration [205]. Exploratory [767]. Exploring [441, 680]. Exponential [862, 711, 206, 19, 358]. Extension [537, 176, 354].

F [679, 705, 547, 241, 342, 410, 829, 187, 457, 264]. F. [373, 379, 411]. Factor [580, 332, 212, 200]. Factorial [251, 694, 732, 283, 396]. Factorials [233, 73]. Factors [59]. Failure [720]. Fairley [708]. Familial [605]. Family [149, 642, 239]. Fast [416, 517, 255]. Faster [207]. Fatigue [707]. features [252]. Federer [377]. Feigl [191]. Feinstein [290]. Fertility [642]. Fertilization [406]. Field [541]. Figures [281]. Fildes [612]. Financial [798]. Finite [345, 861, 88, 138, 218, 622]. Finney [185, 23, 288]. First [673, 545, 664, 531]. First-Order [664, 531]. Fisher [155, 211]. Fit [193, 521, 626, 781, 346, 498, 590, 164, 442, 276]. Fitting [35, 519, 856, 198, 358, 201, 226, 260, 393, 585, 681]. Fitzpatrick [614]. Flow [224]. Fluoridation [598]. Football [443, 391, 334]. Footrule [820]. force [276]. Forecast [525, 695]. forecast-adaptive [695]. Forecasting [575, 302, 724, 612, 330, 835, 831]. Forecasts [568, 630]. Forensic [759]. Forest [753, 66]. Forester [22]. Form [635, 662]. Formalised [379]. Formalization [579]. Format [116, 139]. Forms [586]. Formula [155]. Formulae [885, 297]. Fortmann [802]. FORTRAN [258, 462, 289, 828, 686, 116, 139]. Foundation [833, 490, 240]. Fourier [345, 861, 170, 416, 517, 649]. Fox [108, 550]. Fractions [251]. Fracture [426]. Frank [410]. Fred [772]. Frederick [708, 797]. Free [466, 554, 710, 176]. Freezes [480]. Frequency [884, 842, 651, 742, 215]. Fried [67]. Front [167, 196, 220, 248, 272, 301, 323, 352, 389, 419, 439, 470, 502, 524, 562, 596, 624, 654, 692, 717, 751, 785, 812, 847]. Function [499, 556, 886, 722, 701, 715, 858, 348, 1, 9, 53, 589, 635, 173, 37, 749, 299, 296, 436, 431, 608, 206, 564, 137, 349, 19, 574, 256, 162, 163, 54, 687, 125, 365, 887, 850, 123, 392, 795].

Functional [35, 662, 572, 260, 393]. Functions [600, 403, 744, 447, 186]. Fundamental [545]. Fundamentals [183]. Funke [685]. Further [281, 449, 365].

G [132, 580, 799, 770, 456, 641, 208, 546, 832, 44, 287, 738, 291]. G. [827, 40, 679, 737, 577, 187, 583, 41]. Game [221]. Games [80]. Gamma [68, 573, 822, 882, 816, 853, 849, 876, 336]. Gandy [209]. Gap [98, 311]. Gap-Acceptance [98, 311]. Gardiner [66]. Gary [459]. Gas [16]. Gas-Transmission [16]. Gaussian [402]. Gear [770]. Gehan [109]. GEM [776]. Gene [211]. General [679, 776, 587, 649, 628, 210]. Generalised [341]. Generalization [226, 359]. Generalizations [699]. Generalized [433, 216]. Generating [793, 336, 473]. Generation [132, 755, 777, 573, 462, 713, 866, 818]. Generator [269, 494, 214]. Generators [822, 518, 559, 745, 231, 404]. Geography [647, 184]. Geological [580]. Geomathematics [580]. Geophysical [118]. George [262, 25, 372]. Gerald [454, 62, 182]. Gertsbakh [675]. Gilbert [407]. Gillian [81]. Given [841]. Gleser [374]. GLIM [433]. Glossary [236]. Gokhale [800]. Goldsmith [263]. Goldstein [576]. Gompertz [37, 276, 600]. Goodman [771]. goodness [442]. goodness-of-fit [442]. Gordesch [774]. Government [367]. Graham [832]. Gram [136, 385]. Granger [830]. Graph [808]. Graphical [727, 656, 655, 723, 472]. Graphs [497]. Gray [341]. Greater [777]. Greensted [491]. Groot [86]. Gross [544]. Grossman [835]. Group [723, 789]. Grouped [496, 173, 880, 256, 6, 7, 333, 593, 358, 392]. Growth [817, 206, 564, 19, 600, 278]. Grundy [379]. Guarantee [534]. Guenther [640]. Guide [374, 770, 617].

Gupta [493, 372]. Guttmann [670].

H

[211, 43, 83, 703, 576, 484, 375, 454, 372, 707, 66, 376, 290, 677, 685, 674, 684, 775, 284]. H. [487, 43, 133, 576, 547, 616, 239, 285, 645, 341, 45, 674]. Hahn [62]. Haine [81]. Half [217, 51, 94]. Hall [26]. Halsbury [207]. Hamburg [706]. hand [794, 795]. hand-calculator [794]. Handbook [677, 485]. Harmonic [325]. Harmonics [170]. Harshbarger [682]. Hartigan [486]. Hartkemeier [45]. Harvey [159, 157, 190]. Hausman [674]. Haycocks [133]. Hazard [625, 788]. Health [799, 833, 87, 379, 378, 490]. Healy [576]. Hedetniemi [771]. Height [576, 638, 763]. Heights [537]. Hennes [513]. Herbert [671]. Hertzian [426]. Hess [614]. Hierarchical [619]. Hill [762, 739]. Hills [639]. Hines [235]. Histogram [135, 270]. History [704, 126]. Hit [164]. Hitz [802]. Hoel [641]. Hogg [676]. holiday [505]. Holman [43]. Holt [724]. Holt-Winters [724]. Homogeneous [141]. Honor [262]. Honour [553]. Hormone [308]. Hospitals [376, 614]. Hospitals-Diagnosis [376]. Hotelling [153, 421]. Houthakker [239]. Huang [110]. Huff [286]. Hull [310, 552]. Hurwood [835]. Hyperbola [60]. Hypergeometric [245, 69].

I. [762, 106, 739]. I.B.M. [339]. IASPS [103]. Ideal [155]. Ideas [609]. Identification [366, 142, 510]. Identifying [156]. II [330, 223, 131]. Illness [566]. Illustration [59]. Ilya [675]. Implementing [188]. implicit [512]. Improved [873, 868, 859]. including [255]. Income [238, 290, 287, 11]. Incomplete [886, 68, 589, 882, 853, 849, 295, 296, 876, 887, 395, 331]. Independence [101]. Independent [758, 602, 791]. Index [321, 155]. Indian [638, 763]. Individual [153, 603, 364]. Industrial [26, 615, 147, 460]. Industry [291, 263, 292, 266, 293, 804, 21]. Infection [360, 788]. Infectious [399, 584]. Inference [464, 688, 860, 222, 27, 334, 526, 551, 673, 676, 264]. Infinite [510]. Influence [722]. Information [33, 96, 168, 249, 324, 420, 503, 597, 693, 848, 800, 572, 356, 552]. Ingram [374]. Input [77]. Input/Output [77]. Inspection [40, 640, 329]. Institute [85]. Insurance [146, 683]. Integral [712, 68, 229, 883, 573, 874, 875, 2, 589, 882, 298, 747, 299, 853, 849, 295, 876, 31, 782, 878, 382, 783, 865, 336]. Integrals [1, 9, 53, 299, 370]. Interaction [441, 700, 423, 253]. interactions [395]. Interactive [801, 433]. Interclass [605]. Internal [75]. Interpolating [505]. Interpretation [432, 577, 407]. Interpreting [663, 885, 297, 543]. Intersection [98]. Interval [466, 197, 508, 230, 122]. Interval-Censored [230]. Intervals [554]. Introduction [830, 183, 143, 373, 484, 67, 27, 771, 802, 643, 82, 241, 342, 678, 20, 285, 539, 288, 45, 191, 23, 670, 578]. Introductory [799, 682, 284]. Inverse [886, 887, 589, 667, 757, 508, 296, 574, 795]. Inversion [864, 867, 90, 4, 809, 70]. Investigation [177, 630, 273]. Invitation [811]. Ipsen [191]. Irene [614]. Issacs [454]. Item [458]. iterative [312, 172]. Itzhak [234].

J [830, 63, 579, 129, 373, 374, 576, 188, 27, 670, 65, 641, 62, 544, 577, 23, 832, 109, 239, 541, 584, 516, 288, 105, 460]. J. [47, 580, 83, 412, 703, 491, 576, 836, 26, 737, 409, 577, 678, 131, 677, 541, 831, 486, 22]. Jackknife [341]. Jackson [454]. Jacob [241, 342]. Jamaica [304]. James [513, 684, 775]. Jan [553]. Jardine [491]. Jean [408]. Jean-Marie [408]. Jebe [707]. Jeffers [836]. Jenkins [604]. Jessen [641].

Joan [776, 190]. Johannes [774, 191]. John [458, 827, 548, 703, 767, 797, 514, 237]. Johnson [869, 772, 519, 520, 856]. Johnson-Normal [869, 520, 856]. Joint [666, 651, 369]. Jon [768]. Jones [678]. Joreskog [580]. Judgment [205].

K. [580]. Kapadia [677]. Kendall [318, 339, 456, 204, 215, 291, 610, 611]. Kernel [759]. Keys [732]. King [26]. Kingaby [287]. Kingdom [238, 290]. Kingston [304]. Kirkness [411]. Kleinrock [709]. Klevmarken [266, 293]. Klovan [580]. Kmietowicz [798]. Knapp [799]. Knox [316]. Konijn [616]. Konrad [802]. Kotz [236]. KRANK [339]. Krishnan [547]. Kullback [800]. Kurt [671]. Kurtosis [434].

L [132, 835, 802, 772, 454, 84, 241, 342, 409, 488, 541, 285, 341, 540, 492, 186, 263, 613]. L. [547]. Lake [480]. Language [342, 241]. Languages [74]. Lapin [613]. Large [38, 558, 381]. Large-Scale [38]. Largest [842, 383]. Latent [267, 320, 258]. Law [757, 508]. Lawley [212]. Lawrence [613]. Learning [513]. Least [743, 558, 381, 178, 757, 206, 19, 359, 172]. Lee [737]. Left [152]. Left-Truncated [152]. Lenart [160]. Length [649]. Leon [374]. Leonard [709]. Leone [772]. LEP [475]. Lerman [106]. Leslie [377]. Less [777]. Letter [826, 112, 483, 192, 161, 766, 736, 805, 294]. Letters [702, 343, 669]. Level [636, 482, 668]. Lewes [411]. Lewis [108]. Lie [286]. Life [764, 757, 824]. Likelihood [496, 754, 862, 173, 37, 880, 711, 390, 563, 256, 54, 658, 264, 6, 7, 333, 593, 806, 807, 852, 855, 392, 331]. Limit [181]. Limits [621, 710, 89, 275]. Lindgren [189]. Line [268, 652, 781, 346, 498, 590, 255]. Linear [696, 743, 746, 778, 35, 141, 854, 558, 871, 872, 879, 173, 381, 817, 870, 802, 193, 521, 227, 588, 433, 572, 532, 54, 122, 550, 260, 393, 807, 852, 855, 359, 392]. linearization [401]. Lines [35, 729]. Lines-The [35]. Linkage [555]. List [468, 629]. litter [509]. Litters [174]. Little [707]. Local [367, 819]. Location [261, 496, 279, 533]. Lock [679]. Lockett [770]. Log [193, 521, 359]. Log-Linear [193, 521]. Logistic [564, 600, 365, 402]. Logit [399]. Lognormal [634, 358]. London [287]. Long [760, 831]. Long-Range [831]. Long-Tailed [760]. Longitudinal [36]. Look [700, 673, 814]. Loops [462]. loss [356]. Lowest [607]. Ltd. [291]. Lucy [776]. Lyon [105].

M [47, 458, 487, 513, 548, 184, 576, 768, 547, 159, 157, 190, 411, 108, 540, 237]. M. [799, 87, 703, 576, 24, 484, 456, 287, 291, 182, 550]. MacFarlane [491]. Machin [578]. Machines [228, 102, 207]. MAD [451]. Mail [84]. Maintenance [675]. Major [480]. Making [706]. Makower [24]. Management [209, 235, 150, 159, 829, 378, 46, 665, 26]. Managerial [157, 379]. Managing [108]. Manchester [666]. Mann [269, 216]. Manova [686, 570]. Manual [340]. Many [765, 756]. Map [682]. Mapes [111]. Mapping [753]. Marcel [829]. Marianne [644]. Marie [408]. Market [277, 410]. Marketing [182, 552, 685, 491, 158]. Markov [222, 362, 422, 476]. Marshall [576]. Martin [213]. Masanao [183]. Massey [410]. matched [363]. matched-pair [363]. matches [363]. Mathematical [209, 82, 514, 584, 158, 291, 107, 685, 21, 798]. Mathematicians [23, 43]. Mathematics [111]. Matrices [727, 344, 713, 79, 70, 866]. Matrix [748, 864, 114, 776, 714, 386, 867, 90, 415, 4, 3, 809, 572, 881, 267, 320, 585]. Matter [165, 195, 219, 167, 196, 220, 247, 271, 300, 248, 272, 301, 322, 351, 388, 323, 352, 389, 418, 438, 469, 419, 439, 470, 501, 523, 561, 502, 524, 562, 595, 623, 653, 596, 624, 654, 691, 716, 750, 692, 717, 751, 784, 810, 846, 785, 812, 847]. Maturity [576]. Maurice [610, 611]. Max [778, 870]. Maximum [496, 754, 862, 173, 37, 880, 711, 563, 256, 392, 331, 54, 6, 7, 333, 593, 806, 807, 852, 855]. Maximum-Likelihood [496, 37, 563, 54, 6, 333]. Maxwell [212, 646, 740]. McCall [545]. McGilvray [577]. McLean [492]. McNeil [801]. Mean [307, 114, 197, 344, 537, 303, 194, 507, 472]. Mean-Square [307]. Means [673, 863, 779, 121, 305, 570, 122, 368, 254, 789, 355]. means-results [355]. Measles [360]. Measure [34, 730]. Measurement [487]. Measures [620, 434]. Meddis [485]. median [450]. Mediancentre [840, 384, 467]. Medical [458, 128, 18, 577, 740]. Meghnad [581]. Melvin [454]. Method [405, 718, 307, 229, 43, 753, 576, 197, 212, 663, 758, 601, 825, 260, 393, 473, 361, 254]. Méthodes [408]. Methodology [579]. Methods [565, 759, 292, 580, 400, 612, 656, 655, 671, 62, 200, 829, 378, 583, 532, 540, 515, 263, 613, 684, 775, 266, 293, 512, 814, 484, 379, 110, 41]. Meyer [488]. Michael [639, 539]. Michail [66]. micro [362]. micro-unit [362]. Miller [799]. Min [778, 870]. Min-Max [778, 870]. Minimization [499, 348, 749, 137, 349, 54]. minimize [356]. Minimum [152, 746, 879, 588, 104, 113]. Mink [700]. Mink-Muskrat [700]. Miscellanea [152, 735, 575, 154, 261, 452, 335, 405, 233, 451, 536, 453, 404, 573, 153, 313, 39, 370, 538, 205, 734, 512, 430, 764, 339, 666, 232, 310, 482, 668, 337, 667, 606, 180, 128, 334, 79, 571, 80, 61, 155, 481, 126, 794, 432, 151, 18, 156, 338, 537, 765, 450, 431, 206, 311, 607, 181, 103, 406, 127, 231, 824, 825, 19, 60, 574, 572, 312, 104, 795, 511, 480, 638, 763, 733, 125, 369, 336, 124]. Mises [790, 734, 436]. Missing [306, 660, 424, 250, 71, 172]. Mixed [481, 358, 312]. Mixed-Up [481, 312]. Mixture [731]. Mixtures [689]. MMPI [458]. Mode [177]. Model [304, 721, 661, 179, 57, 399, 232, 360, 681, 603, 720, 508, 563, 686, 291, 226, 480, 21, 585, 807, 852, 855, 507, 788, 278, 422, 398, 276, 444, 199]. Modelling [209, 700, 769, 433, 645]. Models [696, 604, 839, 141, 864, 222, 625, 809, 181, 406, 633, 664, 819, 158, 675, 107, 308, 685, 531, 283, 695, 815, 71, 512, 479, 402, 359, 830, 550]. Modern [378]. Modification [229, 668]. Modifications [453]. Modified [445]. Moment [140]. Moments [519, 856, 274]. Money [662]. Monitoring [259]. Monte [637, 177, 279, 765, 275]. Montgomery [235]. Moore [546, 44]. Morbidity [566]. Morgan [678]. Morris [706, 86]. Mortality [133, 276]. Mosteller [708, 797]. Moving [741, 431, 608]. Muller [404, 473, 231]. Multi [464]. Multi-Variable [464]. Multicollinearity [430, 606]. Multidimensional [243, 477]. Multimodal [744]. Multinomial [842, 254]. Multinormality [421]. Multiple [746, 778, 700, 879, 764, 870, 599, 588, 308, 788, 355]. Multiplicative [231, 600, 404, 398]. Multiresponse [141, 71]. Multivariable [860]. Multivariate [12, 688, 153, 169, 310, 569, 344, 643, 372, 646, 547, 434, 726, 119, 780, 532, 610, 361]. Multiway [32]. Murdick [158]. Muskrat [700]. Myers [186].

N [212, 836, 515, 377]. N. [407, 547]. Naeve [774]. Namboodiri [547]. Nancy [241, 342]. National [130, 566, 280, 290]. Nations [642]. Nature [529]. Navigation [760]. Neale [81]. necessary [511]. Negative [179, 586, 315]. neglected [361]. Neil [803]. Nested [462]. Neter [827]. Network [459]. Neuts [829]. Newell [187]. Newman [132].

Nicos [209]. Nodes [808]. Non

[619, 426, 141, 573, 2, 38, 817, 289, 747, 299, 140, 54, 586, 782, 878, 382, 783, 327, 787, 172, 336, 43, 485]. Non-Central [2, 747, 299, 782, 878, 382, 783]. non-dispersive [327]. Non-Hierarchical [619]. Non-Homogeneous [141]. Non-Integral [573, 336]. non-iterative [172]. Non-Linear [141, 817, 54]. Non-Mathematicians [43]. Non-Negative [586]. Non-Normality [140]. Non-Numerical [289]. Non-Orthogonal [38]. Non-parametric [426, 787]. Non-Statisticians [485]. Noncentral [883, 874, 875]. Nonlinear [183, 512, 401]. Nonparametric [773, 837]. Norbert [186]. Norm [781, 498, 590, 346]. Normal [712, 591, 229, 883, 874, 714, 869, 197, 667, 298, 520, 747, 856, 880, 217, 127, 357, 317, 574, 881, 586, 51, 94, 6, 7, 333, 593, 782, 878, 353, 806, 807, 852, 855, 382, 783, 865, 478, 471, 368, 370, 274, 473, 331, 795, 275, 472]. Normal-Johnson [869, 520, 856]. Normality [140, 602]. Norman [548, 25, 772, 584]. North [304]. Note [335, 350, 233, 764, 281, 155, 311, 181, 627, 124, 448, 450]. Novel [601, 825]. Novick [454]. Null [494, 383]. Number [224, 92, 587, 231, 404, 363]. Numbers [673, 30, 246, 155, 601, 825]. Numerator [620]. Numerical [289, 526].

O.K. [823]. Observation [224, 664, 452]. Observational [309]. Observations [696, 35, 332, 337, 660, 449, 71]. observed [786]. Observer [754]. Obtaining [197, 171]. Occurrence [304, 480]. Odds [621, 89]. Odeh [550]. Odell [132]. Odontology [759]. Office [617]. Official [617]. Olkin [374]. O’Loughlin [130]. O’Muircheataigh [680, 681]. One [307, 464, 688, 860, 710, 204, 657, 164, 744, 777, 444]. one-compartment [444].

One-Dimensional [744]. One-Hit [164]. One-Sided [464, 688, 860, 710]. One-Step [307]. open [444]. Operating [76, 52, 587]. Operation [337, 338, 25]. Operational [24, 770, 375, 546]. Operations [234, 157, 376]. operator [102]. Opinion [698, 402]. Optimal [232, 61, 86]. Optimization [183, 484, 744]. Optimum [335, 440, 350]. Ord [412]. Order [400, 714, 253, 633, 664, 881, 531, 442, 824]. Ordered [620]. Ordinal [620]. Organisation [378]. Orthogonal [5, 115, 117, 38, 59, 251, 415, 713, 866]. Orthogonalization [136]. Osborne [458]. Other [133, 107]. Outlier [722, 529, 816, 657]. Outliers [721, 694, 657, 756, 396]. Output [77, 290]. Owen [229, 704, 677, 263].

P [830, 493, 25, 672, 673, 773, 372, 45, 489, 674]. P. [762, 132, 484, 641, 546, 44]. Painless [673]. pair [363]. Paired [233, 603, 792, 599, 73, 226, 364]. Pairs [793]. Pan [851, 871]. Paper [367, 440]. Papers [811, 262]. Parameter [777, 573, 71, 358, 444]. Parameters [496, 37, 816, 563, 6, 333, 315, 471, 274, 479, 401, 336]. parametric [426, 787]. Part [240, 330]. Partially [720]. Particular [635]. Partitioning [253]. Party [433, 78]. Patel [677]. Patients [528, 614, 458]. Pattern [458, 637]. Patterns [780]. Patterson [684, 775]. Paul [454, 84]. Payne [680, 681]. Pearson [458]. Pentikainen [683]. Percentage [591, 465, 317, 725, 791]. Percentages [673]. Permutational [242]. Permutations [425]. Personal [238, 287]. Pesonen [683]. Peter [64, 774, 237, 263]. Peters [42]. Pfaffenberger [684, 775]. Phillips [409]. Physical [772]. Physicists [213]. piecewise [479]. Pielou [82, 549]. Planning [145, 642, 109]. Plans [587, 55, 329]. Plant [17]. Players [80]. Plots [268, 873, 741, 868, 652, 859, 10, 95, 255]. Plotting [735, 400, 217, 51, 134, 135, 94, 270, 478]. Pneumoconiosis [429]. Pocket [768, 574]. Point [479, 631, 480, 787]. Point-Process [480]. Points [591, 465, 317, 725, 791]. Poisson [152, 755, 818, 225, 125]. Police [240]. Policy [708]. Pollard [284]. Pollution [514]. Polly [191]. Polychoric [592, 437]. Polygonal [793]. Polynomial [256, 356]. Polynomials [5, 115, 117, 495]. population [471, 472]. Populations [197, 88, 138, 218, 622, 353, 254, 275]. Position [367, 760, 440]. positions [478]. Positive [179, 867, 4, 791, 450]. Possible [598]. Posterior [526, 570]. Potential [811]. Powell [287]. Power [701, 715, 858, 571, 757, 508, 198, 194, 850]. Practical [801, 209, 43, 647, 365]. Practice [379, 542]. Prais [239]. Pratt [514]. Pre [149, 377]. Pre-1968 [377]. Pre-Computer [149]. Precipitation [568]. Precision [100]. Predictability [277]. Predicting [307]. Prediction [576, 507]. Predictive [303]. Preference [506]. Preferences [603, 364]. Preparation [29]. Present [657]. Preventive [675]. Price [11]. Prices [277]. Primer [801]. Principal [175, 223, 567, 527, 552, 273]. Principles [159, 157, 44]. Printing [243]. Probabilistic [232]. Probabilities [318, 463, 883, 874, 851, 871, 245, 88, 138, 218, 622, 747, 314, 630, 629, 782, 878, 382, 783, 362]. Probability [735, 712, 400, 235, 374, 676, 180, 79, 829, 607, 568, 614, 285, 160, 865, 255, 478, 83, 704, 457]. Probability-TIMS [829]. Probit [185]. Problem [743, 778, 697, 572, 516, 786, 313, 72, 787]. Problem-Solving [516]. Problems [387, 98, 306, 558, 381, 870, 131, 731, 365, 514]. Procedure [499, 348, 724, 851, 871, 101, 749, 326, 137, 349, 815, 312]. Procedures [58, 529, 188, 558, 67, 381, 72]. Process [259, 728, 758, 480, 579]. Processing [36]. Prodan [66]. Product [603, 140, 791, 364]. Product-Moment [140]. Production [263, 123]. productivity [102]. Professional [84]. Prognosis [376]. Program [12, 776, 38, 156, 433, 315, 451, 123, 252]. Programmes [408]. Programming [670, 61, 208, 241, 342, 227, 77, 63, 183, 548]. Programs [65, 737]. Progressively [357]. Promotion [663]. Properties [476, 60]. Proportional [625, 788]. Proportions [659]. Prospectuses [663]. Pseudo [518, 559, 745, 231, 404, 473]. Pseudo-Random [518, 559, 745, 231, 404, 473]. Psi [556]. Psychology [545, 27]. Public [708]. Purchasing [57]. Purpose [721].

Quadratic [586]. Quality [40, 640, 188, 460]. Quantifying [429]. Quantitative [208, 250, 613, 674]. Quarterly [281, 280]. Quasi [101]. Quasi-Independence [101]. Quenouille [703]. Questionnaires [147]. Questions [698]. Queueing [187, 709, 313]. Quick [571]. quotient [450].

R [64, 801, 803, 213, 371, 373, 25, 576, 682, 484, 836, 26, 454, 516, 341]. R. [580, 493, 87, 373, 576, 48, 545, 26, 641, 208, 410, 707, 20, 108, 287, 340, 550, 107, 683]. R10 [347]. R11 [348]. R12 [386]. R13 [417]. R14 [498]. R15 [499]. R16 [522]. R17 [558]. R18 [559]. R19 [589]. R20 [592]. R21 [593]. R22 [622]. R23 [652]. R24 [745]. R25 [746]. R26 [747]. R27 [748]. R28 [749]. R29 [781]. R3 [93]. R30 [844, 782]. R31 [852]. R32 [855]. R33 [856]. R34 [857]. R35 [858]. R36 [860]. R37 [861]. R38 [862]. R39 [863]. R4 [117]. R40 [864]. R41 [865]. R42 [866]. R44 [867]. R46 [868]. R48 [869]. R49 [870]. R5 [217]. R52 [871]. R53 [872]. R54 [873]. R55 [874]. R56 [875]. R6 [218]. R68 [879].

R7 [246]. R70 [880]. R72 [881]. R79 [882]. R8 [299]. R80 [883]. R81 [884]. R82 [885]. R83 [886, 887]. R9 [319]. Rade [160]. Radioimmunoassay [635]. Radiological [429]. Rainfall [769]. Ralston [671]. Random [761, 762, 132, 755, 777, 841, 854, 871, 872, 518, 559, 713, 745, 793, 231, 739, 866, 818, 791, 404, 473, 336]. Range [712, 831, 865]. Rank [474, 432]. Rank-Correlation [474]. Ranking [30, 246, 204]. Rankits [650]. Ranks [121]. Rates [754, 663, 398, 276, 355]. Ratio [621, 886, 589, 296, 89, 887]. Ratios [673, 658]. Rawstron [184]. Ray [485]. Re [663]. Re-Interpreting [663]. Reading [513]. Readings [42]. Real [223, 517]. Realistic [492]. Rebecca [799]. Recidivism [634]. Reconciliation [240]. Recorded [566]. Records [18]. Rectangular [60]. Recurrence [152]. Recursive [558]. Redpath [287]. Reduction [543]. Reduction-Analysis [543]. Reference [635, 289, 263]. Refinement [176]. Regarding [222]. Regional [184]. Regions [753, 506]. Registration [224]. Regression [743, 746, 778, 839, 335, 307, 864, 153, 729, 59, 385, 879, 173, 817, 870, 606, 178, 599, 625, 809, 227, 170, 567, 588, 110, 256, 664, 394, 100, 350, 356, 479, 527, 392, 401, 797]. Regressions [156, 273]. Regressor [303]. Regressors [153]. Reinke [379]. Rejection [529]. Relation [152]. Relations [537, 260, 393]. Relationship [35, 572]. relationships [786]. Reliability [544, 568, 203, 278, 422]. Remark [746, 748, 592, 886, 860, 883, 864, 858, 348, 874, 875, 117, 589, 882, 873, 877, 884, 869, 863, 386, 558, 871, 872, 879, 857, 246, 867, 861, 862, 870, 868, 218, 622, 747, 749, 856, 880, 299, 745, 93, 319, 652, 217, 885, 876, 881, 417, 498, 593, 866, 782, 878, 522, 852, 855, 865, 887, 844, 746, 748, 592, 499, 860, 864, 858, 348, 874, 875, 117, 589, 882, 873, 884, 869, 863, 386, 558, 871, 872, 879, 857, 246, 867, 861, 862, 559, 870, 868, 218, 622, 747, 749, 856, 880, 299, 745, 93, 319]. Remark [652, 217, 885, 781, 347, 881, 417, 498, 593, 866, 782, 522, 852, 855, 865]. Remarks [321, 656, 781]. Remington [87]. Repeat [506, 265]. Repeat-Buying [265]. Replacement [405, 629]. Replicated [35, 792]. replication [200]. Report [575, 655, 666, 310, 128, 733, 369, 78]. Representation [727]. Research [234, 43, 813, 770, 375, 157, 379, 376, 546, 240, 263, 129, 28, 24, 646, 182]. Residual [280]. Residuals [557, 558, 817]. respiration [509]. Response [198, 181, 226]. Results [337, 365, 355]. Review [41, 761]. Reviews [47, 830, 292, 551, 761, 762, 132, 458, 234, 580, 827, 487, 111, 63, 211, 183, 64, 801, 40, 513, 640, 803, 799, 579, 548, 800, 129, 209, 43, 83, 213, 235, 371, 493, 835, 833, 612, 87, 373, 184, 796, 374, 679, 705, 130, 774, 776, 412, 133, 25, 703, 639, 28, 491, 576, 672, 673, 767, 188, 24, 682, 706, 644, 48, 708, 484, 545, 67, 27, 49, 106, 212, 407, 408, 647, 676, 670, 671, 770, 771, 802, 773, 772, 836, 26, 81, 375, 680, 681, 456, 542, 42, 65, 286, 207, 289]. Reviews [454, 582, 737, 828, 768, 191, 643, 372, 455, 837, 797, 704, 84, 646, 641, 547, 62, 82, 208, 159, 241, 342, 514, 410, 409, 616, 609, 707, 549, 829, 769, 544, 577, 185, 66, 86, 189, 678, 85, 187, 157, 379, 378, 376, 20, 23, 546, 190, 583, 578, 488, 642, 131, 614, 832, 109, 110, 239, 238, 290, 411, 677, 615, 46, 457, 541, 490, 285, 645, 108, 44, 584, 287, 738, 41, 380, 831, 516, 486, 539, 158, 291, 210, 160, 182, 265, 340, 543, 552, 610, 611, 739, 341, 550, 804, 540, 709, 262]. Reviews [264, 288, 492, 515, 740, 675, 459, 485, 834, 107, 237, 377, 186, 240, 581, 45, 105, 263, 683, 553, 489, 617, 613, 685, 674, 684, 775, 21, 22, 284, 460, 798, 266, 293, 236]. Revised [263]. Revisited [151]. Reyment [580]. Rho [463]. Richard [672, 673, 773]. Riedel [614]. Riposte [449]. Risk [683]. Road [224]. Roads [15]. Robert [799, 612, 67, 676, 158, 540, 492]. Robust [718]. Robustness [261, 421]. Roessler [285]. Roger [380, 684, 775]. Rohlf [373]. Role [656]. Romano [804]. Romeder [408]. Rooms [528]. Root [151, 383]. Roots [267, 320, 258]. Rothamsted [252]. Routines [156]. Row [841, 397]. Row-and-column [397]. Rowe [644]. Roy [111]. Rules [823]. Runs [50, 522]. Runyon [672, 673, 773]. Russian [236]. Russian-English [236]. Russian-English/English-Russian [236]. Ruth [376]. RV [532]. RV- [532].

S [458, 835, 491, 24, 671, 42, 62, 616, 110, 239, 265, 543, 540]. S. [551, 800, 771, 26, 643, 239, 287]. Sabine [66]. Salaries [292, 266, 293]. Sales [835]. Sample [405, 466, 114, 534, 306, 52, 61, 616, 816, 140, 69, 710, 99, 550, 329]. Samples [177, 357, 602, 780, 171, 331]. Sampling [58, 40, 640, 269, 52, 609, 587, 314, 614, 13, 55, 738, 340, 629, 328, 329, 471, 275]. Samuel [209, 62, 236]. Satellite [103]. Scale [458, 496, 38, 816, 347, 10, 95]. Scales [764]. Scaling [497]. Scatter [873, 741, 868, 347, 134, 859]. Scheer [644]. Scheme [328]. Schemes [201]. Schmidt [136, 385]. Schork [87]. Schucany [341]. Science [644, 159, 804, 288, 235]. Sciences [487, 833, 772, 829, 799, 87, 67, 643, 544, 490, 645, 515, 798]. Scientific [768, 609, 182, 264, 291]. Scientists [488, 540]. Scores [202]. scoring [401]. Scott [831]. Screening [477]. Search [91]. Seasonal [604, 325]. Seasonality [535]. Second [253, 291]. Second-Order [253]. Security [16]. Segmentation [410]. SELECT [156]. Selected [67]. Selection [59, 636, 482, 668, 567, 723, 627, 10, 95, 22, 585, 353]. Selections [347]. Self [513]. Self-Learning [513]. Sellekaerts [553]. Seltzer [455]. Semi [867, 4]. Semi-Definite [867, 4]. Sequence [162, 163]. Sequences [476, 769]. Sequential [52, 587, 70, 314, 583, 13, 756, 629]. Series [830, 700, 118, 97, 542, 65, 170, 853, 849, 431, 608, 510, 611, 531, 452, 505, 401, 456, 291]. Service [284, 313]. Services [378]. Sets [884, 742]. Several [197, 602, 260, 393]. Severity [34]. Shape [573, 816, 120, 336]. Shapiro [62]. Sharpe [241, 342]. Sheffield [655]. Shell [56]. Shift [631]. Shooting [202]. Short [630]. Short-Term [630]. Sickness [282, 325]. Sided [621, 464, 688, 860, 710, 275]. Sidney [513, 27]. Sign [435]. Significance [839, 536, 153, 582, 423, 242, 225, 366]. Silvey [551]. Simple [743, 405, 637, 822, 734, 232, 101, 429, 853, 849, 497, 427]. Simplex [499, 348, 749, 137, 349]. Simplified [328]. Simulating [462, 780]. Simulation [790, 126, 447, 200, 824, 533, 355]. Simulations [765]. Simultaneous [101, 506, 815]. Single [555, 816, 102]. singular [527]. Six [268, 652]. Six-Line [268, 652]. Size [466, 534, 52, 550, 329]. Sizes [259, 69, 710]. Skeletal [576]. Skewness [434]. Slater [776]. Slope [14]. Small [202, 39, 142, 177, 626, 527, 331]. Small-Bore [202]. Smith [768]. Smoothing [345, 861]. Snedecor [262]. Social [487, 129, 184, 142, 644, 643, 642, 740, 240, 798]. Sociological [13]. Sociology [111]. Software [644, 278]. Sokal [373]. Solving [131, 516]. Som [340]. Some [330, 822, 656, 72, 337, 606, 442, 406, 60, 99, 71, 444]. Somesh [372]. Source [458]. Sources [679, 190, 552, 366]. Space [316]. Space-Time [316]. Spanning [104, 113]. Sparse [753, 558, 381, 651]. Spatial [637, 412, 780]. Spearman [463, 820]. Special [292, 289, 263, 266, 293]. Specification [430]. Spectral [752, 518, 559, 327, 745, 447]. Spectrum [345, 861]. Speed [15]. Spherical [170, 413, 626]. Spurious [696]. Spurr [489].

Square [307, 151, 303, 507]. Squared

[725, 786]. Squares

[558, 178, 757, 206, 19, 359, 172]. Squeeze [536]. Stabilizing [125]. Stable [403]. stage [275]. Stahl [513]. Stand [753]. Standard [648, 794]. Stark [238]. stations [199]. Statistic [153, 39, 269, 494, 734, 851, 871, 216, 602, 394, 383, 786, 341]. Statistical [565, 268, 8, 387, 468, 500, 560, 594, 845, 843, 844, 743, 746, 778, 777, 618, 690, 555, 619, 621, 712, 748, 839, 592, 591, 840, 496, 499, 556, 886, 318, 463, 465, 68, 640, 464, 688, 860, 841, 883, 689, 466, 648, 435, 864, 715, 858, 348, 222, 874, 875, 114, 136, 9, 53, 115, 117, 589, 882, 205, 854, 714, 873, 877, 884, 269, 494, 869, 50, 28, 188, 863, 706, 385, 386, 557, 558, 851, 871, 872, 879, 857, 30, 245, 246, 842, 867, 345, 861, 862, 27, 90, 91, 381, 462]. Statistical [741, 76, 518, 559, 75, 212, 384, 415, 467, 870, 676, 671, 836, 193, 243, 521, 868, 443, 391, 779, 344, 713, 52, 128, 88, 92, 138, 218, 298, 334, 519, 520, 622, 747, 749, 856, 880, 372, 299, 745, 62, 514, 93, 409, 616, 794, 707, 809, 319, 758, 18, 32, 711, 849, 620, 651, 742, 69, 587, 652, 86, 428, 295, 296, 413, 414, 434, 436, 437, 70, 495, 314, 416, 517, 649, 217, 554, 588, 710, 77, 74, 433, 497, 137, 349, 216, 317, 808, 677, 215, 885, 242, 162, 163, 316, 781, 44, 526, 876]. Statistical [780, 532, 113, 297, 347, 686, 346, 687, 41, 881, 586, 214, 120, 51, 134, 135, 94, 244, 267, 270, 320, 417, 543, 194, 498, 590, 540, 859, 264, 7, 593, 866, 31, 288, 515, 10, 95, 89, 164, 782, 878, 485, 834, 237, 383, 186, 45, 263, 489, 617, 684, 775, 522, 22, 585, 650, 798, 236, 806, 807, 852, 855, 315, 116, 139, 382, 783, 744, 865, 78, 814, 444, 292, 551, 705, 262, 266, 293]. Statisticians [454, 85, 485]. Statistics [47, 321, 387, 468, 400, 513, 803, 129, 43, 83, 213, 235, 87, 796, 656, 714, 639, 491, 672, 673, 682, 48, 708, 655, 632, 545, 67, 49, 741, 647, 773, 772, 81, 704, 641, 474, 413, 414, 583, 131, 615, 46, 285, 516, 881, 804, 733, 740, 240, 284, 442, 824, 398, 23, 827, 371, 493, 679, 133, 42, 286, 837, 577, 20, 190, 108, 380, 210, 160, 617, 460]. Step [307]. Stephen [516]. Stimson [376]. Stochastic [603, 769, 181, 406, 303, 480, 683, 199, 364]. Stock [277]. Stone [107]. Stop [697]. Straight [35, 781, 346, 498, 590]. Strata [61]. Strategies [487, 22]. Strategy [232]. Stratification [257, 449, 448]. stratified [471]. Strike [730]. Strong [710]. Structural [661, 409]. Structure [885, 297, 422]. Structured [884, 742]. Structures [75, 32, 680]. Stuart [609, 488]. Student [427, 31]. Students [545, 740]. Studies [184, 309, 36, 829, 120, 275, 789, 639, 109, 46]. Study [811, 721, 126, 821, 566, 447]. Styan [372]. Subject [720]. Subjective [529, 568, 630]. Subroutine [339]. subroutines [258]. Subset [91, 156, 726, 72]. substitution [123]. Successive [58]. sufficient [511]. Sum [746, 879, 588]. Summary [666, 455, 733, 369]. Sums [55, 194]. Supplement [65]. supplementary [448]. Supplies [598]. Supply [16]. Surface [226]. Surfaces [198]. Survey [813, 148, 337, 680, 681, 144, 326, 616, 149, 566, 147, 99, 252, 834]. Surveys [145, 13, 287, 171, 252, 84]. Survival [862, 37, 625, 711, 544, 230]. Susan [262]. Swedish [292, 266, 293]. Swenson [458]. Swindle [279]. swindles [533]. Symbolic [283]. Symmetric [748, 864, 386, 867, 90, 4, 3, 809, 267, 320, 403]. Symposia [103]. Symposium [207]. Synergism [538]. System [76, 18, 16, 203, 530]. Systemic [819]. Systems [211, 528, 819, 291, 709, 459, 327, 422, 802]. Szmidt [705].

T [209, 771, 645, 584, 377]. T. [132, 705, 737, 372, 769, 108, 262, 683].

Table [621, 89, 815]. Tables [761, 618, 690, 592, 841, 39, 877, 857, 101, 193, 243, 521, 253, 85, 437, 739, 820, 798, 403, 366, 477, 395, 359, 762, 454, 800]. Tabulated [832]. Tabulating [884, 651, 742]. Tail [318, 463, 851, 871]. Tailed [760, 474]. Tanis [676]. Tanner [576]. Tau [318]. Taylor [401]. Teach [24]. Teaching [705, 160]. Technique [567, 13]. Techniques [183, 28, 289, 643, 227, 99, 340, 44, 738, 459]. Technological [28]. Technology [796]. Telephones [302]. Teletypewriter [255]. Temperature [568, 199]. temperatures [509]. tended [102]. Tennis [232]. Term [630, 358]. Termination [730]. Terms [663, 600, 236]. Territories [601, 825]. Terry [226]. Test [426, 435, 701, 715, 858, 538, 50, 518, 559, 101, 745, 421, 316, 756, 504, 225, 522, 815, 850, 396, 276]. Testing [153, 764, 603, 506, 423, 242, 602, 260, 393, 364]. Tests [839, 637, 536, 148, 631, 757, 816, 535, 390, 765, 657, 164, 442, 794, 582]. Tetrachoric [648]. Text [516]. Textile [728]. Thad [682]. Them [260, 393]. Theoretical [431, 608]. Theories [487]. Theory [374, 542, 616, 584, 265, 683, 553, 189, 678, 187]. Theory-Essays [553]. Thinking [409]. Thomas [802, 614, 238, 540]. Thought [207]. Three [253]. Three-Way [253]. Thunderstorm [179]. Ties [204]. Time [830, 700, 118, 97, 764, 542, 631, 431, 608, 316, 510, 531, 695, 786, 452, 505, 327, 611, 456]. time-dependent [695]. Time-Series [118, 456]. Times [697, 534, 761, 762, 739]. TIMS [829]. Tinbergen [553]. Today [337]. Tolerance [466, 710, 275]. Tool [532]. Totals [841]. Town [145]. Town-Planning [145]. toxicological [789]. Traffic [98, 224, 311, 15]. Transfer [619]. Transform [345, 861, 416, 517, 649]. Transformation [453, 571, 151, 231, 125, 404]. Transformations [604, 869, 520, 856, 198, 354]. Transient [752]. Transition [79, 362]. Transmission [16]. Transposing [32]. Trapezium [823]. Traversed [228]. Treatment [305, 377, 789]. Tree [113]. Trees [104]. tri [370]. tri-variate [370]. Triangular [748, 386, 867, 3]. Trigamma [687]. Trimming [718, 719]. Truncated [152, 127, 634, 315, 274]. Trunk [15]. Tukey [767, 797]. TW2 [576]. Two [565, 699, 58, 618, 690, 621, 332, 701, 715, 858, 538, 877, 306, 857, 636, 482, 668, 80, 474, 535, 723, 710, 570, 686, 275, 225, 164, 366, 850, 395, 200, 358, 814]. Two-Factor [332, 200]. Two-Group [723]. Two-Hit [164]. Two-Level [636, 482, 668]. two-parameter [358]. Two-Sample [306]. Two-Sided [621, 710, 275]. two-stage [275]. Two-Tailed [474]. two-term [358]. Two-Way [618, 690, 877, 570, 686, 857, 366, 395]. Typescript [29].

U [287, 375]. U-Boat [375]. U.K. [280, 662]. Ullman [803]. Ultra [661]. Ultra-Structural [661]. Ultrametric [555]. Unbalanced [599, 570, 686, 451]. Unbiased [152]. Uncertainties [178]. Uncertainty [457, 162]. Unconfirmed [530]. Unconstrained [879]. Underestimation [14]. Understanding [513]. Unequal [332, 629]. Uniform [56, 793]. Unifying [532]. unit [362]. United [238, 290, 642]. Units [174]. University [411]. Unknown [631, 816]. Updating [114, 113, 530]. Upper [318, 463]. Upton [832]. Ursula [685]. Use [179, 153, 5, 115, 117, 59, 528, 257, 144, 85, 574, 54, 504, 308, 448, 795]. Useful [883, 874, 747, 782, 878, 382, 783]. Uses [577]. Using [499, 307, 754, 749, 61, 241, 231, 137, 602, 226, 459, 530, 634, 348, 404, 442, 342, 349].

V [800, 676, 207]. V. [541, 515]. Value [743, 527]. Values [435, 115, 424, 481, 756, 396, 312, 172]. Variability [821]. Variable [259, 464, 306, 92, 758, 567, 116, 139]. Variables [755, 777, 573, 309, 59, 854, 884, 871, 872, 569, 175, 223, 742, 723, 627, 303, 586, 225, 786, 818, 791, 336]. Variance [152, 141, 38, 635, 344, 631, 628, 510, 125, 531, 283, 451, 363, 172]. Variance-Stabilizing [125]. Variate [229, 822, 214, 125, 370]. Variates [132]. Variations [184, 282, 325, 199]. Vectors [344, 267, 320, 258]. Venkatesan [182]. Verbal [528]. Verbal-Call [528]. Vetting [147]. via [345, 861]. Virgil [492]. Virginia [544]. vital [398]. Vol [234, 680, 681]. Volume [33, 96, 168, 249, 324, 420, 503, 597, 693, 848, 829, 709]. vs [164]. Vulcanism [626].

W [830, 235, 833, 133, 767, 372, 837, 797, 514, 189, 490, 108, 262, 264, 237, 798]. W. [640, 28, 576, 770, 42, 65, 410, 379, 738, 341, 834, 674]. Waddington [375]. Wagner [159, 157]. Waiting [528]. Wald [14]. Wald-Bartlett [14]. Walpole [20]. Walter [109, 377]. War [375]. Wards [449]. Warwick [369]. Wasserman [827]. Water [598]. Waterfowl [99]. Watson [209, 851, 871]. Watson-Gandy [209]. Way [618, 690, 877, 253, 570, 686, 366, 857, 395]. Wayne [833, 837, 490]. weather [199]. Weight [638, 763]. Weighted [558, 381]. Weights [688, 860, 537]. Wendell [458]. Wertz [834]. Wetherill [40, 583, 41]. Wheeler [541]. which [313]. White [63, 579]. Whitehouse [576, 459]. Whitmore [827]. Whitney [269, 216]. Wilf [671]. William [827, 235, 708, 455, 241, 342, 489]. Williamson [24]. Willy [553]. Wind [410]. Winding [728]. Winds [304]. Winning [673]. Winsorized [124]. Winter [34]. Winters [724]. Wishart [214]. Wit [563].

Without [405, 200, 629]. Wood [612]. Wording [698]. Work [737, 78]. Workbook [188]. Working [433, 78]. World [375].

Y. [410, 798]. Yannoulis [798]. Yield [17]. Yourself [24].

Z. [798]. Zonal [495].

References

Cooper:1968:AAA

[1] B. E. Cooper. Algorithms: Algorithm AS 4: An auxiliary function for distribution integrals. Applied Statistics, 17(2): 190–192, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/4. See remarks [299].

Cooper:1968:AAIb

[2] B. E. Cooper. Algorithms: Algorithm AS 5: The integral of the non-central t-distribution. Applied Statistics, 17 (2):193–194, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/5. See remarks

[1, 299, 875].

Healy:1968:AAT

[3] M. J. R. Healy. Algorithms: Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 17 (2):195–197, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/6. See corrigen-

dum [8] and remark [386, 867, 748].

Healy:1968:AAI

[4] M. J. R. Healy. Algorithms: Algorithm AS 7: Inversion of a positive semi-definite symmetric matrix. Applied Statistics, 17(2):198–199, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/7.

See corrigendum [8] and remark [867].

Cooper:1968:AAU

[5] B. E. Cooper. Algorithms: Algorithm AS 10: The use of orthogonal polynomials. Applied Statistics, 17(3):283–287, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/10. See remarks [93, 117].

Swan:1969:CML

[6] A. V. Swan. Computing maximumlikelihood estimates for parameters of the normal distribution from grouped and censored data. Applied Statistics, 18(1):65–69, 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [333].

Swan:1969:SAAa

[7] A. V. Swan. Statistical algorithms: Algorithm AS 16: Maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 18(1): 110–114, March 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/16. See remarks

[880, 593].

Anonymous:1969:SAA

[8] Anonymous. Statistical algorithms: Algorithms AS 6 and AS 7: Corrigenda.

Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [3, 4, 386, 748].

Cooper:1969:SAAb

[9] B. E. Cooper. Statistical algorithms: Algorithm AS 4: An auxiliary function for distribution integrals: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/4. See cor-

rigendum [53].

Thayer:1969:SAA

[10] R. P. Thayer and R. F. Storer. Statistical algorithms: Algorithm AS 21: Scale selection for computer plots. Applied Statistics, 18(2):206–208, June 1969. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/21. See corrigenda [95] and remark [347].

Tipping:1970:PCI

[11] D. G. Tipping. Price changes and income distribution. Applied Statistics, 19(1):1–17, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1970:MAC

[12] A. J. B. Anderson and Bridget I. Lowe. A multivariate analysis computer program. Applied Statistics, 19(1):18–26, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peel:1970:SST

[13] J. Peel and G. E. Skipworth. Sequential sampling as a technique in sociological surveys. Applied Statistics, 19(1):27–33, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carter:1970:UEW

[14] L. F. Carter and H. M. Blalock. Underestimation of error in Wald-Bartlett slope estimation. Applied Statistics, 19 (1):34–41, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wabe:1970:CST

[15] J. S. Wabe. Congestion and the speed of traffic on trunk roads. Applied Statistics, 19(1):42–49, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyness:1970:SSG

[16] F. K. Lyness. Security of supply in a gastransmission system. Applied Statistics, 19(1):50–63, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mead:1970:PDC

[17] R. Mead. Plant density and crop yield. Applied Statistics, 19(1):64–81, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kronmal:1970:MCS

[18] R. A. Kronmal, Linda Bender, and J. Mortensen. Miscellanea: A conversational statistical system for medical records. Applied Statistics, 19(1):82–92, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oliver:1970:MEE

[19] F. R. Oliver. Miscellanea: Estimating the exponential growth function by direct least squares. Applied Statistics, 19 (1):92–100, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Machin:1970:BRI

[20] D. Machin. Book reviews: Introduction to Statistics, by R. E. Walpole. Applied Statistics, 19(1):101, 1970. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Welford:1970:BRM

[21] B. P. Welford. Book reviews: Mathematical Model Building in Economics and Industry. Applied Statistics, 19(1):102, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Welford:1970:BRS

[22] B. P. Welford. Book reviews: Statistical Selection of Business Strategies, by J. Forester. Applied Statistics, 19(1):103, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Machin:1970:BRS

[23] D. Machin. Book reviews: Statistics for Mathematicians — An Introduction, by D. J. Finney. Applied Statistics, 19(1):104, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1970:BRT

[24] I. G. Evans. Book reviews: Teach Yourself Operational Research, by M. S. Makower and E. Williamson. Applied Statistics, 19(1):104–105, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1970:BRE

[25] Paul Davies. Book reviews: Evolutionary Operation, by George E. P. Box and Norman R. Draper. Applied Statistics, 19(1):105–106, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1970:BRE

[26] J. D. Griffiths. Book reviews: Exercises in Industrial Management, by S. Eilon, R. I. Hall and J. R. King. Applied Statistics, 19(1):106, 1970. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gales:1970:BRI

[27] Kathleen Gales. Book reviews: Introduction to Statistical Analysis and Inference for Psychology and Education, by Sidney J. Armore. Applied Statistics, 19(1):107, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Draper:1970:BRS

[28] Norman R. Draper. Book reviews: Statistical Techniques in Technological Research, by W. E. Duckworth. Applied Statistics, 19(1):107–109, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1970:PTA

[29] Anonymous. Preparation of typescript for algorithm descriptions. Applied Statistics, 19(1):110, 1970. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1970:SAA

[30] P. R. Freeman. Statistical algorithms: Algorithm AS 26: Ranking an array of numbers. Applied Statistics, 19(1):111– 113, March 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/26. See remark [246].

Taylor:1970:SAA

[31] G. A. R. Taylor. Statistical algorithms: Algorithm AS 27: The integral of Student’s t-distribution. Applied Statistics, 19(1):113–114, March 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/27.

Krzanowski:1970:SAA

[32] W. J. Krzanowski. Statistical algorithms: Algorithm AS 28: Transposing multiway structures. Applied Statistics, 19(1):115–118, March 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/28.

Anonymous:1970:VIP

[33] Anonymous. Volume information. Applied Statistics, 19(1):[3], 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyness:1970:MWS

[34] F. K. Lyness and E. H. M. Badger. A measure of winter severity. Applied Statistics, 19(2):119–134, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1970:FSL

[35] V. D. Barnett. Fitting straight lines-the linear functional relationship with replicated observations. Applied Statistics, 19(2):135–144, 1970. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldstein:1970:DPL

[36] H. Goldstein. Data processing for longitudinal studies. Applied Statistics, 19 (2):145–151, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garg:1970:MLE

[37] Mohan L. Garg, B. Raja Rao, and Carol K. Redmond. Maximumlikelihood estimation of the parameters of the Gompertz survival function. Applied Statistics, 19(2):152–159, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunningham:1970:PLS

[38] E. P. Cunningham. A program for largescale non-orthogonal analysis of variance and covariance. Applied Statistics, 19 (2):160–172, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Craddock:1970:MDS

[39] J. M. Craddock and C. R. Flood. Miscellanea: The distribution of the χ^{2} statistic in small contingency tables. Applied Statistics, 19(2):173–181, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1970:BRS

[40] A. F. Bissell. Book reviews: Sampling Inspection and Quality Control, by G. Barrie Wetherill. Applied Statistics, 19(2):182, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1970:BRR

[41] J. F. Scott. Book reviews: Review of Elementary Statistical Methods, by G. Barrie Wetherill. Applied Statistics, 19(2):183, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Henshaw:1970:BRR

[42] W. C. Henshaw. Book reviews: Readings in Applied Statistics, by W. S. Peters. Applied Statistics, 19(2):183–184, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chapas:1970:BRB

[43] L. C. Chapas. Book reviews: Biological Research Method: Practical Statistics for Non-Mathematicians, by H. H. Holman. Applied Statistics, 19(2):184–185, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quenouille:1970:BRP

[44] M. H. Quenouille. Book reviews: Principles of Statistical Techniques, by P. G. Moore. Applied Statistics, 19(2):185, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1970:BRI

[45] Denis H. Ward. Book reviews: Introduction to Applied Statistical Analysis, by H. P. Hartkemeier. Applied Statistics, 19(2):186, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1970:BRA

[46] D. J. Page. Book reviews: Applied Statistics for Management Studies, by D. Croft. Applied Statistics, 19(2):

186–187, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Allum:1970:BRS

[47] W. J. Allum. Book reviews: Statistics in the Computer Age, by J. M. Craddock. Applied Statistics, 19(2):187–188, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ferres:1970:BRS

[48] H. M. Ferres. Book reviews: Statistics for Biologists, by R. C. Campbell. Applied Statistics, 19(2):188, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garreau:1970:BRS

[49] G. A. Garreau. Book reviews: Statistics for Experimentalists, by B. E. Cooper. Applied Statistics, 19(2):189, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Downham:1970:SAA

[50] D. Y. Downham. Statistical algorithms: Algorithm AS 29: The runs up and down test. Applied Statistics, 19(2): 190–192, June 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/29. See [522].

Sparks:1970:SAA

[51] D. N. Sparks. Statistical algorithms: Algorithm AS 30: Half normal plotting. Applied Statistics, 19(2):192– 196, June 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/30. See corrigenda [94, 217].

Hill:1970:SAA

[52] I. D. Hill. Statistical algorithms: Algorithm AS 31: Operating characteristic and average sample size for binomial sequential sampling. Applied Statistics, 19 (2):197–203, June 1970. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/31.

Cooper:1970:SAC

[53] B. E. Cooper. Statistical algorithms: Corrigendum: Algorithm AS 4: An auxiliary function for distribution integrals. Applied Statistics, 19(2):204, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/4. See [9].

Ross:1970:EUF

[54] G. J. S. Ross. The efficient use of function minimization in non-linear maximum-likelihood estimation. Applied Statistics, 19(3):205–221, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prairie:1970:CSP

[55] Richard R. Prairie and William J. Zimmer. Continuous sampling plans based on cumulative sums. Applied Statistics, 19(3):222–230, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Doehlert:1970:USD

[56] David H. Doehlert. Uniform shell designs. Applied Statistics, 19(3):231–239, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1970:BBM

[57] C. Chatfield and G. J. Goodhardt. The beta-binomial model for consumer purchasing behaviour. Applied Statistics, 19 (3):240–250, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Avadhani:1970:CTS

[58] M. S. Avadhani and B. V. Sukhatme. A comparison of two sampling procedures with an application to successive sampling. Applied Statistics, 19(3):251–259, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daling:1970:UOF

[59] Janet R. Daling and H. Tamura. Use of orthogonal factors for selection of variables in a regression equation — an illustration. Applied Statistics, 19 (3):260–268, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oliver:1970:MSA

[60] F. R. Oliver. Miscellanea: Some asymptotic properties of Colquhoun’s estimators for a rectangular hyperbola. Applied Statistics, 19(3):269–273, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huddleston:1970:MOS

[61] H. F. Huddleston, P. L. Claypool, and R. R. Hocking. Miscellanea: Optimal sample allocation to strata using convex programming. Applied Statistics, 19 (3):273–278, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1970:BRS

[62] J. N. R. Jeffers. Book reviews: Statistical Methods in Engineering, by Gerald J. Hahn and Samuel S. Shapiro. Applied Statistics, 19(3):279, 1970. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beale:1970:BRD

[63] E. M. L. Beale. Book reviews: Dynamic Programming, by D. J. White. Applied Statistics, 19(3):280–281, 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Benjamin:1970:BRD

[64] Bernard Benjamin. Book reviews: Demography. 4th Edition, by Peter R. Cox. Applied Statistics, 19(3):281–282, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Herne:1970:BRB

[65] H. Herne. Book reviews: Biomedical Computer Programs. X-Series Supplement, by W. J. Dixon. Applied Statistics, 19(3):282–283, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1970:BRF

[66] D. K. Lindley. Book reviews: Forest Biometrics, by Michail Prodan and Sabine H. Gardiner. Applied Statistics, 19(3):283, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1970:BRI

[67] P. R. Freeman. Book reviews: Introduction to Statistics: Selected Procedures for the Behavioural Sciences, by Robert Fried. Applied Statistics, 19(3):284, 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bhattacharjee:1970:SAA

[68] G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 32: The incomplete gamma integral. Applied Statistics, 19(3):285–287, September 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/32. See re-

mark [876, 882].

Leech:1970:SAA

[69] F. B. Leech. Statistical algorithms: Algorithm AS 33: Calculation of hypergeometric sample sizes. Applied Statistics, 19(3):287–289, September 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/33.

McGilchrist:1970:SAA

[70] C. A. McGilchrist. Statistical algorithms: Algorithm AS 34: Sequential inversion of band matrices. Applied Statistics, 19(3):290–292, September 1970. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/34.

Box:1971:PEC

[71] M. J. Box. A parameter estimation criterion for multiresponse models applicable when some observations are missing. Applied Statistics, 20(1):1–7, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Garside:1971:SCP

[72] M. J. Garside. Some computational procedures for the best subset problem. Applied Statistics, 20(1):8–15, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quenouille:1971:PCD

[73] M. H. Quenouille and J. A. John. Paired comparison designs for 2^{n} factorials. Applied Statistics, 20(1):16–24, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:SCS

[74] J. A. Nelder. Statistical computing 1970: Statistical computing and computer languages. Applied Statistics, 20(1):25–32, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1971:SCI

[75] J. C. Gower and I. D. Hill. Statistical computing 1970: Internal data structures. Applied Statistics, 20(1):32–45, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Godfrey:1971:SCO

[76] Michael D. Godfrey. Statistical computing 1970: Operating system considerations for statistical computing. Applied Statistics, 20(1):45–56, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:SCI

[77] J. A. Nelder and B. E. Cooper. Statistical computing 1970: Input/output in statistical programming. Applied Statistics, 20(1):56–73, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanReeken:1971:SCR

[78] A. J. van Reeken. Statistical computing 1970: Report on the Work of the Dutch Working Party on Statistical Computing. Applied Statistics, 20(1):73–79, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hilton:1971:MAD

[79] Gordon Hilton. Miscellanea: An algorithm for detecting differences between transition probability matrices. Applied Statistics, 20(1):80–86, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hsi:1971:MGT

[80] B. P. Hsi and D. M. Burych. Miscellanea: Games of two players. Applied Statistics, 20(1):86–92, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gulliver:1971:BRA

[81] S. Gulliver. Book reviews: Abstract of Statistics 1968–69 for the City of Birmingham, by Alan B. Neale and Gillian Haine. Applied Statistics, 20(1):93, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1971:BRI

[82] J. N. R. Jeffers. Book reviews: An Introduction to Mathematical Ecology, by E. C. Pielou. Applied Statistics, 20(1):93– 94, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1971:BRS

[83] C. Chatfield. Book reviews: Statistics and Probability, by J. H. Durran. Applied Statistics, 20(1):94–95, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holmes:1971:BRP

[84] C. Holmes. Book reviews: Professional Mail Surveys, by Paul L. Erdos. Applied Statistics, 20(1):95–96, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1971:BRA

[85] J. Luckman. Book reviews: Abridged Tables for Use of Examination Candidates, by The Institute of Statisticians. Applied Statistics, 20(1):96, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1971:BRO

[86] D. V. Lindley. Book reviews: Optimal Statistical Decisions, by Morris de Groot. Applied Statistics, 20(1):97, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1971:BRS

[87] G. M. Clarke. Book reviews: Statistics, With Applications to the Biological and Health Sciences, by R. D. Remington and M. A. Schork. Applied Statistics, 20(1):97–98, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1971:SAAa

[88] I. D. Hill and Richard Peto. Statistical algorithms: Algorithm AS 35:

Probabilities derived from finite populations. Applied Statistics, 20(1):99– 105, March 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/35. See corrigenda [138, 218] and remark [622].

Thomas:1971:SAA

[89] Donald G. Thomas. Statistical algorithms: Algorithm AS 36: Exact confidence limits for the odds ratio in a 2×2 table. Applied Statistics, 20(1):105– 110, March 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/36.

Garside:1971:SAAa

[90] M. J. Garside. Statistical algorithms: Algorithm AS 37: Inversion of a symmetric matrix. Applied Statistics, 20 (1):111–112, March 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/37. See remark

[319].

Garside:1971:SAAb

[91] M. J. Garside. Statistical algorithms: Algorithm AS 38: Best subset search. Applied Statistics, 20(1):112– 115, March 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/38.

Hill:1971:SAAb

[92] I. D. Hill. Statistical algorithms: Algorithm AS 39: Arrays with a variable number of dimensions. Applied Statistics, 20(1):115–117, March 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/39.

Jennrich:1971:SAR

[93] R. I. Jennrich and P. I. Sampson. Statistical algorithms: Remark AS R3: A remark on Algorithm AS 10. Applied Statistics, 20(1):117–118, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/10. See [5,

117].

Sparks:1971:SAC

[94] D. N. Sparks. Statistical algorithms: Corrigenda: Algorithm AS 30: Half normal plotting. Applied Statistics, 20(1):118, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/30. See [51, 217].

Thayer:1971:SAC

[95] R. P. Thayer and R. F. Storer. Statistical algorithms: Corrigenda: Algorithm AS 21: Scale selection for computer plots. Applied Statistics, 20(1):118, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/21. See [10, 347].

Anonymous:1971:VIP

[96] Anonymous. Volume information. Applied Statistics, 20(1):[3], 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cohen:1971:AAD

[97] Kalman J. Cohen, Wolfgang Muller, and Manfred W. Padberg. Autoregressive approaches to disaggregation of time series data. Applied Statistics, 20 (2):119–129, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashton:1971:GAP

[98] Winifred D. Ashton. Gap-acceptance problems at a traffic intersection. Applied Statistics, 20(2):130–138, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sen:1971:SRD

[99] A. R. Sen. Some recent developments in waterfowl sample survey techniques. Applied Statistics, 20(2):139–147, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Warren:1971:CRB

[100] W. G. Warren. Correlation or regression: Bias or precision. Applied Statistics, 20 (2):148–164, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goodman:1971:SST

[101] Leo A. Goodman. A simple simultaneous test procedure for quasiindependence in contingency tables. Applied Statistics, 20(2):165–177, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Joffe:1971:PMT

[102] A. D. Joffe. On the productivity of machines tended by a single operator. Applied Statistics, 20(2):178–186, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moran:1971:MIS

[103] P. A. P. Moran and J. Neyman. Miscellanea: The IASPS and its satellite symposia. Applied Statistics, 20(2):187–192, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roger:1971:MCC

[104] J. H. Roger and R. G. Carpenter. Miscellanea: The cumulative construction of minimum spanning trees. Applied Statistics, 20(2):192–194, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1971:BRD

[105] D. H. Ward. Book reviews: Dealing with Data, by Arthur J. Lyon. Applied Statistics, 20(2):195, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1971:BRB

[106] J. C. Gower. Book reviews: Les Bases de la Classification Automatique, by I. C. Lerman. Applied Statistics, 20 (2):195–197, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanSlooten:1971:BRM

[107] R. Van Slooten. Book reviews: Mathematical Models of the Economy and Other Essays, by R. Stone. Applied Statistics, 20(2):198, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pridmore:1971:BRM

[108] W. A. Pridmore. Book reviews: Managing with Statistics, by T. W. Lewis and R. M. Fox. Applied Statistics, 20(2):199, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newman:1971:BRP

[109] G. B. Newman. Book reviews: Planning and Analysis of Clinical Studies, by Walter J. Burdette and Edmund A. Gehan. Applied Statistics, 20(2):199–200, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1971:BRR

[110] R. J. Nicholson. Book reviews: Regression and Econometric Methods, by David S. Huang. Applied Statistics, 20 (2):200–201, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bassett:1971:BRM

[111] E. E. Bassett. Book reviews: Mathematics and Sociology, by Roy Mapes. Applied Statistics, 20(2):201–202, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1971:LE

[112] B. E. Cooper. Letter to the editors. Applied Statistics, 20(2):203, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roger:1971:SAA

[113] J. H. Roger. Statistical algorithms: Algorithm AS 40: Updating a minimum spanning tree. Applied Statistics, 20(2): 204–206, June 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/40.

Clarke:1971:SAA

[114] M. R. B. Clarke. Statistical algorithms: Algorithm AS 41: Updating the sample mean and dispersion matrix. Applied Statistics, 20(2):206– 209, June 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/41.

Cooper:1971:SAA

[115] B. E. Cooper. Statistical algorithms:

Algorithm AS 42: The use of orthogonal polynomials with equal xvalues. Applied Statistics, 20(2):209– 213, June 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/42.

Yates:1971:SAA

[116] F. Yates. Statistical algorithms: Algorithm AS 43: Variable format in Fortran. Applied Statistics, 20(2):213– 216, June 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/43. See corrigenda [139].

Cooper:1971:SAR

[117] B. E. Cooper. Statistical algorithms:

Remark AS R4: A remark on Algorithm AS 10: “The Use of Orthogonal Polynomials”. Applied Statistics, 20(2):216, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/10. See [5, 93].

Chatfield:1971:TSA

[118] C. Chatfield and M. P. G. Pepper. Timeseries analysis: An example from geophysical data. Applied Statistics, 20 (3):217–238, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nathanson:1971:AMA

[119] James A. Nathanson. An application of multivariate analysis in astronomy. Applied Statistics, 20(3):239–249, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snee:1971:SDA

[120] Ronald D. Snee and Horace P. Andrews. Statistical design and analysis of shape studies. Applied Statistics, 20 (3):250–258, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

MacDonald:1971:AEM

[121] Peter MacDonald. The analysis of a 2 n experiment by means of ranks. Applied Statistics, 20(3):259–275, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1971:IEL

[122] A. J. Scott and T. M. F. Smith. Interval estimates for linear combinations of means. Applied Statistics, 20(3): 276–285, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Charatsis:1971:CPE

[123] E. G. Charatsis. A computer program for estimation of the constant elasticity of substitution production function. Applied Statistics, 20(3):286–296, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yuen:1971:MNW

[124] Karen K. Yuen. Miscellanea: A note on Winsorized t. Applied Statistics, 20 (3):297–304, 1971. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Veevers:1971:MVS

[125] A. Veevers and M. C. K. Tweedie.

Miscellanea: Variance-stabilizing transformation of a Poisson variate by a beta function. Applied Statistics, 20 (3):304–308, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnston:1971:MCH

[126] W. Johnston. Miscellanea: The case history of a simulation study. Applied Statistics, 20(3):308–313, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nath:1971:MET

[127] G. Baikunth Nath. Miscellanea: Estimation in truncated bivariate normal distributions. Applied Statistics, 20 (3):313–319, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1971:MRM

[128] I. D. Hill. Miscellanea: Report on a medical statistical conference. Applied Statistics, 20(3):319–321, 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bryant:1971:BRB

[129] D. T. Bryant. Book reviews: Basic Statistics for Social Research, by Dean J. Champion. Applied Statistics, 20 (3):322–323, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1971:BRN

[130] R. A. Cooper. Book reviews: National Economic Accounting, by Carleen O’Loughlin. Applied Statistics, 20 (3):323–324, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Murphy:1971:BRS

[131] B. P. Murphy. Book reviews: Solving Problems in Statistics II, by J. Aitchison. Applied Statistics, 20(3):324–325, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashford:1971:BRG

[132] J. R. Ashford. Book reviews: The Generation of Random Variates, by T. G. Newman and P. L. Odell. Applied Statistics, 20(3):325, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1971:BRA

[133] P. R. Cox. Book reviews: The Analysis of Mortality and Other Actuarial Statistics, by B. Benjamin and H. W. Haycocks. Applied Statistics, 20(3):325–326, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1971:SAAa

[134] D. N. Sparks. Statistical algorithms: Algorithm AS 44: Scatter diagram plotting. Applied Statistics, 20(3):327–331, September 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/44.

Sparks:1971:SAAb

[135] D. N. Sparks. Statistical algorithms:

Algorithm AS 45: Histogram plotting. Applied Statistics, 20(3):332–335, September 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/45. See corrigenda [270].

Clayton:1971:SAA

[136] D. G. Clayton. Statistical algorithms:

Algorithm AS 46: Gram–Schmidt orthogonalization. Applied Statistics, 20 (3):335–338, September 1971. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/46.

ONeill:1971:SAA

[137] R. O’Neill. Statistical algorithms: Algorithm AS 47: Function minimization using a simplex procedure. Applied Statistics, 20(3):338–345, September 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/47. See remarks [348, 499, 749] and corrigendum [349].

Hill:1971:SAC

[138] I. D. Hill and Richard Peto. Statistical algorithms: Corrigenda: Algorithm AS 35: Probabilities derived from finite populations. Applied Statistics, 20(3):346, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/35. See [88, 218, 622].

Yates:1971:SAC

[139] F. Yates. Statistical algorithms: Corrigenda: Algorithm AS 43: Variable format in Fortran. Applied Statistics, 20(3):346, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/43. See [116].

Kowalski:1972:ENN

[140] Charles J. Kowalski. On the effects of non-normality on the distribution of the sample product-moment correlation coefficient. Applied Statistics, 21(1):1–12, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1972:EDC

[141] M. J. Box and N. R. Draper. Estimation and design criteria for multiresponse non-linear models with nonhomogeneous variance. Applied Statistics, 21(1):13–24, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Craig:1972:ICS

[142] J. Craig and Anne Driver. The identification and comparison of small areas of adverse social conditions. Applied Statistics, 21(1):25–35, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burman:1972:DEE

[143] J. P. Burman. Data editing and estimation: [introduction]. Applied Statistics, 21(1):36, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1972:DEE

[144] I. D. Hill. Data editing and estimation: The use of Algol in a bronchitis survey. Applied Statistics, 21(1):37–41, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chapman:1972:DEE

[145] J. S. Chapman. Data editing and estimation: Surveys in town-planning. Applied Statistics, 21(1):42–45, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newing:1972:DEE

[146] C. R. Newing. Data editing and estimation: Data arrangement and checking in the insurance business. Applied Statistics, 21(1):45–53, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Richardson:1972:DEE

[147] M. Richardson and A. D. Cunningham.

Data editing and estimation: Vetting of industrial survey questionnaires by computer. Applied Statistics, 21(1):53–55, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Flaxen:1972:DEE

[148] D. W. Flaxen. Data editing and estimation: Credibility tests and data editing in the new earnings survey. Applied Statistics, 21(1):55–57, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1972:DEE

[149] W. F. F. Kemsley. Data editing and estimation: Pre-computer editing of budgets for the family expenditure survey. Applied Statistics, 21(1):58–64, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1972:DEE

[150] B. E. Cooper. Data editing and estimation: Computing aspects of data management. Applied Statistics, 21(1):65– 75, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kihlberg:1972:MSR

[151] J. K. Kihlberg, J. H. Herson, and W. E. Schotz. Miscellanea: Square root transformation revisited. Applied Statistics, 21(1):76–81, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ahuja:1972:MRR

[152] J. C. Ahuja. Miscellanea: Recurrence relation for minimum variance unbiased estimation of certain left-truncated Poisson distributions. Applied Statistics, 21 (1):81–86, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cherrington:1972:MUH

[153] Norma K. Cherrington and J. V. Smart. Miscellanea: Use of Hotelling’s T^{2} statistic in testing the significance of individual regressors in multivariate regression. Applied Statistics, 21(1):86–88, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:MA

[154] Anonymous. Miscellanea: Addendum. Applied Statistics, 21(1):89, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jazairi:1972:MFI

[155] N. T. Jazairi. Miscellanea: Fisher ideal index numbers formula: A note. Applied Statistics, 21(1):89–92, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

LaMotte:1972:MSR

[156] L. R. LaMotte. Miscellanea: The SE-LECT routines: A program for identifying best subset regressions. Applied Statistics, 21(1):92–93, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1972:BRP

[157] John Luckman. Book reviews: Principles of Operations Research, with Applications to Managerial Decisions, by Harvey M. Wagner. Applied Statistics, 21(1):94, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smithson:1972:BRM

[158] J. F. Smithson. Book reviews: Mathematical Models in Marketing, by Robert B. Murdick. Applied Statistics, 21(1):94, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1972:BRP

[159] J. N. R. Jeffers. Book reviews: Principles of Management Science, with Applications to Executive Decisions, by Harvey M. Wagner. Applied Statistics, 21(1):95, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snell:1972:BRT

[160] E. J. Snell. Book reviews: The Teaching of Probability and Statistics, by Lenart Rade. Applied Statistics, 21(1):95, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Herne:1972:LE

[161] H. Herne. Letter to the editors. Applied Statistics, 21(1):96, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Phillips:1972:SAAa

[162] J. P. N. Phillips. Statistical algorithms: Algorithm AS 48: Uncertainty function for a binary sequence. Applied Statistics, 21(1):97–99, March 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/48.

Phillips:1972:SAAb

[163] J. P. N. Phillips. Statistical algorithms: Algorithm AS 49: Autocorrelation function for a binary sequence. Applied Statistics, 21(1):100– 103, March 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/49.

Thomas:1972:SAA

[164] Donald G. Thomas. Statistical algorithms: Algorithm AS 50: Tests of fit for a one-hit vs. two-hit curve. Applied Statistics, 21(1):103–112, March 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/50.

Anonymous:1972:BMPa

[165] Anonymous. Back matter. Applied Statistics, 21(1):[6], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:CDE

[166] Anonymous. Conference on data editing and estimation. Applied Statistics, 21(1):??, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:FMPa

[167] Anonymous. Front matter. Applied Statistics, 21(1):[3], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:VIP

[168] Anonymous. Volume information. Applied Statistics, 21(1):[3], 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1972:AMB

[169] D. R. Cox. The analysis of multivariate binary data. Applied Statistics, 21 (2):113–120, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kupper:1972:FSS

[170] Lawrence L. Kupper. Fourier series and spherical harmonics regression. Applied Statistics, 21(2):121–130, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sheppard:1972:OSD

[171] D. Sheppard. Obtaining samples of drivers for surveys. Applied Statistics, 21(2):131–135, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rubin:1972:NIA

[172] Donald B. Rubin. A non-iterative algorithm for least squares estimation of missing values in any analysis of variance design. Applied Statistics, 21 (2):136–141, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fryer:1972:MLE

[173] J. G. Fryer and R. J. Pethybridge. Maximum likelihood estimation of a linear regression function with grouped data. Applied Statistics, 21(2):142–154, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1972:ALE

[174] M. J. R. Healy. Animal litters as experimental units. Applied Statistics, 21 (2):155–159, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1972:DVP

[175] I. T. Jolliffe. Discarding variables in a principal component analysis. I: Artificial data. Applied Statistics, 21 (2):160–173, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Richards:1972:RED

[176] Larry E. Richards. Refinement and extension of distribution-free discriminate analysis. Applied Statistics, 21 (2):174–176, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ekblom:1972:MCI

[177] H. Ekblom. A Monte Carlo investigation of mode estimators in small samples. Applied Statistics, 21(2):177–184, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hodges:1972:DUL

[178] S. D. Hodges and P. G. Moore. Data uncertainties and least squares regression. Applied Statistics, 21(2):185–195, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carter:1972:MTA

[179] Michael C. Carter. A model for thunderstorm activity: Use of the compound negative binomial-positive binomial distribution. Applied Statistics, 21 (2):196–201, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1972:MDP

[180] M. J. R. Healy. Miscellanea: Drawing a probability ellipse. Applied Statistics, 21 (2):202–204, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Montgomery:1972:MNL

[181] David B. Montgomery. Miscellanea:

Note on a limit distribution arising in certain stochastic response models. Applied Statistics, 21(2):204–207, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1972:BRS

[182] D. N. Sparks. Book reviews: Scientific Marketing Research, by Gerald Albaum and M. Venkatesan. Applied Statistics, 21(2):208, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beale:1972:BRI

[183] E. M. L. Beale. Book reviews: Introduction to Optimization Techniques: Fundamentals and Applications of Nonlinear Programming, by Masanao Aoki. Applied Statistics, 21(2):208–209, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cliff:1972:BRR

[184] A. D. Cliff. Book reviews: Regional Variations in Britain: Studies in Economic and Social Geography, by B. E. Coates and E. M. Rawstron. Applied Statistics, 21(2):209–210, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:1972:BRP

[185] J. A. Lewis. Book reviews: Probit Analysis (3rd ed), by D. Finney. Applied Statistics, 21(2):210–211, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1972:BRS

[186] Pamela G. Walker. Book reviews: Statistical Functions, by Buddy L. Myers and Norbert L. Enrick. Applied Statistics, 21(2):211, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1972:BRA

[187] John Luckman. Book reviews: Applications of Queueing Theory, by G. F. Newell. Applied Statistics, 21(2):212, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ellis:1972:BRQ

[188] S. H. Ellis. Book reviews: Quality Control Workbook: Procedures for Implementing Statistical Quality Control, by David J. Desmond. Applied Statistics, 21(2):212–213, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1972:BRE

[189] D. V. Lindley. Book reviews: Elements of Decision Theory, by B. W. Lindgren. Applied Statistics, 21(2):213–214, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maunder:1972:BRS

[190] W. F. Maunder. Book reviews: Sources of Statistics, by Joan M. Harvey. Applied Statistics, 21(2):214, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hills:1972:BRB

[191] Michael Hills. Book reviews: Bancroft’s Introduction to Biostatistics, by Johannes Ipsen and Polly Feigl. Applied Statistics, 21(2):214–215, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Golikova:1972:LE

[192] T. Golikova and L. Panchenko. Letter to the editors. Applied Statistics, 21 (2):216–217, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haberman:1972:SAA

[193] S. J. Haberman. Statistical algorithms:

Algorithm AS 51: Log-linear fit for contingency tables. Applied Statistics, 21 (2):218–225, June 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/51.

Spicer:1972:SAA

[194] C. C. Spicer. Statistical algorithms: Algorithm AS 52: Calculation of power sums of deviations about the mean. Applied Statistics, 21(2):226– 227, June 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/52.

Anonymous:1972:BMPb

[195] Anonymous. Back matter. Applied Statistics, 21(2):[7], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:FMPb

[196] Anonymous. Front matter. Applied Statistics, 21(2):[3], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fairweather:1972:MOE

[197] William R. Fairweather. A method of obtaining an exact confidence interval for the common mean of several normal populations. Applied Statistics, 21 (3):229–233, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindsey:1972:FRS

[198] J. K. Lindsey. Fitting response surfaces with power transformations. Applied Statistics, 21(3):234–247, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1972:SMT

[199] C. J. Taylor. A stochastic model of temperature variations at weather stations in Britain. Applied Statistics, 21 (3):248–260, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1972:MAB

[200] K. E. Kemp and D. F. Butcher. Methods of analysing binomial data in a two-factor experiment without replication compared by computer simulation. Applied Statistics, 21(3):261–273, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pagano:1972:AFA

[201] Marcello Pagano. An algorithm for fitting autoregressive schemes. Applied Statistics, 21(3):274–281, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conway-Jones:1972:ASB

[202] A. H. Conway-Jones. Analysis of smallbore shooting scores. Applied Statistics, 21(3):282–296, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

ONeill:1972:SRA

[203] T. S. O’Neill. System reliability assessment from its components. Applied Statistics, 21(3):297–320, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Panneton:1972:EDK

[204] M. Panneton and P. Robillard. The exact distribution of Kendall’s S with ties in one ranking. Applied Statistics, 21 (3):321–323, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davenport:1972:MSA

[205] M. Davenport and G. Studdert-Kennedy. Miscellanea: The statistical analysis of aesthetic judgment: An exploration. Applied Statistics, 21(3): 324–333, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Michelini:1972:MEE

[206] C. Michelini. Miscellanea: Estimating the exponential growth function by direct least squares: A comment. Applied Statistics, 21(3):333–335, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1972:BRF

[207] I. D. Hill. Book reviews: Faster Than Thought. A Symposium on Digital Computing Machines, by The Earl of Halsbury and B. V. Bowden. Applied Statistics, 21(3):336, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1972:BRC

[208] J. N. R. Jeffers. Book reviews: Computer Programming in Quantitative Biology, by R. G. Davies. Applied Statistics, 21(3):336–337, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burrows:1972:BRD

[209] C. P. Burrows. Book reviews: Distribution Management: Mathematical Modelling and Practical Analysis, by Samuel Eilon, C. D. T. Watson-Gandy and Nicos Christofides. Applied Statistics, 21 (3):337–338, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snell:1972:BRA

[210] E. J. Snell. Book reviews: Advanced General Statistics, by B. C. Erricker. Applied Statistics, 21(3):338, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beale:1972:BRC

[211] E. M. L. Beale. Book reviews: Cost Considerations in Systems Analysis, by Gene H. Fisher. Applied Statistics, 21 (3):338–339, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1972:BRF

[212] J. C. Gower. Book reviews: Factor Analysis as a Statistical Method. 2nd ed, by D. N. Lawley and A. E. Maxwell. Applied Statistics, 21(3):339–340, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1972:BRS

[213] C. Chatfield. Book reviews: Statistics for Physicists, by B. R. Martin. Applied Statistics, 21(3):340, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1972:SAA

[214] W. B. Smith and R. R. Hocking.

Statistical algorithms: Algorithm AS 53: Wishart variate generator. Applied Statistics, 21(3):341–345, September 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/53.

Panneton:1972:SAA

[215] M. Panneton and P. Robillard. Statistical algorithms: Algorithm AS 54: Kendall’s S frequency distribution. Applied Statistics, 21(3):345–348, September 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/54.

Odeh:1972:SAA

[216] R. E. Odeh. Statistical algorithms: Algorithm AS 55: The generalized Mann– Whitney U-statistic. Applied Statistics, 21(3):348–351, September 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/55.

Munford:1972:SAR

[217] A. G. Munford. Statistical algorithms: Remark AS R5: A remark on Algorithm AS 30: Half normal plotting. Applied Statistics, 21(3):351, 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/30. See [51, 94].

Hill:1972:SAR

[218] I. D. Hill and Richard Peto. Statistical algorithms: Remark AS R6: A remark on Algorithm AS 35: Probabilities derived from finite populations. Applied Statistics, 21(3):352, 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/35. See [88,

138, 622].

Anonymous:1972:BMPc

[219] Anonymous. Back matter. Applied Statistics, 21(3):[8], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1972:FMPc

[220] Anonymous. Front matter. Applied Statistics, 21(3):[3], 1972. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mead:1973:EG

[221] R. Mead and K. H. Freeman. An experiment game. Applied Statistics, 22(1):1– 6, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1973:SIR

[222] Christopher Chatfield. Statistical inference regarding Markov chain models. Applied Statistics, 22(1):7–20, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1973:DVP

[223] I. T. Jolliffe. Discarding variables in a principal component analysis. II: Real data. Applied Statistics, 22(1):21–31, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1973:ERT

[224] Andrew W. Evans. The estimation of road traffic flow by registration number observation. Applied Statistics, 22 (1):32–49, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sichel:1973:STT

[225] H. S. Sichel. On a significance test for two Poisson variables. Applied Statistics, 22(1):50–58, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1973:RSF

[226] A. Springall. Response surface fitting using a generalization of the Bradley-Terry paired comparison model. Applied Statistics, 22(1):59–68, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kiountouzis:1973:LPT

[227] E. A. Kiountouzis. Linear programming techniques in regression analysis. Applied Statistics, 22(1):69–73, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bunday:1973:EBD

[228] B. D. Bunday and C. Mack. Efficiency of bi-directionally traversed machines. Applied Statistics, 22(1):74–81, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Borth:1973:MOM

[229] David M. Borth. A modification of Owen’s method for computing the bivariate normal integral. Applied Statistics, 22(1):82–85, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peto:1973:ESC

[230] Richard Peto. Experimental survival curves for interval-censored data. Applied Statistics, 22(1):86–91, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neave:1973:MUB

[231] Henry R. Neave. Miscellanea: On using the Box-Muller transformation with multiplicative congruential pseudorandom number generators. Applied Statistics, 22(1):92–97, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

George:1973:MOS

[232] S. L. George. Miscellanea: Optimal strategy in tennis: A simple probabilistic model. Applied Statistics, 22 (1):97–104, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boyd:1973:MNP

[233] K. T. Boyd. Miscellanea: A note on paired comparison designs for 2 n factorials. Applied Statistics, 22(1):104–106, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1973:BRD

[234] A. C. Atkinson. Book reviews: Developments in Operations Research Vol. I, by B. Avi-Itzhak. Applied Statistics, 22(1):107, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1973:BRP

[235] C. Chatfield. Book reviews: Probability and Statistics in Engineering and Management Science, by William W. Hines and Douglas C. Montgomery. Applied Statistics, 22(1):107–108, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wishart:1973:BRR

[236] D. M. G. Wishart. Book reviews: Russian-English/English-Russian Glossary of Statistical Terms, by Samuel Kotz. Applied Statistics, 22(1):108, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Veevers:1973:BRS

[237] A. Veevers. Book reviews: Statistical Design and Analysis of Experiments, by Peter W. M. John. Applied Statistics, 22 (1):108–109, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1973:BRD

[238] R. J. Nicholson. Book reviews: The Distribution of Personal Income in the United Kingdom 1949–63, by Thomas Stark. Applied Statistics, 22(1):109–110, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1973:BRA

[239] R. J. Nicholson. Book reviews: The Analysis of Family Budgets, by S. J. Prais and H. S. Houthakker. Applied Statistics, 22(1):110, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1973:BRA

[240] Monica A. Walker. Book reviews: The Amalgamation of Police Districts. Reconciliation of Statistics: 1960–1970. Part A, by Christian Economic and Social Research Foundation. Applied Statistics, 22(1):110–111, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1973:BRI

[241] J. N. R. Jeffers. Book reviews: An Introduction to Computer Programming Using the BASIC Language, by William F. Sharpe and Nancy L. Jacob. Applied Statistics, 22(1):111, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peto:1973:SAA

[242] Richard Peto. Statistical algorithms:

Algorithm AS 56: Permutational significance testing. Applied Statistics, 22 (1):112–118, March 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/56.

Haberman:1973:SAA

[243] S. J. Haberman. Statistical algorithms:

Algorithm AS 57: Printing multidimensional tables. Applied Statistics, 22(1): 118–126, March 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/57.

Sparks:1973:SAAa

[244] D. N. Sparks. Statistical algorithms:

Algorithm AS 58: Euclidean cluster analysis. Applied Statistics, 22(1):126– 130, March 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/58. See remark [417].

Freeman:1973:SAA

[245] P. R. Freeman. Statistical algorithms:

Algorithm AS 59: Hypergeometric probabilities. Applied Statistics, 22(1):130– 133, March 1973. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/59.

Freeman:1973:SAR

[246] P. R. Freeman. Statistical algorithms: Remark AS R7: A remark on Algorithm AS 26: Ranking an array of numbers. Applied Statistics, 22(1):133, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

                                        /lib.stat.cmu.edu/apstat/26. See

[30].

Anonymous:1973:BMPa

[247] Anonymous. Back matter. Applied Statistics, 22(1):[6], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1973:FMPa

[248] Anonymous. Front matter. Applied Statistics, 22(1):[3], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1973:VIP

[249] Anonymous. Volume information. Applied Statistics, 22(1):[3], 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shearer:1973:MDQ

[250] P. R. Shearer. Missing data in quantitative designs. Applied Statistics, 22 (2):135–140, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:1973:CAF

[251] P. L. Goldsmith and R. Boddy. Critical analysis of factorial experiments and orthogonal fractions. Applied Statistics, 22(2):141–160, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yates:1973:ASC

[252] F. Yates. The analysis of surveys on computers-features of the Rothamsted survey program. Applied Statistics, 22 (2):161–171, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kullback:1973:PSO

[253] S. Kullback and Marian Fisher. Partitioning second-order interaction in three-way contingency tables. Applied Statistics, 22(2):172–184, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lochner:1973:BMC

[254] Robert H. Lochner. A Bayesian method for comparing the means of multinomial populations. Applied Statistics, 22 (2):185–191, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Andrews:1973:TPD

[255] David F. Andrews and John W. Tukey.

Teletypewriter plots for data analysis can be fast: 6-line plots, including probability plots. Applied Statistics, 22 (2):192–202, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pethybridge:1973:MLE

[256] R. J. Pethybridge. Maximum likelihood estimation of a polynomial regression function with grouped data. Applied Statistics, 22(2):203–212, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Golder:1973:UCA

[257] P. A. Golder and K. A. Yeomans. The use of cluster analysis for stratification. Applied Statistics, 22(2):213–219, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1973:CFS

[258] D. N. Sparks and A. D. Todd. A comparison of FORTRAN subroutines for calculating latent roots and vectors. Applied Statistics, 22(2):220–225, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1973:PMV

[259] A. F. Bissell. Process monitoring with variable element sizes. Applied Statistics, 22(2):226–238, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1973:MFS

[260] J. Taylor. A method of fitting several linear functional relations and of testing for differences between them. Applied Statistics, 22(2):239–248, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ansell:1973:MRL

[261] Margaret J. G. Ansell. Miscellanea: Robustness of location estimators to asymmetry. Applied Statistics, 22(2):249–254, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stone:1973:BRS

[262] M. Stone. Book reviews: Statistical Papers in Honor of George W. Snedecor, by T. A. Bancroft and Susan Alice Brown. Applied Statistics, 22(2):255, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1973:BRS

[263] D. H. Ward. Book reviews: Statistical Methods in Research and Production with Special Reference to the Chemical Industry. 4th Revised, by Owen L. Davies and Peter L. Goldsmith. Applied Statistics, 22(2):256, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stuart:1973:BRL

[264] Alan Stuart. Book reviews: Likelihood.

An Account of the Statistical Concept of Likelihood and Its Application to Scientific Inference, by A. W. F. Edwards. Applied Statistics, 22(2):257, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1973:BRR

[265] D. N. Sparks. Book reviews: Repeat-Buying: Theory and Applications, by A. S. C. Ehrenberg. Applied Statistics, 22(2):258, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkinson:1973:BRSa

[266] R. K. Wilkinson. Book reviews: Statistical Methods for the Analysis of Earnings Data with Special Application to Salaries in Swedish Industry, by Anders Klevmarken. Applied Statistics, 22 (2):258–259, 1973. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1973:SAAb

[267] D. N. Sparks and A. D. Todd. Statistical algorithms: Algorithm AS 60: Latent roots and vectors of a symmetric matrix. Applied Statistics, 22(2): 260–265, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/60. See corrigendum [320].

Andrews:1973:SAA

[268] David F. Andrews and John W. Tukey. Statistical algorithms: Algorithm AS 61: Six-line plots. Applied Statistics, 22(2): 265–269, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/61. See remark [652].

Dinneen:1973:SAA

[269] L. C. Dinneen and B. C. Blakesley. Statistical algorithms: Algorithm AS 62: A generator for the sampling distribution of the Mann–Whitney U statistic. Applied Statistics, 22(2):269– 273, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/62.

Sparks:1973:SAC

[270] D. N. Sparks. Statistical algorithms: Corrigendum: Algorithm AS 45: Histogram plotting. Applied Statistics, 22(2):274, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [135].

Anonymous:1973:BMPb

[271] Anonymous. Back matter. Applied Statistics, 22(2):[8], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1973:FMPb

[272] Anonymous. Front matter. Applied Statistics, 22(2):[3], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1973:IAR

[273] Douglas M. Hawkins. On the investigation of alternative regressions by principal component analysis. Applied Statistics, 22(3):275–286, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dyer:1973:MEP

[274] Danny D. Dyer. On moments estimation of the parameters of a truncated bivariate normal distribution. Applied Statistics, 22(3):287–291, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seeger:1973:TST

[275] Paul Seeger and Ulf Thorsson. Twosided tolerance limits with two-stage sampling from normal populations: Monte Carlo studies of the distribution of coverages. Applied Statistics, 22 (3):292–300, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prentice:1973:MMR

[276] R. L. Prentice and A. El. Shaarawi. A model for mortality rates and a test of fit for the Gompertz force of mortality. Applied Statistics, 22(3):301–314, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunningham:1973:PBS

[277] S. W. Cunningham. The predictability of British stock market prices. Applied Statistics, 22(3):315–331, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Littlewood:1973:BRG

[278] B. Littlewood and J. L. Verrall. A Bayesian reliability growth model for computer software. Applied Statistics, 22(3):332–346, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gross:1973:MCS

[279] A. M. Gross. A Monte Carlo swindle for estimators of location. Applied Statistics, 22(3):347–353, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McDonald:1973:ARE

[280] John McDonald. An analysis of residual error in the quarterly national accounts of the U.K. Applied Statistics, 22 (3):354–367, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ginsburgh:1973:FND

[281] Victor A. Ginsburgh. A further note on the derivation of quarterly figures consistent with annual data. Applied Statistics, 22(3):368–374, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pocock:1973:DVS

[282] S. J. Pocock. Daily variations in sickness absence. Applied Statistics, 22 (3):375–391, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkinson:1973:SDF

[283] G. N. Wilkinson and C. E. Rogers. Symbolic description of factorial models for analysis of variance. Applied Statistics, 22(3):392–399, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1973:BRI

[284] G. B. Wetherill. Book reviews: Introductory Statistics: A Service Course, by A. H. Pollard. Applied Statistics, 22(3):400, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Preece:1973:BRI

[285] D. A. Preece. Book reviews: Introduction to Probability and Statistics. 5th Ed, by H. L. Alder and E. B. Roessler. Applied Statistics, 22(3):400–401, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Herne:1973:BRH

[286] H. Herne. Book reviews: How to Lie with Statistics, by Darrell Huff. Applied Statistics, 22(3):401–402, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Richardson:1973:BRS

[287] Harry W. Richardson. Book reviews:

Surveys of Personal Income in London, by R. U. Redpath, M. G. Powell and S. Kingaby. Applied Statistics, 22 (3):402–403, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1973:BRI

[288] J. Taylor. Book reviews: An Introduction to Statistical Science in Agriculture. 4th ed, by D. J. Finney. Applied Statistics, 22(3):403, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1973:BRF

[289] I. D. Hill. Book reviews: Fortran Techniques, With Special Reference to Non-Numerical Applications, by A. Colin Day. Applied Statistics, 22(3):404, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1973:BRN

[290] R. J. Nicholson. Book reviews: National Income, Expenditure and Output of the United Kingdom 1855–1965, by C. H. Feinstein. Applied Statistics, 22 (3):404–405, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smithson:1973:BRM

[291] J. F. Smithson. Book reviews: Mathematical Model Building in Economics and Industry. Second Series, by Scientific Control Systems Ltd. and M. G. Kendall. Applied Statistics, 22(3): 405–406, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1973:BRS

[292] Anonymous. Book reviews: Statistical Methods for the Analysis of Earnings Data with Special Application to Salaries in Swedish Industry: Corrigendum. Applied Statistics, 22(3):406, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkinson:1973:BRSb

[293] R. K. Wilkinson. Book reviews: Statistical Methods for the Analysis of Earnings Data with Special Application to Salaries in Swedish Industry, by Anders Klevmarken. Applied Statistics, 22(3):406, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1973:LE

[294] D. A. Williams. Letter to the editors. Applied Statistics, 22(3):407–408, 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Majumder:1973:SAAa

[295] K. L. Majumder and G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 63: The incomplete beta integral. Applied Statistics, 22(3):409–411, September 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/63. See remark [589].

Majumder:1973:SAAb

[296] K. L. Majumder and G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 64: Inverse of the incomplete beta function ratio. Applied Statistics, 22(3): 411–414, September 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/64. See remark

Rogers:1973:SAA

[297] C. E. Rogers. Statistical algorithms: Algorithm AS 65: Interpreting structure formulae. Applied Statistics, 22(3): 414–424, September 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/65. See remark

[885].

Hill:1973:SAA

[298] I. D. Hill. Statistical algorithms: Algorithm AS 66: The normal integral. Applied Statistics, 22(3):424–427, September 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/66.

Hitchin:1973:SAR

[299] D. Hitchin. Statistical algorithms: Remark AS R8: A remark on algorithms AS 4 and AS 5: An auxiliary function for distribution integrals, and the integral of the non-central t- distribution. Applied Statistics, 22(3):428, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [1, 2].

Anonymous:1973:BMPc

[300] Anonymous. Back matter. Applied Statistics, 22(3):[6], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1973:FMPc

[301] Anonymous. Front matter. Applied Statistics, 22(3):[3], 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bhattacharyya:1974:FDT

[302] M. N. Bhattacharyya. Forecasting the demand for telephones in australia. Applied Statistics, 23(1):1–10, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Narula:1974:PMS

[303] Subhash C. Narula. Predictive mean square error and stochastic regressor variables. Applied Statistics, 23(1):11– 17, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ali:1974:MON

[304] Zaffar Ali. A model for the occurrence of north winds at kingston, jamaica. Applied Statistics, 23(1):18–21, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pearce:1974:ETM

[305] S. C. Pearce. The estimation of treatment means by designed experiments. Applied Statistics, 23(1):22–25, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Elashoff:1974:TSP

[306] Janet Dixon Elashoff and R. M. Elashoff.

Two-sample problems for a dichotomous variable with missing data. Applied Statistics, 23(1):26–34, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bhansali:1974:AMS

[307] R. J. Bhansali. Asymptotic mean-square error of predicting more than one-step ahead using the regression method. Applied Statistics, 23(1):35–42, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walters:1974:UAM

[308] D. E. Walters. The use of asymptotic models in multiple hormone assays. Applied Statistics, 23(1):43–50, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1974:CVO

[309] D. R. Cox and E. J. Snell. The choice of variables in observational studies. Applied Statistics, 23(1):51–59, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1974:MMA

[310] J. C. Gower and K. V. Mardia. Miscellanea: Multivariate analysis and its applications: A report on the hull conference, 1973. Applied Statistics, 23(1):60– 66, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Miller:1974:MNA

[311] Alan J. Miller. Miscellanea: A note on the analysis of gap-acceptance in traffic. Applied Statistics, 23(1):66–73, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Preece:1974:MIC

[312] D. A. Preece and J. C. Gower. Miscellanea: An iterative computer procedure for mixed-up values in experiments. Applied Statistics, 23(1):73–74, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collings:1974:MQP

[313] T. W. R. Collings. Miscellanea: A queueing problem in which customers have different service distributions. Applied Statistics, 23(1):75–82, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McPherson:1974:SAA

[314] Klim McPherson. Statistical algorithms:

Algorithm AS 67: The evaluation of absorption probabilities in sequential binomial sampling. Applied Statistics, 23(1): 83–86, March 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/67.

Wyshak:1974:SAA

[315] G. Wyshak. Statistical algorithms: Algorithm AS 68: A program for estimating the parameters of the truncated negative binomial distribution. Applied Statistics, 23(1):87–91, March 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/68.

Pike:1974:SAA

[316] M. C. Pike and Diana Bull. Statistical algorithms: Algorithm AS 69: Knox test for space-time clustering in epidemiology. Applied Statistics, 23(1):92– 95, March 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/69.

Odeh:1974:SAA

[317] R. E. Odeh and J. O. Evans. Statistical algorithms: Algorithm AS 70: The percentage points of the normal distribution. Applied Statistics, 23(1):96– 97, March 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/70.

Best:1974:SAA

[318] D. J. Best and P. G. Gipps. Statistical algorithms: Algorithm AS 71: The upper tail probabilities of Kendall’s tau. Applied Statistics, 23(1):98– 100, March 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/71.

Knight:1974:SAR

[319] William Knight. Statistical algorithms: Remark AS R9: A remark on Algorithm AS 37. Applied Statistics, 23 (1):100–101, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/37. See [90].

Sparks:1974:SAC

[320] D. N. Sparks and A. D. Todd. Statistical algorithms: Corrigendum: Algorithm AS 60: Latent roots and vectors of a symmetric matrix. Applied Statistics, 23(1):101–102, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [267].

Anonymous:1974:IRC

[321] Anonymous. Index of remarks and corrigenda on applied statistics algorithms. Applied Statistics, 23(1):102, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:BMPa

[322] Anonymous. Back matter. Applied Statistics, 23(1):[6], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:FMPa

[323] Anonymous. Front matter. Applied Statistics, 23(1):[3], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:VIP

[324] Anonymous. Volume information. Applied Statistics, 23(1):[3], 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pocock:1974:HAA

[325] S. J. Pocock. Harmonic analysis applied to seasonal variations in sickness absence. Applied Statistics, 23(2):103–120, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hocking:1974:PES

[326] R. R. Hocking, H. F. Huddleston, and H. H. Hunt. A procedure for editing survey data. Applied Statistics, 23 (2):121–133, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hamon:1974:SET

[327] B. V. Hamon and E. J. Hannan. Spectral estimation of time delay for dispersive and non-dispersive systems. Applied Statistics, 23(2):134–142, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1974:SAS

[328] G. B. Wetherill and W. K. Chiu. A simplified attribute sampling scheme. Applied Statistics, 23(2):143–148, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1974:ECS

[329] V. Barnett. Economic choice of sample size for sampling inspection plans. Applied Statistics, 23(2):149–157, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1974:SRA

[330] G. E. P. Box, G. M. Jenkins, and J. F. MacGregor. Some recent advances in forecasting and control. Part II. Applied Statistics, 23(2):158–179, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ratkowsky:1974:MLE

[331] D. A. Ratkowsky. Maximum likelihood estimation in small incomplete samples from the bivariate normal distribution. Applied Statistics, 23(2):180–189, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Butcher:1974:ABD

[332] D. F. Butcher and K. E. Kemp. Analysing binomial data in a two-factor experiment with unequal observations per cell. Applied Statistics, 23(2): 190–202, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1974:CCM

[333] A. V. Swan. Corrigendum: Computing maximum-likelihood estimates for parameters of the normal distribution from grouped and censored data. Applied Statistics, 23(2):202, 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [6].

Hill:1974:MAF

[334] I. D. Hill. Miscellanea: Association football and statistical inference. Applied Statistics, 23(2):203–208, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beale:1974:MNC

[335] E. M. L. Beale and P. C. Hutchinson.

Miscellanea: Note on constrained optimum regression. Applied Statistics, 23 (2):208–210, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittaker:1974:MGG

[336] J. Whittaker. Miscellanea: Generating gamma and beta random variables with non-integral shape parameters. Applied Statistics, 23(2):210–214, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hahn:1974:MEO

[337] Gerald J. Hahn and Arthur F. Dershowitz. Miscellanea: Evolutionary operation today — some survey results and observations. Applied Statistics, 23 (2):214–218, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lowe:1974:MEO

[338] C. W. Lowe. Miscellanea: Evolutionary operation in action. Applied Statistics, 23(2):218–226, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fleckenstein:1974:MCE

[339] Fritz Von Fleckenstein. Miscellanea:

Correction of an error in I.B.M. subroutine KRANK the computation of Kendall’s τ. Applied Statistics, 23(2): 226–228, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1974:BRM

[340] D. N. Sparks. Book reviews: A manual of sampling techniques, by R. K. Som. Applied Statistics, 23(2):229, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1974:BRG

[341] A. Springall. Book reviews: The Generalised Jackknife Statistic, by H. L. Gray and W. R. Schucany. Applied Statistics, 23(2):229–230, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1974:BRB

[342] J. N. R. Jeffers. Book reviews: BA-SIC: An introduction to computer programming using the BASIC language, by William F. Sharpe and Nancy L. Jacob. Applied Statistics, 23(2):230–231, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1974:LE

[343] J. A. Nelder and David A. Swick. Letters to the editors. Applied Statistics, 23 (2):232–233, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1974:SAA

[344] Douglas M. Hawkins. Statistical algorithms: Algorithm AS 72: Computing mean vectors and dispersion matrices in multivariate analysis of variance. Applied Statistics, 23(2):234– 238, June 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/72.

Frome:1974:SAA

[345] E. L. Frome. Statistical algorithms:

Algorithm AS 73: Cross-spectrum smoothing via the finite Fourier transform. Applied Statistics, 23(2):238– 244, June 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/73. See remark [861].

Sadovski:1974:SAA

[346] A. N. Sadovski. Statistical algorithms:

Algorithm AS 74: L_{1}-norm fit of a straight line. Applied Statistics, 23(2): 244–248, June 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/74. See remark [498].

Rowlands:1974:SAR

[347] G. J. Rowlands and Rita M. Pocock.

Statistical algorithms: Remark AS R10: On scale selections for scatter diagrams. Applied Statistics, 23(2):248–249, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [10, 95].

Chambers:1974:SAR

[348] John M. Chambers and J. E. Ertel. Statistical algorithms: Remark AS R11: A remark on Algorithm AS 47 “Function Minimization using a Simplex Procedure”. Applied Statistics, 23 (2):250–251, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/47. See [137, 349, 499, 749].

ONeill:1974:SAC

[349] R. O’Neill. Statistical algorithms: Corrigendum: Algorithm AS 47: Function minimization using a simplex procedure. Applied Statistics, 23(2):252, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [137, 348, 499, 749].

Beale:1974:NCO

[350] E. M. L. Beale and P. C. Hutchinson. Note on constrained optimum regression. Applied Statistics, 23(2): 2108–2110, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:BMPb

[351] Anonymous. Back matter. Applied Statistics, 23(2):[6], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:FMPb

[352] Anonymous. Front matter. Applied Statistics, 23(2):[3], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1974:SBN

[353] G. B. Wetherill and J. B. Ofosu. Selection of the best of k normal populations. Applied Statistics, 23(3):253–277, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wood:1974:EAT

[354] J. T. Wood. An extension of the analysis of transformations of Box and Cox. Applied Statistics, 23(3):278–283, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1974:ERM

[355] D. A. H. Thomas. Error rates in multiple comparisons among means-results of a simulation exercise. Applied Statistics, 23(3):284–294, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dyke:1974:DML

[356] G. V. Dyke. Designs to minimize loss of information in polynomial regression. Applied Statistics, 23(3):295–299, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nath:1974:PCB

[357] G. Baikunth Nath. Progressively censored bivariate normal samples. Applied Statistics, 23(3):300–312, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marshall:1974:FTT

[358] Malcomn L. Marshall. Fitting the two-term mixed exponential and twoparameter lognormal distributions to grouped and censored data. Applied Statistics, 23(3):313–322, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1974:LLM

[359] J. A. Nelder. Log linear models for contingency tables: a generalization of classical least squares. Applied Statistics, 23 (3):323–329, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Griffiths:1974:CMI

[360] D. A. Griffiths. A catalytic model of infection for measles. Applied Statistics, 23(3):330–339, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1974:CAN

[361] M. O. Hill. Correspondence analysis: A neglected multivariate method. Applied Statistics, 23(3):340–354, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collins:1974:EMT

[362] Lyndhurst Collins. Estimating Markov transition probabilities from micro-unit data. Applied Statistics, 23(3):355–371, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McKinlay:1974:ENM

[363] Sonja M. McKinlay. The expected number of matches and its variance for matched-pair designs. Applied Statistics, 23(3):372–383, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wierenga:1974:PCP

[364] B. Wierenga. Paired comparison product testing when individual preferences are stochastic. Applied Statistics, 23 (3):384–396, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See comment [792].

Anderson:1974:DLD

[365] J. A. Anderson. Diagnosis by logistic discriminant function: further practical problems and results. Applied Statistics, 23(3):397–404, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1974:ISS

[366] Morton B. Brown. Identification of the sources of significance in two-way contingency tables. Applied Statistics, 23 (3):405–413, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brook:1974:BLG

[367] D. Brook and G. J. G. Upton. Biases in local government elections due to position on the ballot paper. Applied Statistics, 23(3):414–419, 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chiu:1974:EDC

[368] W. K. Chiu. The economic design of cusum charts for controlling normal means. Applied Statistics, 23(3): 420–433, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1974:MJC

[369] D. H. Ward. Miscellanea: Joint Conference at Warwick — a summary report. Applied Statistics, 23(3):434–435, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daley:1974:MCB

[370] D. J. Daley. Miscellanea: Computation of bi- and tri-variate normal integrals. Applied Statistics, 23(3):435–438, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1974:BRE

[371] C. Chatfield. Book reviews: Elements of Statistics, by D. R. Byrkit. Applied Statistics, 23(3):439, 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinkley:1974:BRB

[372] David Hinkley. Book reviews: A bibliography of multivariate statistical analysis, by T. W. Anderson, Somesh Das Gupta and George P. H. Styan. Applied Statistics, 23(3):439–440, 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1974:BRI

[373] G. M. Clarke. Book reviews: Introduction to Biostatistics, by R. R. Sokal and F. J. Rohlf. Applied Statistics, 23 (3):440–441, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coleman:1974:BRG

[374] Rodney Coleman. Book reviews: A guide to probability theory and application, by Cyrus Derman, Leon J. Gleser and Ingram Olkin. Applied Statistics, 23(3):441, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harris:1974:BRW

[375] R. Harris. Book reviews: OR in World War 2: Operational Research Against the U-Boat, by C. H. Waddington. Applied Statistics, 23(3):441–442, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1974:BRO

[376] John Luckman. Book reviews: Operations Research in Hospitals-Diagnosis and Prognosis, by David H. Stimson and Ruth H. Stimson. Applied Statistics, 23 (3):442–443, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Veevers:1974:BRB

[377] A. Veevers. Book reviews: Bibliography on Experiment and Treatment Design Pre-1968, by Walter T. Federer and Leslie N. Balaam. Applied Statistics, 23 (3):443–444, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1974:BRM

[378] John Luckman. Book reviews: Modern Management Methods and the Organisation of the Health Services. Applied Statistics, 23(3):444–445, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Luckman:1974:BRH

[379] John Luckman. Book reviews: Health Practice Research and Formalised Managerial Methods, by F. Grundy and W. A. Reinke. Applied Statistics, 23 (3):445–446, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1974:BRS

[380] J. F. Scott. Book reviews: Statistics, by Roger A. Carlson. Applied Statistics, 23 (3):446–447, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1974:SAA

[381] W. Morven Gentleman. Statistical algorithms: Algorithm AS 75: Basic procedures for large, sparse or weighted linear least problems. Applied Statistics, 23(3):

448–454, September 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/75. See remark

[558].

Young:1974:SAA

[382] J. C. Young and Ch. E. Minder. Statistical algorithms: Algorithm AS 76: An integral useful in calculating noncentral t and bivariate normal probabilities. Applied Statistics, 23(3):455– 457, September 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/76. See remarks

[747, 782, 874, 878, 883].

Venables:1974:SAA

[383] W. N. Venables. Statistical algorithms: Algorithm AS 77: Null distribution of the largest root statistic. Applied Statistics, 23(3):458–465, September 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/77.

Gower:1974:SAA

[384] J. C. Gower. Statistical algorithms: Algorithm AS 78: The mediancentre. Applied Statistics, 23(3):466–470, September 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/78.

Farebrother:1974:SAA

[385] R. W. Farebrother. Statistical algorithms: Algorithm AS 79: Gram– Schmidt regression. Applied Statistics, 23(3):470–476, September 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/79.

Farebrother:1974:SAR

[386] R. W. Farebrother and G. Berry. Statistical algorithms: Remark AS R12: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 23(3):477, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/6. See [3, 8,

748].

Anonymous:1974:SAC

[387] Anonymous. Statistical algorithms:

Computational problems in statistics. Applied Statistics, 23(3):478, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:BMPc

[388] Anonymous. Back matter. Applied Statistics, 23(3):[6], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1974:FMPc

[389] Anonymous. Front matter. Applied Statistics, 23(3):[3], 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindsey:1975:LAT

[390] J. K. Lindsey. Likelihood analyses and tests for binary data. Applied Statistics, 24(1):1–16, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hart:1975:SAA

[391] R. A. Hart, J. Hutton, and T. Sharot.

A statistical analysis of association football attendances. Applied Statistics, 24 (1):17–27, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pethybridge:1975:MLE

[392] R. J. Pethybridge. Maximum likelihood estimation of a linear regression function with grouped data. Applied Statistics, 24(1):28–41, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1975:CMF

[393] J. Taylor. Correction: A method of fitting several linear functional relations and of testing for differences between them. Applied Statistics, 24(1):41, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1975:RAS

[394] J. R. Scott. A regression assessment statistic. Applied Statistics, 24(1):42–45, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1975:AII

[395] G. H. Freeman. Analysis of interactions in incomplete two-way tables. Applied Statistics, 24(1):46–55, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1975:CVT

[396] J. A. John and P. Prescott. Critical values of a test to detect outliers in factorial experiments. Applied Statistics, 24(1):56–59, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pearce:1975:RD

[397] S. C. Pearce. Row-and-column designs. Applied Statistics, 24(1):60–74, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Osborn:1975:MMA

[398] John Osborn. A multiplicative model for the analysis of vital statistics rates. Applied Statistics, 24(1):75–84, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1975:LMI

[399] Martin Crowder and Paul R. Grob. A logit model for infectious diseases. Applied Statistics, 24(1):85–94, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1975:PPM

[400] V. Barnett. Probability plotting methods and order statistics. Applied Statistics, 24(1):95–108, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ratkowsky:1975:TSL

[401] D. A. Ratkowsky and G. R. Dolby.

Taylor series linearization and scoring for parameters in nonlinear regression. Applied Statistics, 24(1):109–111, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McKenzie:1975:CGL

[402] J. McKenzie. A comparison of Gaussian and logistic categorical opinion distribution models. Applied Statistics, 24 (1):112–122, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Worsdale:1975:TCD

[403] Graham J. Worsdale. Tables of cumulative distribution functions for symmetric stable distributions. Applied Statistics, 24(1):123–131, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chay:1975:MUB

[404] S. C. Chay, R. D. Fardo, and M. Mazumdar. Miscellanea: On using the Box-Muller transformation with multiplicative congruential pseudo-random number generators. Applied Statistics, 24 (1):132–135, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bebbington:1975:MSM

[405] A. C. Bebbington. Miscellanea: A simple method of drawing a sample without replacement. Applied Statistics, 24(1):136, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1975:MSS

[406] R. W. Morgan. Miscellanea: Some stochastic models to describe the fertilization of an egg. Applied Statistics, 24 (1):137–138, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1975:BRB

[407] J. C. Gower. Book reviews: Biometrical Interpretation, by N. E. Gilbert. Applied Statistics, 24(1):139, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gower:1975:BRM

[408] J. C. Gower. Book reviews: Méthodes et Programmes d’Analyse Discriminante, by Jean-Marie Romeder. Applied Statistics, 24(1):139–140, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jephcott:1975:BRS

[409] J. St G. Jephcott. Book reviews: Statistical Thinking: A Structural Approach, by J. L. Phillips. Applied Statistics, 24 (1):140–141, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jephcott:1975:BRM

[410] J. St G. Jephcott. Book reviews: Market Segmentation, by R. E. Frank, W. F. Massey and Y. Wind. Applied Statistics, 24(1):141–142, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1975:BRE

[411] R. J. Nicholson. Book reviews: Exeter — University and City, by F. M. M. Lewes and Anne Kirkness. Applied Statistics, 24(1):142, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cormack:1975:BRS

[412] R. M. Cormack. Book reviews: Spatial Autocorrelation, by A. D. Cliff and J. K. Ord. Applied Statistics, 24(1):142–143, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mardia:1975:SAAa

[413] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 80:

Spherical statistics. Applied Statistics, 24(1):144–146, March 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/80.

Mardia:1975:SAAb

[414] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 81: Circular statistics. Applied Statistics, 24 (1):147–150, March 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/81.

Gower:1975:SAA

[415] J. C. Gower. Statistical algorithms: Algorithm AS 82: The determinant of an orthogonal matrix. Applied Statistics, 24 (1):150–153, March 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/82.

Monro:1975:SAA

[416] Donald M. Monro. Statistical algorithms: Algorithm AS 83: Complex discrete fast Fourier transform. Applied Statistics, 24(1):153–160, March 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/83.

Sparks:1975:SAR

[417] D. N. Sparks. Statistical algorithms: Remark AS R13: A remark on Algorithm AS 58: Euclidean cluster analysis. Applied Statistics, 24(1): 160–161, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/58. See [244].

Anonymous:1975:BMPa

[418] Anonymous. Back matter. Applied Statistics, 24(1):[6], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:FMPa

[419] Anonymous. Front matter. Applied Statistics, 24(1):[7], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:VIP

[420] Anonymous. Volume information. Applied Statistics, 24(1):[3], 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mardia:1975:AMR

[421] K. V. Mardia. Assessment of multinormality and the robustness of Hotelling’s T^{2} test. Applied Statistics, 24(2): 163–171, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Littlewood:1975:RMS

[422] B. Littlewood. A reliability model for systems with Markov structure. Applied Statistics, 24(2):172–177, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kass:1975:STA

[423] G. V. Kass. Significance testing in automatic interaction detection (A.I.D.). Applied Statistics, 24(2):178–189, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1975:EMV

[424] J. A. John and P. Prescott. Estimating missing values in experiments. Applied Statistics, 24(2):190–192, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Plackett:1975:AP

[425] R. L. Plackett. The analysis of permutations. Applied Statistics, 24(2):193–202, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bather:1975:NPT

[426] J. A. Bather. A non-parametric test for Hertzian fracture data. Applied Statistics, 24(2):203–209, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1975:SAS

[427] P. Prescott. A simple alternative to Student’s t. Applied Statistics, 24(2): 210–217, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1975:EED

[428] D. V. Lindley. The effect of ethical design considerations on statistical analysis. Applied Statistics, 24(2):218–228, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jacobsen:1975:QRC

[429] M. Jacobsen. Quantifying radiological changes in simple pneumoconiosis. Applied Statistics, 24(2):229–249, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fabrycy:1975:MMC

[430] Mark Z. Fabrycy. Miscellanea: Multicollinearity caused by specification errors. Applied Statistics, 24(2):250–254, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See comments [606].

McLeod:1975:MDT

[431] Ian McLeod. Miscellanea: Derivation of the theoretical autocovariance function of autoregressive moving average time series. Applied Statistics, 24(2):255–256, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kerridge:1975:MIR

[432] D. Kerridge. Miscellanea: The interpretation of rank correlations. Applied Statistics, 24(2):257–258, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1975:AWP

[433] J. A. Nelder. Announcement by the working party on statistical computing: GLIM (generalized linear interactive modelling program). Applied Statistics, 24(2):259–261, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mardia:1975:SAAc

[434] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 84: Measures of multivariate skewness and kurtosis. Applied Statistics, 24(2):262– 265, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/84.

Bukac:1975:SAA

[435] J. Bukač. Statistical algorithms: Algorithm AS 85: Critical values of the sign test. Applied Statistics, 24(2):265– 267, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/85.

Mardia:1975:SAAd

[436] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 86: The von Mises distribution function. Applied Statistics, 24(2):268– 272, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/86.

Martinson:1975:SAA

[437] E. O. Martinson and M. A. Hamdan. Statistical algorithms: Algorithm AS 87: Calculation of the polychoric estimate of correlation in contingency tables. Applied Statistics, 24(2):272– 278, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/87. See remark [592].

Anonymous:1975:BMPb

[438] Anonymous. Back matter. Applied Statistics, 24(2):[6], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:FMPb

[439] Anonymous. Front matter. Applied Statistics, 24(2):[3], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Upton:1975:DOP

[440] G. J. G. Upton and D. Brook. The determination of the optimum position on a ballot paper. Applied Statistics, 24 (3):279–287, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1975:EIE

[441] Morton B. Brown. Exploring interaction effects in the ANOVA. Applied Statistics, 24(3):288–298, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hegazy:1975:SNG

[442] Y. A. S. Hegazy and J. R. Green.

Some new goodness-of-fit tests using order statistics. Applied Statistics, 24 (3):299–308, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hart:1975:CSA

[443] R. A. Hart, J. Hutton, and T. Sharot.

Correction: A statistical analysis of association football attendances. Applied Statistics, 24(3):308, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rodda:1975:OCO

[444] B. E. Rodda, C. B. Sampson, and D. W. Smith. The one-compartment open model: some statistical aspects of parameter estimation. Applied Statistics, 24(3):309–318, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carter:1975:EMB

[445] Michael C. Carter and William O. Williford. Estimation in a modified binomial distribution. Applied Statistics, 24 (3):319–328, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robinson:1975:DAC

[446] D. Robinson. Divination in ancient china. Applied Statistics, 24(3):329–332, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chan:1975:SSE

[447] W.-Y. T. Chan and H. Tong. A simulation study of the estimation of evolutionary spectral functions. Applied Statistics, 24(3):333–341, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heeler:1975:SNU

[448] Roger M. Heeler and George S. Day. A supplementary note on the use of cluster analysis for stratification. Applied Statistics, 24(3):342–344, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yeomans:1975:FOS

[449] K. A. Yeomans and P. A. Golder. Further observations on the stratification of birmingham wards by clustering: A riposte. Applied Statistics, 24(3):345–346, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McCormick:1975:MNE

[450] G. F. McCormick and V. A. Sposito. Miscellanea: A note on L_{1} estimation based on the median positive quotient.

Applied Statistics, 24(3):347–350, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bryce:1975:MMA

[451] G. Rex Bryce. Miscellanea: MAD: an analysis of variance program for unbalanced designs. Applied Statistics, 24 (3):350–352, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Balmer:1975:MOD

[452] D. W. Balmer. Miscellanea: An observation on the disaggregation of annual time series data. Applied Statistics, 24 (3):352–354, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chanter:1975:MMA

[453] D. O. Chanter. Miscellanea: Modifications of the angular transformation. Applied Statistics, 24(3):354–359, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1975:BRT

[454] I. D. Hill. Book reviews: Tables for Bayesian Statisticians, by Gerald L. Issacs, David E. Christ, Melvin R. Novick and Paul H. Jackson. Applied Statistics, 24(3):360–361, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hiorns:1975:BRD

[455] R. W. Hiorns. Book reviews: Demographic Data Collection: A Summary of Experience, by William Seltzer. Applied Statistics, 24(3):361, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1975:BRT

[456] M. J. R. Healy. Book reviews: Time-Series, by M. G. Kendall. Applied Statistics, 24(3):361–362, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paige:1975:BRU

[457] W. J. Paige. Book reviews: Uncertainty, Calibration and Probability, by C. F. Dietrich. Applied Statistics, 24(3):362–363, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashford:1975:BRM

[458] J. R. Ashford. Book reviews: An MMPI Source Book. Basic Item, Scale and Pattern Data on 50,000 Medical Patients, by Wendell M. Swenson, John S. Pearson and David Osborne. Applied Statistics, 24(3):363, 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thewlis:1975:BRS

[459] D. J. Thewlis. Book reviews: Systems Analysis and Design Using Network Techniques, by Gary E. Whitehouse. Applied Statistics, 24(3):363–364, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wetherill:1975:BRQ

[460] G. B. Wetherill. Book reviews: Quality Control and Industrial Statistics, by Acheson J. Duncan. Applied Statistics, 24(3):364–365, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Computing:1975:CDA

[461] The Working Party on Statistical Computing. The construction and description of algorithms. Applied Statistics, 24 (3):366–373, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1975:SAA

[462] Jane F. Gentleman. Statistical algorithms: Algorithm AS 88: Generation of all _{N}C_{R} combinations by simulating nested Fortran DO loops. Applied Statistics, 24(3):374–376, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/88.

Best:1975:SAAa

[463] D. J. Best and D. E. Roberts. Statistical algorithms: Algorithm AS 89: The upper tail probabilities of Spearman’s rho. Applied Statistics, 24(3):377–379, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/89.

Bohrer:1975:SAA

[464] Robert Bohrer. Statistical algorithms:

Algorithm AS 90: One-sided multivariable inference. Applied Statistics, 24(3):380–384, September 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/90.

Best:1975:SAAb

[465] D. J. Best and D. E. Roberts. Statistical algorithms: Algorithm AS 91: The percentage points of the χ^{2} distribution. Applied Statistics, 24(3):385–388, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/91.

Brooker:1975:SAA

[466] P. Brooker and M. J. P. Selby. Statistical algorithms: Algorithm AS 92: The sample size for a distribution-free tolerance interval. Applied Statistics, 24(3): 388–390, September 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/92.

Gower:1975:SAC

[467] J. C. Gower. Statistical algorithms: Correction: AS 78: The mediancentre. Applied Statistics, 24(3):390, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:SAL

[468] Anonymous. Statistical algorithms: List of applied statistics algorithms. Applied Statistics, 24(3):391–394, 1975. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:BMPc

[469] Anonymous. Back matter. Applied Statistics, 24(3):[2], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1975:FMPc

[470] Anonymous. Front matter. Applied Statistics, 24(3):[3], 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1976:ESS

[471] G. H. Brown. The efficiency of stratified sampling in the estimation of parameters for a bivariate normal population. Applied Statistics, 25(1):1–7, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Selvin:1976:GEP

[472] Steve Selvin. A graphical estimate of the population mean from censored normal data. Applied Statistics, 25(1):8–11, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Golder:1976:BMM

[473] E. R. Golder and J. G. Settle. The Box-Muller method for generating pseudorandom normal deviates. Applied Statistics, 25(1):12–20, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hutchinson:1976:CTT

[474] T. P. Hutchinson. Combining twotailed rank-correlation statistics. Applied Statistics, 25(1):21–25, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Charlton:1976:CL

[475] P. Charlton and A. S. C. Ehrenberg.

Customers of the LEP. Applied Statistics, 25(1):26–30, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1976:MPS

[476] B. J. T. Morgan. Markov properties of sequences of behaviours. Applied Statistics, 25(1):31–36, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1976:SEM

[477] Morton B. Brown. Screening effects in multidimensional contingency tables. Applied Statistics, 25(1):37–46, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1976:CPP

[478] V. Barnett. Convenient probability plotting positions for the normal distribution. Applied Statistics, 25(1):47–50, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1976:PEP

[479] Douglas M. Hawkins. Point estimation of the parameters of piecewise regression models. Applied Statistics, 25(1):51–57, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Steinijans:1976:MSP

[480] Volker W. Steinijans. Miscellanea: A stochastic point-process model for the occurrence of major freezes in lake constance. Applied Statistics, 25(1):58–61, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1976:MMV

[481] J. A. John and S. M. Lewis. Miscellanea: Mixed-up values in experiments. Applied Statistics, 25(1):61–63, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1976:MSD

[482] A. A. Greenfield. Miscellanea: Selection of defining contrasts in two-level experiments. Applied Statistics, 25(1):64–67, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farebrother:1976:LE

[483] R. W. Farebrother. Letter to the editors.

Applied Statistics, 25(1):68, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fletcher:1976:BRI

[484] R. Fletcher. Book reviews: Introduction to Optimization Methods, by P. R. Adby and M. A. H. Dempster. Applied Statistics, 25(1):69, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1976:BRS

[485] L. H. C. Tippett. Book reviews: Statistical Handbook for Non-Statisticians, by Ray Meddis. Applied Statistics, 25(1): 69, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sibson:1976:BRC

[486] Robin Sibson. Book reviews: Clustering Algorithms, by J. A. Hartigan. Applied Statistics, 25(1):70, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartholomew:1976:BRM

[487] D. J. Bartholomew. Book reviews: Measurement in the Social Sciences: Theories and Strategies, by H. M. Blalock. Applied Statistics, 25(1):70–71, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1976:BRD

[488] R. W. Morgan. Book reviews: Data Analysis for Scientists and Engineers, by Stuart L. Meyer. Applied Statistics, 25(1):71, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1976:BRS

[489] Graham Watts. Book reviews: Statistical Analysis for Business Decisions, by William A. Spurr and Charles P. Bonini. Applied Statistics, 25(1):71–72, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pollock:1976:BRB

[490] K. H. Pollock. Book reviews: Biostatistics: A Foundation for Analysis in the Health Sciences, by Wayne W. Daniel. Applied Statistics, 25(1):72, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1976:BRE

[491] A. S. C. Ehrenberg. Book reviews: Essentials of Statistics in Marketing, by C. S. Greensted, A. K. S. Jardine and J. D. MacFarlane. Applied Statistics, 25(1): 73, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1976:BRD

[492] J. Taylor. Book reviews: Design of Experiments: A Realistic Approach, by Virgil L. Anderson and Robert A. McLean. Applied Statistics, 25(1):73–74, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1976:BRA

[493] C. Chatfield. Book reviews: Applied Statistics, by R. P. Gupta. Applied Statistics, 25(1):74, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinneen:1976:SAA

[494] L. C. Dinneen and B. C. Blakesley. Statistical algorithms: Algorithm AS 93: A generator for the null distribution of the Ansari-Bradley W statistic. Applied Statistics, 25(1):75–81, March 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/93.

McLaren:1976:SAA

[495] Mara L. McLaren. Statistical algorithms: Algorithm AS 94: Coefficients of the zonal polynomials. Applied Statistics, 25(1):82–87, March 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/94.

Benn:1976:SAA

[496] R. T. Benn and Susan Sidebottom. Statistical algorithms: Algorithm AS 95: Maximum-likelihood estimation of location and scale parameters from grouped data. Applied Statistics, 25(1):88– 93, March 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/95.

Nelder:1976:SAA

[497] J. A. Nelder. Statistical algorithms: Algorithm AS 96: A simple algorithm for scaling graphs. Applied Statistics, 25(1): 94–96, March 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/96.

Sposito:1976:SAR

[498] V. A. Sposito. Statistical algorithms:

Remark AS R14: A remark on Algorithm AS 74: “L_{1}-Norm Fit of a Straight Line”. Applied Statistics, 25 (1):96–97, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/74. See [346].

Benyon:1976:SAR

[499] P. R. Benyon. Statistical algorithms: Remark AS R15: Function minimization using a simplex procedure. Applied Statistics, 25(1):97, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [137, 348, 349].

Anonymous:1976:SARa

[500] Anonymous. Statistical algorithms: Recent statistical algorithms. Applied Statistics, 25(1):98–101, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:BMPa

[501] Anonymous. Back matter. Applied Statistics, 25(1):[4], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:FMPa

[502] Anonymous. Front matter. Applied Statistics, 25(1):[5], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:VIP

[503] Anonymous. Volume information. Applied Statistics, 25(1):[3], 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:1976:ATU

[504] A. J. Scott and M. Knott. An approximate test for use with AID. Applied Statistics, 25(2):103–106, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brubacher:1976:ITS

[505] S. R. Brubacher and G. Tunnicliffe Wilson. Interpolating time series with application to the estimation of holiday effects on electricity demand. Applied Statistics, 25(2):107–116, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalmijn:1976:SCR

[506] W. M. Kalmijn. Simultaneous confidence regions for repeat preference testing. Applied Statistics, 25(2):117–122, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yamamoto:1976:AMS

[507] Taku Yamamoto. Asymptotic mean square prediction error for an autoregressive model with estimated coefficients. Applied Statistics, 25(2):123–127, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawless:1976:CIE

[508] J. F. Lawless. Confidence interval estimation in the inverse power law model. Applied Statistics, 25(2):128–138, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1976:ALR

[509] J. N. R. Jeffers, Doreen M. Howard, and P. J. A. Howard. An analysis of litter respiration at different temperatures. Applied Statistics, 25(2):139–146, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenfeld:1976:ITS

[510] Gerald Rosenfeld. Identification of time series with infinite variance. Applied Statistics, 25(2):147–153, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sposito:1976:MSC

[511] V. A. Sposito and W. C. Smith. Miscellanea: On a sufficient condition and a necessary condition for L_{1} estimation. Applied Statistics, 25(2):154–157, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dolby:1976:MCB

[512] G. R. Dolby. Miscellanea: The connection between methods of estimation in implicit and explicit nonlinear models. Applied Statistics, 25(2):157–162, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1976:BRR

[513] Derek Bissell. Book reviews: Reading and Understanding Applied Statistics: A Self-Learning Approach, by Sidney M. Stahl and James D. Hennes. Applied Statistics, 25(2):163, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jeffers:1976:BRS

[514] J. N. R. Jeffers. Book reviews: Statistical and Mathematical Aspects of Pollution Problems, by John W. Pratt. Applied Statistics, 25(2):164, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1976:BRS

[515] J. Taylor. Book reviews: Statistical Methods in Animal Sciences, by V. N. Amble. Applied Statistics, 25 (2):164–165, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sharot:1976:BRS

[516] T. Sharot. Book reviews: Statistics and Econometrics: A Problem-Solving Text, by Barry R. Chiswick and Stephen J. Chiswick. Applied Statistics, 25(2):165, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Monro:1976:SAA

[517] Donald M. Monro. Statistical algorithms: Algorithm AS 97: Real discrete fast Fourier transform. Applied Statistics, 25(2):166–172, June 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/97.

Golder:1976:SAA

[518] E. R. Golder. Statistical algorithms: Algorithm AS 98: The spectral test for the evaluation of congruential pseudorandom generators. Applied Statistics, 25(2):173–180, June 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/98. See remark

[559, 745].

Hill:1976:SAAa

[519] I. D. Hill, R. Hill, and R. L. Holder.

Statistical algorithms: Algorithm AS 99: Fitting Johnson curves by moments. Applied Statistics, 25(2):180– 189, June 1976. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/99. See remark [856].

Hill:1976:SAAb

[520] I. D. Hill. Statistical algorithms: Algorithm AS 100: Normal-Johnson and Johnson-normal transformations. Applied Statistics, 25(2):190–192, June 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/100. See remark [856, 869].

Haberman:1976:SAC

[521] S. J. Haberman. Statistical algorithms: Correction to AS 51: Log-linear fit for contingency tables. Applied Statistics, 25(2):193, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wedderburn:1976:SAR

[522] R. W. M. Wedderburn. Statistical algorithms: Remark AS R16: A remark on Algorithm AS 29: “The Runs Up and Down Test”. Applied Statistics, 25(2):193, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/29. See [50].

Anonymous:1976:BMPb

[523] Anonymous. Back matter. Applied Statistics, 25(2):[2], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:FMPb

[524] Anonymous. Front matter. Applied Statistics, 25(2):[3], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1976:CFA

[525] G. E. P. Box and G. C. Tiao. Comparison of forecast and actuality. Applied Statistics, 25(3):195–200, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reilly:1976:NCP

[526] Park M. Reilly. The numerical computation of posterior distributions in Bayesian statistical inference. Applied Statistics, 25(3):201–209, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nash:1976:PCR

[527] J. C. Nash and L. P. Lefkovitch. Principal components and regression by singular value decomposition on a small computer. Applied Statistics, 25(3):210–216, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilchrist:1976:UVC

[528] Robert Gilchrist. The use of verbal-call systems to call patients from waiting areas to consulting rooms. Applied Statistics, 25(3):217–227, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collett:1976:SNO

[529] D. Collett and T. Lewis. The subjective nature of outlier rejection procedures. Applied Statistics, 25(3):228–237, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Titterington:1976:UDS

[530] D. M. Titterington. Updating a diagnostic system using unconfirmed cases. Applied Statistics, 25(3):238–247, 1976.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wichern:1976:CVF

[531] Dean W. Wichern, Robert B. Miller, and Der-Ann Hsu. Changes of variance in first-order autoregressive time series models — with an application. Applied Statistics, 25(3):248–256, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robert:1976:UTL

[532] P. Robert and Y. Escoufier. A unifying tool for linear multivariate statistical methods: The RV- coefficient. Applied Statistics, 25(3):257–265, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simon:1976:CSS

[533] Gary Simon. Computer simulation swindles, with applications to estimates of location and dispersion. Applied Statistics, 25(3):266–274, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Desu:1976:SSD

[534] M. M. Desu, Subhash C. Narula, and Bernardo Villarreal. Sample size determination for estimation of guarantee times. Applied Statistics, 25(3):275–279, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leger:1976:CTT

[535] A. S. St Leger. Comparison of two tests for seasonality in epidemiological data. Applied Statistics, 25(3):280–286, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burch:1976:MSS

[536] C. R. Burch and I. T. Parsons. Miscellanea: “squeeze” significance tests. Applied Statistics, 25(3):287–291, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lozy:1976:MEE

[537] Mohamed El Lozy. Miscellanea: An extension of Ehrenberg’s relations between mean weights and heights. Applied Statistics, 25(3):291–295, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Darby:1976:MTS

[538] S. C. Darby and M. J. Ellis. Miscellanea: A test for synergism between two drugs. Applied Statistics, 25(3):296–299, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Skegg:1976:BRI

[539] D. C. G. Skegg. Book reviews: An Introduction to Epidemiology, by Michael Alderson. Applied Statistics, 25(3):300, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stern:1976:BRS

[540] G. J. A. Stern. Book reviews: Statistical Methods for Engineers and Scientists, by Robert M. Bethea, Benjamin S. Duran and Thomas L. Boullion. Applied Statistics, 25(3):300–301, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Patterson:1976:BRD

[541] H. D. Patterson. Book reviews: Developments in Field Experiment Design and Analysis, by V. J. Bofinger and J. L.

Wheeler. Applied Statistics, 25(3):301, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1976:BRA

[542] M. J. R. Healy. Book reviews: The Analysis of Time Series: Theory and Practice, by C. Chatfield. Applied Statistics, 25(3):301–302, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1976:BRD

[543] D. N. Sparks. Book reviews: Data Reduction-Analysis and Interpreting Statistical Data, by A. S. C. Ehrenberg. Applied Statistics, 25(3):302–303, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1976:BRS

[544] P. N. Lee. Book reviews: Survival Distributions: Reliability Applications in the Biomedical Sciences, by Alan J. Gross and Virginia A. Clark. Applied Statistics, 25(3):303, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fox:1976:BRF

[545] W. Fox. Book reviews: Fundamental Statistics for Psychology: First Course for Students of Psychology and Education, by R. B. McCall. Applied Statistics, 25(3):304, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maher:1976:BRB

[546] Michael J. Maher. Book reviews: Basic Operational Research. 2 Edition, by P. G. Moore. Applied Statistics, 25 (3):304–305, 1976. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Horsfield:1976:BRA

[547] S. Horsfield. Book reviews: Applied Multivariate Analysis and Experimental Designs, by N. Krishnan Namboodiri, L. F. Carter and H. M. Blalock, *Jr. Applied Statistics, 25(3):305–306, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brigden:1976:BRE

[548] M. E. Brigden. Book reviews: Elementary Dynamic Programming, by John M. Norman. Applied Statistics, 25(3):306, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kempton:1976:BRE

[549] R. A. Kempton. Book reviews: Ecological Diversity, by E. C. Pielou. Applied Statistics, 25(3):306–307, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1976:BRS

[550] A. Springall. Book reviews: Sample Size Choice: Charts for Experiments with Linear Models, by R. E. Odeh and M. Fox. Applied Statistics, 25(3):307–308, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:BRS

[551] Anonymous. Book reviews: Statistical Inference, by S. D. Silvey. Applied Statistics, 25(3):308, 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1976:BRP

[552] D. N. Sparks. Book reviews: Principal Sources of Marketing Information, by Christine Hull. Applied Statistics, 25(3):308, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1976:BRE

[553] G. Watts. Book reviews: Econometrics and Economic Theory-Essays in Honour of Jan Tinbergen, by Willy Sellekaerts. Applied Statistics, 25(3):308, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Narula:1976:SAA

[554] Subhash C. Narula and M. M. Desu. Statistical algorithms: Algorithm AS 101: Distribution-free confidence intervals. Applied Statistics, 25(3):309–312, September 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/101.

Banfield:1976:SAA

[555] C. F. Banfield. Statistical algorithms: Algorithm AS 102: Ultrametric distances for a single linkage dendrogram. Applied Statistics, 25(3):313–315, September 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/102.

Bernardo:1976:SAA

[556] J. M. Bernardo. Statistical algorithms: Algorithm AS 103: Psi (digamma) function. Applied Statistics, 25(3):315–317, September 1976. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/103.

Farebrother:1976:SAA

[557] R. W. Farebrother. Statistical algorithms: Algorithm AS 104: BLUS residuals. Applied Statistics, 25(3):317–322, September 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/104.

Farebrother:1976:SAR

[558] R. W. Farebrother. Statistical algorithms: Remark AS R17: Recursive residuals — A remark on Algorithm AS 75: Basic procedures for large, sparse or weighted linear least squares problems. Applied Statistics, 25(3): 323–324, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/75. See [381].

Golder:1976:SAR

[559] E. R. Golder. Statistical algorithms:

Remark AS R18: The spectral test for the evaluation of congruential pseudorandom generators. Applied Statistics, 25(3):324, 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [518, 745].

Anonymous:1976:SARb

[560] Anonymous. Statistical algorithms: Recent statistical algorithms. Applied Statistics, 25(3):325–327, 1976. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:BMPc

[561] Anonymous. Back matter. Applied Statistics, 25(3):[2], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1976:FMPc

[562] Anonymous. Front matter. Applied Statistics, 25(3):[3], 1976. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Machin:1977:CML

[563] David Machin and Barry Sanderson. Computing maximum-likelihood estimates for the parameters of the de Wit competition model. Applied Statistics, 26(1):1–8, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nash:1977:DAL

[564] J. C. Nash. A discrete alternative to the logistic growth function. Applied Statistics, 26(1):9–14, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitchison:1977:CCT

[565] J. Aitchison, J. D. F. Habbema, and J. W. Kay. A critical comparison of two methods of statistical discrimination. Applied Statistics, 26(1):15–25, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kilpatrick:1977:ESD

[566] S. James Kilpatrick, Jr. An empirical study of the distribution of episodes of illness recorded in the 1970–71 national morbidity survey. Applied Statistics, 26(1):26–33, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mansfield:1977:AVS

[567] E. R. Mansfield, J. T. Webster, and R. F. Gunst. An analytic variable selection technique for principal component regression. Applied Statistics, 26(1):34– 40, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Murphy:1977:RSP

[568] Allan H. Murphy and Robert L. Winkler. Reliability of subjective probability forecasts of precipitation and temperature. Applied Statistics, 26(1):41–47, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gudmundsson:1977:MAE

[569] G. Gudmundsson. Multivariate analysis of economic variables. Applied Statistics, 26(1):48–59, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rubin:1977:CPD

[570] Donald B. Rubin and T. W. F. Stroud.

The calculation of the posterior distribution of the cell means in a two-way unbalanced MANOVA. Applied Statistics, 26(1):60–66, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinkley:1977:MQC

[571] David Hinkley. Miscellanea: On quick choice of power transformation. Applied Statistics, 26(1):67–69, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Patefield:1977:MIM

[572] W. M. Patefield. Miscellanea: On the information matrix in the linear functional relationship problem. Applied Statistics, 26(1):69–70, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheng:1977:MGG

[573] R. C. H. Cheng. Miscellanea: The generation of gamma variables with non-integral shape parameter. Applied Statistics, 26(1):71–75, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1977:MAC

[574] E. Page. Miscellanea: Approximations to the cumulative normal function and its inverse for use on a pocket calculator. Applied Statistics, 26(1):75–76, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1977:MRC

[575] O. D. Anderson. Miscellanea: A report on the 1976 Cambridge Forecasting Conference. Applied Statistics, 26(1):76–79, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1977:BRA

[576] A. S. C. Ehrenberg. Book reviews: Assessment of Skeletal Maturity and Prediction of Adult Height (TW2 Method), by J. M. Tanner, R. H. Whitehouse, W. A. Marshall, M. J. R. Healy and H. Goldstein. Applied Statistics, 26(1):80, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1977:BRI

[577] P. N. Lee. Book reviews: Interpretation and Uses of Medical Statistics, by G. J. Bourke and J. McGilvray. Applied Statistics, 26(1):81, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mead:1977:BRB

[578] R. Mead. Book reviews: Biomathematics: An Introduction, by David Machin. Applied Statistics, 26(1):81–82, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bowen:1977:BRD

[579] K. C. Bowen. Book reviews: Decision Methodology: A Formalization of the OR Process, by D. J. White. Applied Statistics, 26(1):82, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Banfield:1977:BRG

[580] C. F. Banfield. Book reviews: Geological Factor Analysis (Methods in Geomathematics 1), by K. G. Joreskog, J. E. Klovan and R. A. Reyment. Applied Statistics, 26(1):82–83, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wanhill:1977:BRA

[581] Stephen Wanhill. Book reviews: Applied Econometrics, by Meghnad Desai. Applied Statistics, 26(1):83–84, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1977:BRS

[582] I. D. Hill. Book reviews: Significance Tests, by Evelyn Caulcott. Applied Statistics, 26(1):84–85, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McPherson:1977:BRS

[583] Klim McPherson. Book reviews: Sequential Methods in Statistics. 2nd Edition, by G. Barrie Wetherill. Applied Statistics, 26(1):85, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Radcliffe:1977:BRM

[584] J. Radcliffe. Book reviews: The Mathematical Theory of Infectious Diseases and its Applications. 2nd Edition, by Norman T. J. Bailey. Applied Statistics, 26(1):85–87, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wermuth:1977:SAA

[585] Nanny Wermuth and Eberhard Scheidt. Statistical algorithms: Algorithm AS 105: Fitting a covariance selection model to a matrix. Applied Statistics, 26(1):88–92, March 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/105.

Sheil:1977:SAA

[586] J. Sheil and I. O’Muircheartaigh. Statistical algorithms: Algorithm AS 106: The distribution of non-negative quadratic forms in normal variables. Applied Statistics, 26(1):92–98, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/106. See remark [872].

Leventhal:1977:SAA

[587] B. Leventhal. Statistical algorithms:

Algorithm AS 107: Operating characteristics and average sampling number for a general class of sequential sampling plans. Applied Statistics, 26(1): 98–106, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/107.

Narula:1977:SAA

[588] Subhash C. Narula and John F. Wellington. Statistical algorithms: Algorithm AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 26(1):106– 111, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/108. See remark [746, 879].

Cran:1977:SAR

[589] G. W. Cran, K. J. Martin, and G. E.

Thomas. Statistical algorithms: Remark AS R19 and Algorithm AS 109: A remark on algorithms: AS 63: The incomplete beta integral AS 64: Inverse of the incomplete beta function ratio. Applied Statistics, 26(1):111–114, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [295, 296, 886].

Sposito:1977:SAA

[590] V. A. Sposito, W. J. Kennedy, and J. E. Gentle. Statistical algorithms: Algorithm AS 110: L_{p} norm fit of a straight line. Applied Statistics, 26(1): 114–118, March 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/110. See remark

[781].

Beasley:1977:SAA

[591] J. D. Beasley and S. G. Springer. Statistical algorithms: Algorithm AS 111: The percentage points of the normal distribution. Applied Statistics, 26(1):118– 121, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/111.

Beardwood:1977:SAR

[592] Jillian E. Beardwood. Statistical algorithms: Remark AS R20: A remark on Algorithm AS 87: Calculation of the polychoric estimate of correlation in contingency tables. Applied Statistics, 26(1):121, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/87. See [437].

Swan:1977:SAR

[593] A. V. Swan. Statistical algorithms: Remark AS R21: A remark on Algorithm AS 16 maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 26(1):122, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/16. See [7,

880].

Anonymous:1977:SAR

[594] Anonymous. Statistical algorithms: Recent statistical algorithms. Applied Statistics, 26(1):122–123, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1977:BMPa

[595] Anonymous. Back matter. Applied Statistics, 26(1):[4], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1977:FMPa

[596] Anonymous. Front matter. Applied Statistics, 26(1):[3], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1977:VIP

[597] Anonymous. Volume information. Applied Statistics, 26(1):[3], 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oldham:1977:FWS

[598] P. D. Oldham and D. J. Newell. Fluoridation of water supplies and cancer — A possible association? Applied Statistics, 26(2):125–135, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnston:1977:AUP

[599] B. Johnston and C. Schwartz. The analysis of an unbalanced paired comparison experiment by multiple regression. Applied Statistics, 26(2):136–142, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vieira:1977:CLG

[600] S. Vieira and R. Hoffmann. Comparison of the logistic and the Gompertz growth functions considering additive and multiplicative error terms. Applied Statistics, 26(2):143–148, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

North:1977:NCM

[601] Philip M. North. A novel clustering method for estimating numbers of bird territories. Applied Statistics, 26 (2):149–155, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1977:TNS

[602] A. N. Pettitt. Testing the normality of several independent samples using the Anderson–Darling statistic. Applied Statistics, 26(2):156–161, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Horsnell:1977:PCP

[603] Gareth Horsnell. Paired comparison product testing when individual preferences are stochastic: An alternative model. Applied Statistics, 26(2): 162–172, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ansley:1977:CTB

[604] C. F. Ansley, W. A. Spivey, and W. J. Wrobleski. A class of transformations for Box-Jenkins seasonal models. Applied Statistics, 26(2):173–178, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosner:1977:EIC

[605] B. Rosner, A. Donner, and C. H. Hennekens. Estimation of interclass correlation from familial data. Applied Statistics, 26(2):179–187, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harvey:1977:MSC

[606] A. C. Harvey. Miscellanea: Some comments on multicollinearity in regression. Applied Statistics, 26(2):188–191, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [430].

Mitchell:1977:MPB

[607] M. S. Mitchell. Miscellanea: The probability of being the lowest bidder. Applied Statistics, 26(2):191–194, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McLeod:1977:CDT

[608] Ian McLeod. Correction: Derivation of the theoretical autocovariance function of autoregressive-moving average time series. Applied Statistics, 26(2):194, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1977:BRB

[609] Graham Kalton. Book reviews: Basic Ideas of Scientific Sampling (2nd Edition), by Alan Stuart. Applied Statistics, 26(2):195, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1977:BRM

[610] D. N. Sparks. Book reviews: Multivariate Analysis, by Maurice Kendall. Applied Statistics, 26(2):195, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1977:BRT

[611] D. N. Sparks. Book reviews: Time Series (2nd Edition), by Maurice Kendall.

Applied Statistics, 26(2):195, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clark:1977:BRF

[612] C. J. Clark. Book reviews: Forecasting for Business: Methods and Applications, by Douglas Wood and Robert Fildes. Applied Statistics, 26(2):196, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1977:BRQ

[613] G. Watts. Book reviews: Quantitative Methods for Business Decisions, by Lawrence L. Lapin. Applied Statistics, 26(2):196, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neal:1977:BRP

[614] Don Neal. Book reviews: Probability Sampling of Hospitals and Patients (2nd Edition), by Irene Hess, Donald C. Riedel and Thomas B. Fitzpatrick. Applied Statistics, 26(2):196–197, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paddle:1977:BRA

[615] G. M. Paddle. Book reviews: Applications of Statistics to Industrial Experimentation, by C. Daniel. Applied Statistics, 26(2):197, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jephcott:1977:BRS

[616] J. St G. Jephcott. Book reviews: Statistical Theory of Sample Survey Design and Analysis, by H. S. Konijn. Applied Statistics, 26(2):198, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1977:BRG

[617] G. Watts. Book reviews: Guide to Official Statistics, by Central Statistical Office. Applied Statistics, 26(2):198, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:1977:SAA

[618] R. J. Baker. Statistical algorithms:

Algorithm AS 112: Exact distributions derived from two-way tables. Applied Statistics, 26(2):199–206, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/112. See remarks [857, 877].

Banfield:1977:SAA

[619] C. F. Banfield and L. C. Bassill. Statistical algorithms: Algorithm AS 113: A transfer for non-hierarchical classification. Applied Statistics, 26(2):206– 210, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/113.

Leathers:1977:SAAa

[620] Burton L. Leathers. Statistical algorithms: Algorithm AS 114: Computing the numerator of ordinal measures of association when the data are ordered categories. Applied Statistics, 26(2):211– 213, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/114.

Baptista:1977:SAA

[621] Joyce Baptista and M. C. Pike. Statistical algorithms: Algorithm AS 115: Exact two-sided confidence limits for the odds ratio in a 2×2 table. Applied Statistics, 26(2):214–220, June 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/115.

Hill:1977:SAR

[622] I. D. Hill and Richard Peto. Statistical algorithms: Remark AS R22: A remark on Algorithm AS 35: “Probabilities Derived from Finite Populations”. Applied Statistics, 26(2):221–226, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/35. See [88,

138, 218].

Anonymous:1977:BMPb

[623] Anonymous. Back matter. Applied Statistics, 26(2):[5], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1977:FMPb

[624] Anonymous. Front matter. Applied Statistics, 26(2):[3], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1977:PHR

[625] R. Kay. Proportional hazard regression models and the analysis of censored survival data. Applied Statistics, 26(3):227–237, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mardia:1977:SCB

[626] K. V. Mardia and R. J. Gadsden. A small circle of best fit for spherical data and areas of vulcanism. Applied Statistics, 26(3):238–245, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Murray:1977:CNS

[627] Gordon D. Murray. A cautionary note on selection of variables in discriminant analysis. Applied Statistics, 26 (3):246–250, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Payne:1977:GAA

[628] R. W. Payne and G. N. Wilkinson. A general algorithm for analysis of variance. Applied Statistics, 26(3):251–260, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sunter:1977:LSS

[629] A. B. Sunter. List sequential sampling with equal or unequal probabilities without replacement. Applied Statistics, 26 (3):261–268, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OCarroll:1977:SPS

[630] F. M. O’Carroll. Subjective probabilities and short-term economic forecasts: An empirical investigation. Applied Statistics, 26(3):269–278, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hsu:1977:TVS

[631] D. A. Hsu. Tests for variance shift at an unknown time point. Applied Statistics, 26(3):279–284, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finney:1977:DS

[632] D. J. Finney. Dimensions of statistics. Applied Statistics, 26(3):285–289, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ozaki:1977:ODA

[633] T. Ozaki. On the order determination of ARIMA models. Applied Statistics, 26 (3):290–301, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Witte:1977:ARU

[634] Ann D. Witte and Peter Schmidt. An analysis of recidivism, using the truncated lognormal distribution. Applied Statistics, 26(3):302–311, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finney:1977:FEV

[635] D. J. Finney and Patricia Phillips. The form and estimation of a variance function, with particular reference to radioimmunoassay. Applied Statistics, 26 (3):312–320, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Franklin:1977:SDC

[636] M. F. Franklin and R. A. Bailey. Selection of defining contrasts and confounded effects in two-level experiments. Applied Statistics, 26(3):321–326, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Besag:1977:SMC

[637] Julian Besag and Peter J. Diggle. Simple Monte Carlo tests for spatial pattern. Applied Statistics, 26(3):327–333, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sundaram:1977:MHW

[638] K. R. Sundaram, I. M. S. Lamba, and A. D. Taskar. Miscellanea: The height and weight of Indian children. Applied Statistics, 26(3):334–335, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Doyle:1977:BRS

[639] Pat Doyle. Book reviews: Statistics for Comparative Studies, by Michael Hills. Applied Statistics, 26(3):336, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1977:BRS

[640] Derek Bissell. Book reviews: Sampling Inspection in Statistical Quality Control, by W. C. Guenther. Applied Statistics, 26(3):337, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hope:1977:BRB

[641] R. P. Hope. Book reviews: Basic Statistics for Business and Economics 2nd Edition, by P. G. Hoel and R. J. Jessen. Applied Statistics, 26(3):337–338, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morton-Williams:1977:BRF

[642] Jean Morton-Williams. Book reviews:

Fertility and Family Planning in Europe Around 1970, by United Nations and Department of Economic and Social Affairs. Applied Statistics, 26(3):338–339, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hills:1977:BRI

[643] M. Hills. Book reviews: An Introduction to Multivariate Techniques for Social and Behavioural Sciences, by S. Bennett and D. Bowers. Applied Statistics, 26(3):339–340, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farlie:1977:BRC

[644] D. J. G. Farlie. Book reviews: Computer Software for Social Science Data, by Beverley Rowe and Marianne Scheer. Applied Statistics, 26(3):340, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1977:BRM

[645] R. J. Prescott. Book reviews: Modelling and Control in the Biomedical Sciences, by H. T. Banks. Applied Statistics, 26 (3):340–341, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holt:1977:BRM

[646] D. Holt. Book reviews: Multivariate Analysis in Behavioural Research, by A. E. Maxwell. Applied Statistics, 26(3):341, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:1977:BRS

[647] J. R. Green. Book reviews: Statistics in Geography: A Practical Approach, by David Ebdon. Applied Statistics, 26 (3):341–342, 1977. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1977:SAA

[648] Morton B. Brown. Statistical algorithms: Algorithm AS 116: The tetrachoric correlation and its asymptotic standard error. Applied Statistics, 26(3): 343–351, September 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/116.

Monro:1977:SAA

[649] Donald M. Monro and John L. Branch.

Statistical algorithms: Algorithm AS 117: The chirp discrete Fourier transform of general length. Applied Statistics, 26(3):351–361, September 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/117.

Westcott:1977:SAA

[650] Brian Westcott. Statistical algorithms: Algorithm AS 118: Approximate rankits. Applied Statistics, 26(3): 362–364, September 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/118.

Leathers:1977:SAAb

[651] Burton L. Leathers. Statistical algorithms: Algorithm AS 119: Tabulating sparse joint frequency distributions. Applied Statistics, 26(3):364–368, September 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/119.

Ley:1977:SAR

[652] J. P. Ley. Statistical algorithms: Remark AS R23: A remark on Algorithm AS 61: “Six-Line Plots”. Applied Statistics, 26(3):368, 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/61. See [268].

Anonymous:1977:BMPc

[653] Anonymous. Back matter. Applied Statistics, 26(3):[2], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1977:FMPc

[654] Anonymous. Front matter. Applied Statistics, 26(3):[3], 1977. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fieller:1978:RSC

[655] N. R. J. Fieller. Report on the 1977 sheffield conference on graphical methods in statistics. Applied Statistics, 27(1):1–3, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1978:SRR

[656] D. R. Cox. Some remarks on the role in statistics of graphical methods. Applied Statistics, 27(1):4–9, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1978:EBT

[657] P. Prescott. Examination of the behaviour of tests for outliers when more than one outlier is present. Applied Statistics, 27(1):10–25, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Silverman:1978:DRE

[658] B. W. Silverman. Density ratios, empirical likelihood and cot death. Applied Statistics, 27(1):26–33, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1978:BBA

[659] Martin J. Crowder. Beta-binomial Anova for proportions. Applied Statistics, 27(1):34–37, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jarrett:1978:ADE

[660] Richard G. Jarrett. The analysis of designed experiments with missing observations. Applied Statistics, 27(1):38–46, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1978:CUS

[661] G. H. Brown. Calibration with an ultrastructural model. Applied Statistics, 27 (1):47–51, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mills:1978:FFU

[662] T. C. Mills. The functional form of the U.K. demand for money. Applied Statistics, 27(1):52–57, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1978:MRI

[663] D. G. Jones. A method of re-interpreting promotion rates in terms of promotion prospectuses. Applied Statistics, 27(1):58–66, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Poirier:1978:EFO

[664] Dale J. Poirier. The effect of the first observation in regression models with firstorder autoregressive disturbances. Applied Statistics, 27(1):67–68, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Powell:1978:BBM

[665] N. K. Powell and J. B. Westwood. Buyer-behaviour in management education. Applied Statistics, 27(1):69–72, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1978:MJC

[666] P. R. Freeman. Miscellanea: Joint conference at Manchester — A summary report. Applied Statistics, 27(1):73–75, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hamaker:1978:MAC

[667] Hugo C. Hamaker. Miscellanea: Approximating the cumulative normal distribution and its inverse. Applied Statistics, 27(1):76–77, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1978:MSD

[668] A. A. Greenfield. Miscellanea: Selection of defining contrasts in two-level experiments — A modification. Applied Statistics, 27(1):78, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Theobald:1978:LE

[669] C. M. Theobald and R. G. Hunt. Letters to the editors. Applied Statistics, 27(1):79–80, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1978:BRP

[670] A. A. Greenfield. Book reviews: Programming and Algorithms: An Introduction, by A. J. Guttmann. Applied Statistics, 27(1):81, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1978:BRS

[671] A. A. Greenfield. Book reviews: Statistical Methods for Digital Computers, by Kurt Enslein, Anthony Ralston and Herbert S. Wilf. Applied Statistics, 27 (1):81–83, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1978:BRD

[672] A. S. C. Ehrenberg. Book reviews:

Descriptive Statistics: A Contemporary Approach, by Richard P. Runyon. Applied Statistics, 27(1):84, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1978:BRW

[673] A. S. C. Ehrenberg. Book reviews: Winning with Statistics: A Painless First Look at Numbers, Ratios, Percentages, Means, and Inference, by Richard P. Runyon. Applied Statistics, 27(1):84, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1978:BRQ

[674] Graham Watts. Book reviews: Quantitative Analysis for Business Decisions, by H. Bierman, C. P. Bonini and W. H. Hausman. Applied Statistics, 27(1):84, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tennant-Smith:1978:BRM

[675] J. Tennant-Smith. Book reviews: Models of Preventive Maintenance, by Ilya B. Gertsbakh. Applied Statistics, 27(1): 85, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:1978:BRP

[676] P. J. Green. Book reviews: Probability and Statistical Inference, by Robert V. Hogg and Elliot A. Tanis. Applied Statistics, 27(1):85–86, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ord:1978:BRH

[677] Keith Ord. Book reviews: Handbook of Statistical Distributions, by J. K. Patel, C. H. Kapadia and D. B. Owen. Applied Statistics, 27(1):86–87, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1978:BRI

[678] D. V. Lindley. Book reviews: Introduction to Decision Theory, by J. Morgan Jones. Applied Statistics, 27(1):87, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collings:1978:BRG

[679] S. N. Collings. Book reviews: General Sources of Statistics, by G. F. Lock. Applied Statistics, 27(1):87–88, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harris:1978:BRAa

[680] P. T. Harris. Book reviews: The Analysis of Survey Data. Vol. 1: Exploring Data Structures, by C. A. O’Muircheataigh and C. Payne. Applied Statistics, 27(1):88, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harris:1978:BRAb

[681] P. T. Harris. Book reviews: The Analysis of Survey Data. Vol. 2: Model Fitting, by C. A. O’Muircheataigh and C. Payne. Applied Statistics, 27(1):88, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1978:BRI

[682] I. G. Evans. Book reviews: Introductory Statistics: A Decision Map, by Thad R. Harshbarger. Applied Statistics, 27(1): 89, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Waters:1978:BRR

[683] H. R. Waters. Book reviews: Risk Theory: The Stochastic Basis of Insurance, by R. E. Beard, T. Pentikainen and E. Pesonen. Applied Statistics, 27(1):89–90, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1978:BRS

[684] Graham Watts. Book reviews: Statistical Methods for Business and Economics, by Roger C. Pfaffenberger and James H. Patterson. Applied Statistics, 27(1):90, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1978:BRM

[685] Graham Watts. Book reviews: Mathematical Models in Marketing — A Collection of Abstracts, by Ursula H. Funke. Applied Statistics, 27(1):90–91, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rubin:1978:SAA

[686] Donald B. Rubin, Thomas W. F. Stroud, and Dorothy Thayer. Statistical algorithms: Algorithm AS 120: A Fortran algorithm for the additive model in a two-way unbalanced manova. Applied Statistics, 27(1):92–97, March 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/120.

Schneider:1978:SAA

[687] B. E. Schneider. Statistical algorithms: Algorithm AS 121: Trigamma function. Applied Statistics, 27(1):97– 99, March 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/121.

Bohrer:1978:SAA

[688] Robert Bohrer and Winston Chow. Statistical algorithms: Algorithm AS 122: Weights for one-sided multivariate inference. Applied Statistics, 27(1):100– 104, March 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/122. See remark [860].

Bremner:1978:SAA

[689] J. M. Bremner. Statistical algorithms: Algorithm AS 123: Mixtures of beta distributions. Applied Statistics, 27(1):104– 109, March 1978. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/123.

Baker:1978:SAC

[690] R. J. Baker. Statistical algorithms: Correction: AS 112: Exact distributions derived from two-way tables. Applied Statistics, 27(1):109, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1978:BMPa

[691] Anonymous. Back matter. Applied Statistics, 27(1):[4], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1978:FMPa

[692] Anonymous. Front matter. Applied Statistics, 27(1):[3], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1978:VIP

[693] Anonymous. Volume information. Applied Statistics, 27(1):[3], 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1978:OFE

[694] J. A. John. Outliers in factorial experiments. Applied Statistics, 27(2): 111–119, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abraham:1978:DFA

[695] Bovas Abraham and George E. P.

Box. Deterministic and forecastadaptive time-dependent models. Applied Statistics, 27(2):120–130, 1978.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abraham:1978:LMS

[696] Bovas Abraham and George E. P. Box. Linear models and spurious observations. Applied Statistics, 27(2):131–138, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cohen:1978:PEB

[697] G. Cohen and K. M. Crawford. A problem in estimating bus stop times. Applied Statistics, 27(2):139–148, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1978:EWO

[698] Graham Kalton, Martin Collins, and Lindsay Brook. Experiments in wording opinion questions. Applied Statistics, 27 (2):149–161, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altham:1978:TGB

[699] Patricia M. E. Altham. Two generalizations of the binomial distribution. Applied Statistics, 27(2):162–167, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chan:1978:MTS

[700] W.-Y. T. Chan and Kenneth F. Wallis. Multiple time series modelling: Another look at the mink-muskrat interaction. Applied Statistics, 27(2):168–175, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Casagrande:1978:PFE

[701] J. T. Casagrande, M. C. Pike, and P. G. Smith. The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 27(2):176–180, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See correction [850].

Best:1978:LE

[702] D. J. Best and R. W. Farebrother. Letters to the editors. Applied Statistics, 27 (2):181–182, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dean:1978:BRE

[703] Angela Dean. Book reviews: Experiments: Design and Analysis (2nd Edition), by J. A. John and M. H. Quenouille. Applied Statistics, 27(2): 182–183, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holgate:1978:BRH

[704] P. Holgate. Book reviews: On the History of Statistics and Probability, by D. B. Owen. Applied Statistics, 27(2):183, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conway:1978:BRS

[705] Freda Conway. Book reviews: Statistical Teaching Aids (2nd Edition), by T. Szmidt and A. F. Bissell. Applied Statistics, 27(2):183–184, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evans:1978:BRS

[706] I. G. Evans. Book reviews: Statistical Analysis for Decision Making (2nd Edition), by Morris Hamburg. Applied Statistics, 27(2):184–185, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1978:BRS

[707] E. Kay. Book reviews: Statistical Design of Fatigue Experiments, by R. E. Little and E. H. Jebe. Applied Statistics, 27(2):185, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fessey:1978:BRS

[708] M. C. Fessey. Book reviews: Statistics and Public Policy, by William B. Fairley and Frederick Mosteller. Applied Statistics, 27(2):186, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stern:1978:BRQ

[709] G. J. A. Stern. Book reviews: Queueing Systems, Volume 2: Computer Applications, by Leonard Kleinrock. Applied Statistics, 27(2):186–187, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neave:1978:SAA

[710] Henry R. Neave. Statistical algorithms: Algorithm AS 124: Sample sizes for onesided and strong two-sided distributionfree tolerance limits. Applied Statistics, 27(2):188–190, June 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/124.

Lagakos:1978:SAA

[711] S. W. Lagakos and M. H. Kuhns. Statistical algorithms: Algorithm AS 125: Maximum likelihood estimation for censored exponential survival data with covariates. Applied Statistics, 27(2):190– 197, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/125. See remark [862].

Barnard:1978:SAA

[712] J. Barnard. Statistical algorithms: Algorithm AS 126: Probability integral of the normal range. Applied Statistics, 27(2): 197–198, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/126. See remark [865].

Heiberger:1978:SAA

[713] Richard M. Heiberger. Statistical algorithms: Algorithm AS 127: Generation of random orthogonal matrices. Applied Statistics, 27(2):199– 206, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/127. See remark [866].

Davis:1978:SAA

[714] C. S. Davis and M. A. Stephens. Statistical algorithms: Algorithm AS 128: Approximating the covariance matrix of normal order statistics. Applied Statistics, 27(2):206–212, June 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/128. See

remark [881].

Casagrande:1978:SAA

[715] J. T. Casagrande, M. C. Pike, and P. G.

Smith. Statistical algorithms: Algorithm AS 129: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 27(2): 212–219, June 1978. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/129. See correction [850] and remark [858].

Anonymous:1978:BMPb

[716] Anonymous. Back matter. Applied Statistics, 27(2):[4], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1978:FMPb

[717] Anonymous. Front matter. Applied Statistics, 27(2):[3], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bebbington:1978:MBT

[718] A. C. Bebbington. A method of bivariate trimming for robust estimation of the correlation coefficient. Applied Statistics, 27(3):221–226, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Titterington:1978:ECC

[719] D. M. Titterington. Estimation of correlation coefficients by ellipsoidal trimming. Applied Statistics, 27(3):227–234, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lagakos:1978:CMP

[720] S. W. Lagakos. A covariate model for partially censored data subject to competing causes of failure. Applied Statistics, 27(3):235–241, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1978:SOP

[721] Vic Barnett. The study of outliers: Purpose and model. Applied Statistics, 27 (3):242–250, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1978:IFA

[722] Norm A. Campbell. The influence function as an aid in outlier detection in discriminant analysis. Applied Statistics, 27 (3):251–258, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McKay:1978:GAS

[723] R. J. McKay. A graphical aid to selection of variables in two-group discriminant analysis. Applied Statistics, 27 (3):259–263, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1978:HWF

[724] C. Chatfield. The Holt-Winters forecasting procedure. Applied Statistics, 27 (3):264–279, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zar:1978:APP

[725] Jerrold H. Zar. Approximations for the percentage points of the chi-squared distribution. Applied Statistics, 27(3): 280–290, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McHenry:1978:CBS

[726] C. E. McHenry. Computation of a best subset in multivariate analysis. Applied Statistics, 27(3):291–296, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Constantine:1978:GRA

[727] A. G. Constantine and J. C. Gower. Graphical representation of asymmetric matrices. Applied Statistics, 27 (3):297–304, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bunday:1978:ETW

[728] B. D. Bunday and R. A. Lee. The efficiency of a textile winding process. Applied Statistics, 27(3):305–309, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1978:CRL

[729] Martin J. Crowder. On concurrent regression lines. Applied Statistics, 27 (3):310–318, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sahin:1978:SDM

[730] Izzet Sahin and David J. Hendrick. On strike durations and a measure of termination. Applied Statistics, 27(3): 319–324, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Murray:1978:EPD

[731] G. D. Murray and D. M. Titterington. Estimation problems with data from a mixture. Applied Statistics, 27 (3):325–334, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Patterson:1978:DKF

[732] H. D. Patterson and R. A. Bailey. Design keys for factorial experiments. Applied Statistics, 27(3):335–343, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tavare:1978:MCE

[733] S. Tavare. Miscellanea: Conference on ecological statistics — A summary report. Applied Statistics, 27(3):344–345, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dobson:1978:MSA

[734] Annette J. Dobson. Miscellanea: Simple approximations for the von Mises concentration statistic. Applied Statistics, 27(3):345–347, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altman:1978:MPP

[735] D. G. Altman. Miscellanea: Plotting probability ellipses. Applied Statistics, 27(3):347–349, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nandy:1978:LE

[736] S. Nandy. Letter to the editors. Applied Statistics, 27(3):350, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1978:BRC

[737] I. D. Hill. Book reviews: Computer Programs that Work!, by J. D. Lee, G. Beech and T. D. Lee. Applied Statistics, 27 (3):351–352, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sampford:1978:BRS

[738] M. R. Sampford. Book reviews: Sampling Techniques, by W. G. Cochran.

Applied Statistics, 27(3):352, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1978:BRT

[739] D. N. Sparks. Book reviews: Tables of Random Times, by I. D. Hill. Applied Statistics, 27(3):352–353, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1978:BRB

[740] J. Taylor. Book reviews: Basic Statistics for Medical and Social Students, by A. E. Maxwell. Applied Statistics, 27(3):353, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1978:SAA

[741] Jane F. Gentleman. Statistical algorithms: Algorithm AS 130: Moving statistics for enhanced scatter plots. Applied Statistics, 27(3):354–358, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/130.

Leathers:1978:SAA

[742] Burton L. Leathers. Statistical algorithms: Algorithm AS 131: Tabulating frequency distributions for variables with structured code sets. Applied Statistics, 27(3):359–362, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/131. See remark [884].

Armstrong:1978:SAA

[743] Ronald D. Armstrong and Mabel Tam Kung. Statistical algorithms: Algorithm AS 132: Least absolute value estimates for a simple linear regression problem. Applied Statistics, 27(3):363– 366, September 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/132.

Zilinskas:1978:SAA

[744] A. Žilinskas. Statistical algorithms: Algorithm AS 133: Optimization of onedimensional multimodal functions. Applied Statistics, 27(3):367–375, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/133.

Hoaglin:1978:SAR

[745] David C. Hoaglin and Max L. King. Statistical algorithms: Remark AS R24: A remark on Algorithm AS 98: The spectral test for the evaluation of congruential pseudo-random generators. Applied Statistics, 27(3):375–377, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/98. See

[518, 559].

Armstrong:1978:SAR

[746] Ronald D. Armstrong and Edward L.

Frome. Statistical algorithms: Remark AS R25: A remark on Algorithm AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 27(3):378, 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/108. See [588,

879].

Hill:1978:SARa

[747] I. D. Hill. Statistical algorithms: Remark AS R26: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 27(3):379, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [382, 782, 874, 883, 878].

Barrett:1978:SAR

[748] J. C. Barrett and M. J. R. Healy. Statistical algorithms: Remark AS R27: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 27(3): 379–380, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/6. See [386, 3, 8].

Hill:1978:SARb

[749] I. D. Hill. Statistical algorithms: Remark AS R28: A remark on Algorithm AS 47: Function minimization using a simplex procedure. Applied Statistics, 27 (3):380–382, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/47. See [137, 348, 349].

Anonymous:1978:BMPc

[750] Anonymous. Back matter. Applied Statistics, 27(3):[6], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1978:FMPc

[751] Anonymous. Front matter. Applied Statistics, 27(3):[3], 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bennett:1979:ATS

[752] W. F. Bennett. The analysis of transient spectral components with the autoregressive spectral estimator. Applied Statistics, 28(1):1–13, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1979:MMD

[753] Trevor F. Cox. A method for mapping the dense and sparse regions of a forest stand. Applied Statistics, 28(1):14–19, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dawid:1979:MLE

[754] A. P. Dawid and A. M. Skene. Maximum likelihood estimation of observer error-rates using the EM algorithm. Applied Statistics, 28(1):20–28, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1979:CGP

[755] A. C. Atkinson. The computer generation of Poisson random variables. Applied Statistics, 28(1):29–35, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1979:CVS

[756] P. Prescott. Critical values for a sequential test for many outliers. Applied Statistics, 28(1):36–39, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kahn:1979:LSE

[757] Henry D. Kahn. Least squares estimation for the inverse power law for accelerated life tests. Applied Statistics, 28(1):40–46, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Koons:1979:MSA

[758] G. F. Koons, J. J. Geissler, and B. Perlic. A method for the statistical analysis of a changeable independent variable in a batch process. Applied Statistics, 28(1):47–54, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitken:1979:ADK

[759] Colin Aitken and D. Gordon MacDonald. An application of discrete kernel methods to forensic odontology. Applied Statistics, 28(1):55–61, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hsu:1979:LTD

[760] D. A. Hsu. Long-tailed distributions for position errors in navigation. Applied Statistics, 28(1):62–72, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:BRR

[761] Anonymous. Book reviews: Review of Tables of Random Times: Corrections. Applied Statistics, 28(1):72, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:BRT

[762] Anonymous. Book reviews: Tables of Random Times, by I. D. Hill and P. A. Hill. Applied Statistics, 28(1):72, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sundaram:1979:MCH

[763] K. R. Sundaram, I. M. S. Lamba, and A. D. Taskar. Miscellanea: Corrections: The height and weight of Indian children. Applied Statistics, 28(1):72, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farewell:1979:MNM

[764] V. T. Farewell and D. R. Cox. Miscellanea: A note on multiple time scales in life testing. Applied Statistics, 28(1):73– 75, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marriott:1979:MBM

[765] F. H. C. Marriott. Miscellanea:

Barnard’s Monte Carlo tests: How many simulations? Applied Statistics, 28 (1):75–77, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1979:LE

[766] P. N. Lee. Letter to the editors. Applied Statistics, 28(1):78, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1979:BRE

[767] A. S. C. Ehrenberg. Book reviews: Exploratory Data Analysis, by John W. Tukey. Applied Statistics, 28(1):79–83, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1979:BRS

[768] I. D. Hill. Book reviews: Scientific Analysis on the Pocket Calculator (2nd Edition), by Jon M. Smith. Applied Statistics, 28(1):83–84, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawrance:1979:BRS

[769] A. J. Lawrance. Book reviews: Stochastic Modelling of Daily Rainfall Sequences, by T. A. Buishand. Applied Statistics, 28(1):84, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRG

[770] Tony Greenfield. Book reviews: A guide to operational research (3rd edition), by W. E. Duckworth, A. E. Gear and A. G. Lockett. Applied Statistics, 28(1):85, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRIa

[771] Tony Greenfield. Book reviews: Introduction to the Design and Analysis of Algorithms, by S. E. Goodman and S. T. Hedetniemi. Applied Statistics, 28(1):85–86, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRSa

[772] Tony Greenfield. Book reviews: Statistics and Experimental Design in Engineering and the Physical Sciences (2nd Edition), by Norman L. Johnson and Fred C. Leone. Applied Statistics, 28 (1):86–87, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRN

[773] Tony Greenfield. Book reviews: Nonparametric Statistics: A Contemporary Approach, by Richard P. Runyon. Applied Statistics, 28(1):87, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1979:BRC

[774] B. E. Cooper. Book reviews: COMP-STAT 1976, by Johannes Gordesch and Peter Naeve. Applied Statistics, 28 (1):87–88, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1979:BRS

[775] Graham Watts. Book reviews: Statistical Methods for Business and Economics, by Roger C. Pfaffenberger and James H. Patterson. Applied Statistics, 28(1):88–89, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cooper:1979:BRG

[776] B. E. Cooper. Book reviews: GEM.

A general econometric matrix program, by Lucy Joan Slater. Applied Statistics, 28(1):89, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1979:SAA

[777] A. C. Atkinson and J. Whittaker. Statistical algorithms: Algorithm AS 134: The generation of beta random variables with one parameter greater than and one parameter less than 1. Applied Statistics, 28(1):90–93, March 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/134.

Armstrong:1979:SAA

[778] Ronald D. Armstrong and David S.

Kung. Statistical algorithms: Algorithm AS 135: Min-max estimates for a linear multiple regression problem. Applied Statistics, 28(1):93– 100, March 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/135. See remark [870].

Hartigan:1979:SAA

[779] J. A. Hartigan and M. A. Wong. Statistical algorithms: Algorithm AS 136: A K-means clustering algorithm. Applied Statistics, 28(1):100– 108, March 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/136. See remark [863].

Ripley:1979:SAA

[780] B. D. Ripley. Statistical algorithms: Algorithm AS 137: Simulating spatial patterns: Dependent samples from a multivariate density. Applied Statistics, 28 (1):109–112, March 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/137.

Porter:1979:SAR

[781] M. A. Porter and D. J. Winstanley. Statistical algorithms: Remark AS R29: Remarks on AS 110: L_{p} norm fit of a straight line. Applied Statistics, 28 (1):112–113, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [590].

Thomas:1979:SAR

[782] G. E. Thomas. Statistical algorithms: Remark AS R30: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 28(1):113, 1979. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [747, 382, 883, 874, 878].

Young:1979:SAC

[783] J. C. Young and Ch. E. Minder. Statistical algorithms: Correction: AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 28(1):113, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:BMPa

[784] Anonymous. Back matter. Applied Statistics, 28(1):[4], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:FMPa

[785] Anonymous. Front matter. Applied Statistics, 28(1):[3], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altham:1979:DRB

[786] Patricia M. E. Altham. Detecting relationships between categorical variables observed over time: a problem of deflating a chi-squared statistic. Applied Statistics, 28(2):115–125, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1979:NPA

[787] A. N. Pettitt. A non-parametric approach to the change-point problem. Applied Statistics, 28(2):126–135, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farewell:1979:ACP

[788] V. T. Farewell. An application of Cox’s proportional hazard model to multiple infection data. Applied Statistics, 28 (2):136–143, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shirley:1979:CTC

[789] Eryl A. C. Shirley. The comparison of treatment with control group means in toxicological studies. Applied Statistics, 28(2):144–151, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Best:1979:ESM

[790] D. J. Best and N. I. Fisher. Efficient simulation of the von Mises distribution. Applied Statistics, 28(2):152–157, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bache:1979:APP

[791] Niels Bache. Approximate percentage points for the distribution of a product of independent positive random variables. Applied Statistics, 28(2):158–162, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hutchinson:1979:CRP

[792] T. P. Hutchinson. A comment on replicated paired comparisons. Applied Statistics, 28(2):163–169, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [364].

Hsuan:1979:GUP

[793] Francis C. Hsuan. Generating uniform polygonal random pairs. Applied Statistics, 28(2):170–172, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1979:MST

[794] G. H. Jowett. Miscellanea: Statistical tests with hand-calculator standard deviations. Applied Statistics, 28 (2):173–174, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schmeiser:1979:MAI

[795] Bruce W. Schmeiser. Miscellanea: Approximations to the inverse cumulative normal function for use on hand calculators. Applied Statistics, 28(2):175–176, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clifford:1979:BRS

[796] F. E. Clifford. Book reviews: Statistics for Technology. 2nd Edition, by C. Chatfield. Applied Statistics, 28(2):177, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holder:1979:BRD

[797] R. L. Holder. Book reviews: Data Analysis and Regression, by Frederick Mosteller and John W. Tukey. Applied Statistics, 28(2):177–178, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittaker:1979:BRM

[798] J. C. Whittaker. Book reviews: Mathematical, Statistical and Financial Tables for the Social Sciences, by Z. W. Kmietowicz and Y. Yannoulis. Applied Statistics, 28(2):178–179, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bithell:1979:BRI

[799] J. F. Bithell. Book reviews: Introductory Biostatistics for the Health Sciences, by Robert C. Duncan, Rebecca G. Knapp and M. Clinton Miller. Applied Statistics, 28(2):179–180, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1979:BRI

[800] P. J. Brown. Book reviews: The Information in Contingency Tables, by D. V. Gokhale and S. Kullback. Applied Statistics, 28(2):180–181, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berry:1979:BRI

[801] J. Brian Berry. Book reviews: Interactive Data Analysis: A Practical Primer, by D. R. McNeil. Applied Statistics, 28(2):181, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRIb

[802] Tony Greenfield. Book reviews: An Introduction to Linear Control Systems, by Thomas E. Fortmann and Konrad L. Hitz. Applied Statistics, 28(2):181–182, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1979:BRE

[803] Derek Bissell. Book reviews: Elementary Statistics: An Applied Approach, by Neil R. Ullman. Applied Statistics, 28 (2):182–183, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Springall:1979:BRA

[804] A. Springall. Book reviews: Applied Statistics for Science and Industry, by Albert Romano. Applied Statistics, 28(2):183, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oldham:1979:LE

[805] P. D. Oldham and D. J. Newell. Letter to the editors. Applied Statistics, 28(2):184, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wolynetz:1979:SAAa

[806] M. S. Wolynetz. Statistical algorithms:

Algorithm AS 138: Maximum likelihood estimation from confined and censored normal data. Applied Statistics, 28(2): 185–195, June 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/138.

Wolynetz:1979:SAAb

[807] M. S. Wolynetz. Statistical algorithms: Algorithm AS 139: Maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 28(2):195– 206, June 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/139. See remarks [852, 855].

Oehlert:1979:SAA

[808] Gary W. Oehlert. Statistical algorithms:

Algorithm AS 140: Clustering the nodes of a directed graph. Applied Statistics, 28(2):206–214, June 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/140.

Kent:1979:SAA

[809] Philippe Kent. Statistical algorithms: Algorithm AS 141: Inversion of a symmetric matrix in regression models. Applied Statistics, 28(2):214– 217, June 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/141. See remark [864].

Anonymous:1979:BMPb

[810] Anonymous. Back matter. Applied Statistics, 28(2):[8], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:EIP

[811] Anonymous. Editorial: Invitation to potential authors of “case study” papers. Applied Statistics, 28(2):[1], 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:FMPb

[812] Anonymous. Front matter. Applied Statistics, 28(2):[3], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunliffe:1979:EAS

[813] Stella Y. Cunliffe and H. Goldstein. Ethical aspects of survey research. Applied Statistics, 28(3):219–222, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moran:1979:CLT

[814] M. A. Moran and B. J. Murphy. A closer look at two alternative methods of statistical discrimination. Applied Statistics, 28(3):223–232, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1979:STP

[815] Murray Aitkin. A simultaneous test procedure for contingency table models. Applied Statistics, 28(3):233–242, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1979:TSO

[816] A. C. Kimber. Tests for a single outlier in a gamma sample with unknown shape and scale parameters. Applied Statistics, 28(3):243–250, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glasbey:1979:CRN

[817] C. A. Glasbey. Correlated residuals in non-linear regression applied to growth data. Applied Statistics, 28(3):251–259, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1979:RDC

[818] A. C. Atkinson. Recent developments in the computer generation of Poisson random variables. Applied Statistics, 28 (3):260–263, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1979:MLS

[819] R. J. Prescott. Models for local and systemic effects in bilateral systems. Applied Statistics, 28(3):264–270, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ury:1979:TDS

[820] Hans K. Ury and David C. Kleinecke. Tables of the distribution of Spearman’s footrule. Applied Statistics, 28 (3):271–275, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kalton:1979:SCV

[821] Graham Kalton and Richard Stowell. A study of coder variability. Applied Statistics, 28(3):276–289, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheng:1979:SSG

[822] R. C. H. Cheng and G. M. Feast. Some simple gamma variate generators. Applied Statistics, 28(3):290–295, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walley:1979:TRK

[823] D. Walley and T. Fearn. Trapezium rules, O.K. Applied Statistics, 28(3): 296–297, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newby:1979:MSO

[824] M. J. Newby. Miscellanea: The simulation of order statistics from life distributions. Applied Statistics, 28(3):298–300, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

North:1979:MNC

[825] Philip M. North. Miscellanea: A novel clustering method for estimating numbers of bird territories: An addendum. Applied Statistics, 28(3):300–301, 1979.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1979:LE

[826] Ian Clarke. Letter to the editors. Applied Statistics, 28(3):302, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1979:BRA

[827] V. Barnett. Book reviews: Applied Statistics, by John Neter, William Wasserman and G. A. Whitmore. Applied Statistics, 28(3):303, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1979:BRC

[828] I. D. Hill. Book reviews: Compatible Fortran, by A. Colin Day. Applied Statistics, 28(3):304, 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1979:BRA

[829] Trevor Lambert. Book reviews: Algorithmic Methods in Probability-TIMS Studies in Management Sciences, Volume 7, by Marcel F. Neuts. Applied Statistics, 28(3):304–305, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1979:BRI

[830] Oliver D. Anderson. Book reviews:

An Introduction to Bilinear Time Series Models, by C. W. J. Granger and A. P. Andersen. Applied Statistics, 28 (3):305–306, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Settle:1979:BRL

[831] J. G. Settle. Book reviews: Long-Range Forecasting, by J. Scott Armstrong. Applied Statistics, 28(3):306, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1979:BRA

[832] J. A. Nelder. Book reviews: The Analysis of Cross-Tabulated Data, by Graham J. G. Upton. Applied Statistics, 28 (3):306–307, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chinn:1979:BRB

[833] Susan Chinn. Book reviews: Biostatistics: A Foundation for Analysis in the Health Sciences. 2nd Edition, by Wayne W. Daniel. Applied Statistics, 28 (3):307–308, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Titterington:1979:BRS

[834] D. M. Titterington. Book reviews: Statistical Density Estimation: A Survey, by W. Wertz. Applied Statistics, 28(3):308, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1979:BRS

[835] Chris Chatfield. Book reviews: Sales Forecasting, by David L. Hurwood, Elliott S. Grossman and Earl L. Bailey. Applied Statistics, 28(3):308–309, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1979:BRSb

[836] Tony Greenfield. Book reviews: Statistical Checklists, by J. N. R. Jeffers. Applied Statistics, 28(3):309, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holder:1979:BRA

[837] R. L. Holder. Book reviews: Applied Nonparametric Statistics, by Wayne W. Daniel. Applied Statistics, 28(3): 309–310, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Computing:1979:CDA

[838] The Working Party on Statistical Computing. The construction and description of algorithms. Applied Statistics, 28 (3):311–318, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bayer:1979:SAA

[839] Leonard Bayer and Christopher Cox.

Statistical algorithms: Algorithm AS 142: Exact tests of significance in binary regression models. Applied Statistics, 28(3):319–324, September 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/142.

Bedall:1979:SAA

[840] F. K. Bedall and H. Zimmermann. Statistical algorithms: Algorithm AS 143: The mediancentre. Applied Statistics, 28(3):325–328, September 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/143.

Boyett:1979:SAA

[841] James M. Boyett. Statistical algorithms: Algorithm AS 144: Random R × C tables with given row and column totals. Applied Statistics, 28(3):329–332, September 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/144.

Freeman:1979:SAA

[842] P. R. Freeman. Statistical algorithms: Algorithm AS 145: Exact distribution of the largest multinomial frequency. Applied Statistics, 28(3):333– 336, September 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/145.

Anonymous:1979:SACa

[843] Anonymous. Statistical algorithms: Corrections: Algorithm AS 141. Applied Statistics, 28(3):336, 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/141.

Anonymous:1979:SACb

[844] Anonymous. Statistical algorithms: Corrections: (Algorithm AS 76): Remark AS R30. Applied Statistics, 28(3):336, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/76.

Anonymous:1979:RSA

[845] Anonymous. Recent statistical algorithm. Applied Statistics, 28(3):337–340, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:BMPc

[846] Anonymous. Back matter. Applied Statistics, 28(3):[2], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:FMPc

[847] Anonymous. Front matter. Applied Statistics, 28(3):[3], 1979. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1979:VIP

[848] Anonymous. Volume information. Applied Statistics, 28(3):[3], 1979. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lau:1980:SAA

[849] Chi-Leung L. Lau. Statistical algorithms: Algorithm AS 147: A simple series for the incomplete gamma integral. Applied Statistics, 29(1):113– 114, March 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/147. See corrigenda [853] and remarks [876, 882].

Casagrande:1980:CPF

[850] J. T. Casagrande, M. C. Pike, and P. G.

Smith. Correction to: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 29(1):118, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [701, 715].

Farebrother:1980:SAA

[851] R. W. Farebrother. Statistical algorithms: Algorithm AS 153: Pan’s procedure for the tail probabilities of the Durbin-Watson statistic. Applied Statistics, 29(2):224–227, June 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/153. See

remark [871, 872].

Wolynetz:1980:SAR

[852] M. S. Wolynetz. Statistical algorithms: Remark AS R31: A remark on Algorithm AS 139. maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 29(2):228, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/139. See

[807, 855].

Lau:1980:CAS

[853] Chi-Leung Lau. Corrigenda: Algorithm AS 147. A simple series for the incomplete gamma integral. Applied Statistics, 29(2):229, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/147. See [849, 882].

Davies:1980:SAA

[854] Robert B. Davies. Statistical algorithms: Algorithm AS 155: The distribution of a linear combination of χ^{2} random variables. Applied Statistics, 29(3):323–333, September 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/155. See remark [872].

Wolynetz:1981:SAR

[855] M. S. Wolynetz. Statistical algorithms: Remark AS R32: A remark on Algorithm AS 139: Maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 30(1):105, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/139. See

[807, 852].

Hill:1981:SAR

[856] I. D. Hill and R. E. Wheeler. Statistical algorithms: Remark AS R33: A remark on algorithms AS 99: Fitting Johnson curves by moments and AS 100: Normal-Johnson and Johnsonnormal transformations. Applied Statistics, 30(1):106, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [519, 520, 869].

Farley:1981:SAR

[857] T. M. M. Farley. Statistical algorithms:

Remark AS R34: A remark on Algorithm AS 112: Exact distributions derived from two-way tables. Applied Statistics, 30(1):106–107, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/112. See

[618, 877].

Casagrande:1981:SAR

[858] J. T. Casagrande. Statistical algorithms:

Remark AS R35: A remark on Algorithm AS 129: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 30(1):108, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/129. See [715].

Stirling:1981:SAAc

[859] W. Douglas Stirling. Statistical algorithms: Algorithm AS 169: An improved algorithm for scatter plots. Applied Statistics, 30(3):345–349, September 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/169. See remarks [868, 873].

Bohrer:1981:SAR

[860] Robert Bohrer. Statistical algorithms: Remark AS R36: A remark on AS 122: Weights for one-sided multivariable inference. Applied Statistics, 30 (3):352–353, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [688].

Frome:1981:SARa

[861] E. L. Frome. Statistical algorithms: Remark AS R37: A remark on AS 73: Cross-spectrum smoothing via the finite Fourier transform. Applied Statistics, 30(3):354, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [345].

Frome:1981:SARb

[862] E. L. Frome. Statistical algorithms: Remark AS R38: A remark on AS 125: Maximum likelihood estimation for censored exponential survival data with covariates. Applied Statistics, 30(3):355, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [711].

England:1981:SAR

[863] R. England and D. Beynon. Statistical algorithms: Remark AS R39: A remark on AS 136: A K-means clustering algorithm. Applied Statistics, 30(3):355–356, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See remark [779].

Campbell:1981:SAR

[864] M. J. Campbell. Statistical algorithms:

Remark AS R40: A remark on AS 141: Inversion of a symmetric matrix in regression models. Applied Statistics, 30(3):356, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [809].

elLozy:1982:SAR

[865] Mohamed el Lozy. Statistical algorithms: Remark AS R41: A remark on Algorithm AS 126: Probability integral of the normal range. Applied Statistics, 31(1):99, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/126. See [712].

Tanner:1982:SAR

[866] Martin A. Tanner and Ronald A.

Thisted. Statistical algorithms: Remark AS R42: A remark on AS 127. generation of random orthogonal matrices. Applied Statistics, 31(2):190–192, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [713].

Freeman:1982:SAR

[867] P. R. Freeman. Statistical algorithms:

Remark AS R44: A remark on AS 6 and AS 7: Triangular decomposition of a symmetric matrix and inversion of a positive semi-definite symmetric matrix. Applied Statistics, 31(3):336–339, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [3, 4].

Harris:1982:SAR

[868] R. G. Harris. Statistical algorithms: Remark AS R46: Remark on AS 169: An improved algorithm for scatter plots. Applied Statistics, 31(3):340, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [859, 873].

Dodgson:1983:SAR

[869] J. H. Dodgson and I. D. Hill. Statistical algorithms: Remark AS R48: A remark on algorithm 100: Normal-Johnson and Johnson-normal transformations. Applied Statistics, 32(3):345, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [520, 856].

Grant:1983:SAR

[870] P. M. Grant and T. R. Hopkins. Statistical algorithms: Remark AS R49: A remark on Algorithm AS 135: Minmax estimates for linear multiple regression problems. Applied Statistics, 32 (3):345–347, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/135. See [778].

Farebrother:1984:SARa

[871] R. W. Farebrother. Statistical algorithms: Remark AS R52: The distribution of a linear combination of central χ^{2} random variables: A remark on AS 153: Pan’s procedure for the tail probabilities of the Durbin-Watson statistic. Applied Statistics, 33(3):363–366, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [851, 872].

Farebrother:1984:SARb

[872] R. W. Farebrother. Statistical algorithms: Remark AS R53: A remark on algorithms AS 106, AS 153 and AS 155: The distribution of a linear combination of χ^{2} random variables. Applied Statistics, 33(3):366–369, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [586, 851, 854, 871].

Dewey:1984:SAR

[873] Michael E. Dewey. Statistical algorithms: Remark AS R54: A remark on algorithms AS 169: An improved algorithm for scatter plots. Applied Statistics, 33(3):370–372, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [859, 868, 883].

Chou:1985:SARa

[874] Youn-Min Chou. Statistical algorithms:

Remark AS R55: A remark on Algorithm AS 76: An integral useful in calculating noncentral T and bivariate normal probabilities. Applied Statistics, 34 (1):100–101, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [382, 747, 782, 878, 883].

Chou:1985:SARb

[875] Youn-Min Chou. Statistical algorithms:

Remark AS R56: A remark on Algorithm AS 5: The integral of the noncentral t-distribution. Applied Statistics, 34(1):102, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/5. See [2].

Rice:1985:SAR

[876] Lory R. Rice and Rose E. Gaines Das.

Statistical algorithms: A remark on algorithms AS 32 and AS 147. the incomplete gamma integral. Applied Statistics, 34(3):326, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [68, 849].

Dewey:1986:SAR

[877] M. E. Dewey. Statistical algorithms: Remark ASR 60: A remark on AS 112 exact distributions derived from two-way tables. Applied Statistics, 35(1):86–88, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [618, 857].

Thomas:1986:SAR

[878] G. E. Thomas. Statistical algorithms: Remark ASR 65: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 35 (3):310–312, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [382, 747, 782, 874, 883].

Farebrother:1987:SAR

[879] R. W. Farebrother. Statistical algorithms: Remark AS R68: Constrained and unconstrained L_{1} estimation A remark on AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 36(1):118, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [588, 746].

Hill:1987:SAR

[880] I. D. Hill. Statistical algorithms: Remark AS R70: A remark on Algorithm AS 16. maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 36 (1):119–120, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/16. See [7, 593].

Shea:1988:SAR

[881] B. L. Shea and A. J. Scallon. Statistical algorithms: Remark AS R72: A remark on Algorithm AS 128. approximating the covariance matrix of normal order statistics. Applied Statistics, 37 (1):151–155, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/128. See [714].

Cran:1989:SAR

[882] G. W. Cran. Statistical algorithms: Remark AS R79: Comment on: A remark on algorithms AS 32 and AS 147: The incomplete gamma integral. Applied Statistics, 38(2):423, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [853, 849, 68].

Boys:1989:SAA

[883] Richard J. Boys. Statistical algorithms:

Algorithm AS R80: A remark on Algorithm AS 76: An integral useful in calculating noncentral t and bivariate normal probabilities. Applied Statistics, 38 (3):580–582, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [382, 747, 782, 873, 874, 878].

Dewey:1989:SAR

[884] Michael E. Dewey. Statistical algorithms: Remark AS R81: A remark on Algorithm AS 131: Tabulating frequency distributions for variables with structured code sets. Applied Statistics, 38(3):582, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/131. See [742].

Payne:1990:SAR

[885] R. W. Payne. Statistical algorithms: Remark AS R82: A remark on Algorithm AS 65: Interpreting structure formulae. Applied Statistics, 39 (1):167–175, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/65. See [297].

Berry:1990:SAA

[886] K. J. Berry, P. W. Mielke, Jr., and G. W. Cran. Statistical algorithms: Algorithm AS R83: A remark on Algorithm AS 109: Inverse of the incomplete beta function ratio. Applied Statistics, 39(2):309–310, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/109. See [589] and correction [887].

Berry:1991:CAR

[887] K. J. Berry, P. W. Mielke, and G. W. Cran. Correction of Algorithm AS R83 — A remark on Algorithm AS 109: Inverse of the incomplete beta function ratio. Applied Statistics, 40 (1):236, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/109. See [886].